A Complete Bibliography of Publications in Computers and Mathematics with Applications (1980–1989)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

25 August 2025 Version 1.05

Title word cross-reference

(R^{2}) [EC83]. 0 [Ric81]. 1 [KSTK84, Ric81]. $15.95 [Lev82]. $15.96 [Lov82]. 2 [Bos87, MS88f]. $20 [Ste81d]. $23.50 [Ste81a]. 230 [Bok88]. 2^{10} [Ste81c]. 2 × 2 [Kar81]. 3 [Fol87, dVWY87]. $39.50 [Sha82]. 3 × 3 [Fra89]. 4 [Fol87]. $40 [Ste82]. $47 [Ste82]. 515 [dF89]. $6.95 [Ste81c]. $65.00 [Eld87]. $7.95 [Ste81c]. 820 [Pau89]. 3 [FWPT87]. _{1} [KA89a]. _{3}F_{2}(1) [Wim83]. _{60}^{B} [Kro89]. A [UA85]. α [Web89]. β [Web89]. c^{0} [Ber82]. C^{1} [Blo87]. d [Wal85]. Δ^{2} [Jur83]. exp(x) [AJMT81]. G [KA89a, Zen89]. Γ [Mar86c]. I [GS81, Med88b]. K [GS81, GHN89, Mor89]. L_{2} [LM86a]. λ [Pan85c]. ln(x) [MTP82]. LU [NT85]. N [CH89, Fer82, KO88, AMR89, CK88b, HWYL88, LT87, Pur83a, TSY88, Ubh84, Ubh87]. O [Ubh84, Ubh87]. P [Sha88]. q [PM84]. R^{3} [OP87, TB87]. ⇀ ↽ [Zou89]. T [WCB85, JV88]. τ [Luk81]. → [Pou88]. wxyz [Kar81]. X [BHM88, Wai81]. XY (X^{2} + Y^{2})^{1}2 [BHM88]. Y [BHM88, SE88b]. y′′ = f(x, y) [DOS88]. y′′ = f(x, y, y′) [DOS88].

-atom [Pau89]. -axis [AMR89]. -breaking [MM86a]. -cube [HWYL88]. -curves [Web89]. -cyclic [GHN89]. -D [Fol87, MS88f, dVWY87]. -dimensional [Bos87, CK88b, Mor89, Wal85]. -factors [KSTK84]. -matrix [WCB85]. -method [Luk81]. -person [LT87]. -point [Ubh84]. -projection [Wai81]. -stable [Sha88, UA85]. -th [CH89]. -transformals [Kar81]. -unidirectional [JV88].

0-916460-08-8 [Lev82].

1 [Pou88]. 10/11 [BM90]. 11th [Ste82]. 12A [Ano86i]. 12B [Ano86j]. 156pp [Ste81b]. 17 [Kou90]. 175pp [Ste81c]. 18 [BM90, Zad92]. 1989 [Ano88e, Zad92].

255pp [Ste81c]. 2nd [O’C87a].

6/7 [Zad92].

7 [Rob87b].

85721 [Ste82]. 89 [Ano88f].

9-12 [Rob87b].

= [SE88b, Zen89].

A. [Los92, Vai88]. A.D. [Bér89a]. Abelian [Ric88]. absence [Kra87]. absorbing [GK88]. Abstract [CvdC83, Mar86a]. academic [O’C87b]. Academy [Tar88]. accelerated [Mar88]. accelerating [KW81]. acceleration [CS89, MdlS86, MS88b, Pan81b, Pan82c, Wai81, Wai82]. accuracy [CF84b, DGL88, Lay86]. achievement [Str89]. Acoustic [EG85, Ame85, Ano87-27, BBPW85, CSL85b, FG80, KTPB85, LK85, NP81, PO85, Sch85]. acoustic/elastic [KTPB85]. acoustics [Ano88-33, BGT85, BS85, Bir85, CSL85a, DeS85, Fel85a, Pie85]. action [MS88b]. activity [ML88]. acute [SMH87, SvDHH89]. Ad [Zvi89]. Adams [LP81, LJP84, RT81]. Adaptation [CLF86, dlS87]. Adaptive [Cas85a, CLS87, CL89c, HS87, HS88, BdSM88, Kan88, MdlS86, Sen88a, dlS85, dlSGA87]. adaptive-control [Sen88a]. Addendum [Ano89n, Rob87a]. adding [Sch86c]. ADI [Duf83, Wac84a, YE87a]. adjoint [Eva81a]. adjusting [Han86]. adjustment [Kit89, YNA89]. adjustments [SDB89]. adriamycin [WLR87]. Advanced [Con87a]. Advances [Ano88r, Ano88-43]. advantage [Wit87a]. advection [HS87, HS88]. aesthetic [Osb86]. Affine [Tom85]. after [MR88d, SDB89]. against [IM89a]. Age [WSR87, Bla83, Bra83, Cus85, Cus86, Eld83, Eld87, Eva88, KSW85, Lev83, Li88, Mar83, Mar85a, Mur83, Prü83, Sho80, Swa88, Ven87, Web83, YE87a]. age-dependent [Bra83, Eld83, Eld87, KSW85, Lev83, Sho80, Swa88, Ven87]. Age-size [WSR87]. age-specific [Prü83]. age-structered [Cus86]. age-structured [Bla83, Cus85, Li88]. aggregating [Pan82c]. aggregation [Eva81b]. aggressive [Jär89]. aging [AA85]. agriculture [Ste81b]. Aided [Rei81]. AIDS [Tho89]. aim [SA87]. air [LRG89, MPW87, RLM^{+}87, ZB88]. air-combat [MPW87]. Aircraft [PR87, AR87c, Jär89]. airfoil [CF88]. Aitken [Jur83]. ale [Nic87]. Algebra [Rei81, Pan85c, Wil87]. Algebraic [Pan87a, Pan83, Pan85d, Sla86, TG88]. algebras [BM84b]. Algorithm [Fra89, AS88, AS89, AR87b, AMR89, AB81b, BM89a, BB87, Boz81, Cas83b, CF84a, EW84, EW86, EW88, EO82, FHS84, GM88, Hag88, Jul88, KTW83, KLT80, Kob89a, Lin89, Lou83, OP89, ODS86, Pan85b, Pas82, Pur83a, Pur83b, Rab86, Rus89, SB88, SSS87, Sha81a, Ubh87, WW89a, YE86, YE87a, YE87b]. Algorithmic [VO85, Ng81]. Algorithms [TSY88, BL81, BP86, BS80b, CB88, Cod88, CP88, CdC85, Dev81b, Dev81a, HGP86, KA89b, LLM84, LO82, MdlS86, Muk81, Ort86, Pan82b, Pan85d, Pan85c, Pet89, Pug89, SLPC87, Ubh83, Ubh84, ZN88, Ste81a]. Alhambra [PG87]. all-furthest- [Tou83]. allocation [McL86, RPL87, TC89]. alloys [Pau89]. Almost [Blo88a, CKB83, CK89]. alternate [Bas89]. Alternating [Eva81a, CS88a, Hay80, YE87b]. alternating-direction [CS88a, Hay80]. alternative [KW89, Mei86a]. Amsterdam [Ste81d]. AN-stable [Zed87]. analogue [IM89b, TT89]. analogues [Bro89]. Analysis [Ame88, CN89, LL89, PAC85, SHPT87, St.85, Ano87-29, BM85a, BM85b, CW85, CdC85, EL88, Gra87, Gre87a, Gri83, Had81, Jur83, KKK88, KO88, Kou89, Kou90, Krz86, Li89, LV86, Lou84, ML88, MR88a, MS88b, MS88c, Mit89a, MRV85, MM86b, ND88, O’C87b, RL88, Ros87, Ser80, Sha81a, SI89, SD87, TG88, Tuz88, Ver89, Vla86, WAYU88, O’C87a]. Analytic [Zad89, Zad92, SF88]. Analytical [KESC89, BD85, CF88]. analyzed [BGG^{+}87]. analyzer [Had85]. analyzers [SHPT87]. Anderson [O’C87a]. aneurysms [MS88e]. Angell [Con87a]. angle [Ano87-35]. Announcement [Ano80a, Ano83a, Ano88b]. Announcements [Ano88c]. annual [Ano88r]. answers [Ste81c]. antenna [LK85]. Antisymmetry [Shu88a, Zam88]. APA [Pan82b]. APA-algorithms [Pan82b]. apparent [Pau89]. appear [Ano86k, Ano89r]. appendices [Ste81c]. Applic [BM90, Kou90]. Applicable [Ano87c]. Application [BD85, CKB83, IE88, MN88, MdlS86, Neu80, Rok85, Rud88, WAYU88, YE87a, AS88, Akr81, And86, Ano89n, Asm87, BKL88, GZ89, HC89, LL88a, OP86, Pan85d, PB82, PB81a, PB81b, Rob87a, Rob87b, SF88, SI89, Sod85, UM85, Yak89]. Applications [Ano88f, Ano88e, Ano88d, (Co85, McD85, TR88, Tuz88, Zad92, Ano84a, Asz80, BEW86a, CM88, Cus83, DD83, Ede80, GP89, Gyö89, Mar86a, RW87, RR88a, Sha81a, SU88, Wit83c, Ano88r]. Applied [Mea86b, Web86b, BS85, Eld87, Lam86, Mak89, MC85, Møl86, SCSS85]. Applying [IR89]. appraisal [Rod88a]. approach [AS89, Ado87a, Ado88a, Ami89b, Ano89n, AR87c, DV87, Har86d, Kou89, Kou90, Mar83, McC83, MM88, Ng81, OS84, PM87, Per87, Ran88, Rob87a, Rob87b, Sen88a, Sod87, TD88, YNA89, Zad83]. approaches [WCB85]. approx [Lev82]. approximants [AS88, AS89, Cuy82, CvdC83]. approximants-the [AS89]. Approximate [Ben85, Cho85, HZ86, BSPSM81, Lip84, Pan82a, Pan87b, PR86b, Sar84, Sar88, ZL80]. Approximation [PB82, SSB88, Ell86, Fol87, Gyö88b, KLT80, Man87, MS88a, Mei86a, OvD89, PMS86, SL88, Sza87a, Sza87b, Tom85, Ubh83, Ubh84, Ubh87, Was84]. approximations [Bal86b, Bal88, BE85, CK89, Die84, DS85, FGP83, Hay80, Lay83, Lay85, OK82, Pie85, Ser80, TG88, Vic85, Wer86]. aquatic [Ant89]. Arabian [CARG89, LBC82]. arbitrarily [Ser80]. arbitrary [Bha82, BP86]. archaeological [Bér89a]. archaeology [BHW86]. architectures [CP88, Pug89]. area [Lou84, RPL87, RL88]. Argot [All83]. arguments [AC86]. arising [FS80, LK85, McC85, MN83, Ras85]. arithmetic [BB83, Bos87, JáJ85, Pan82a]. Arizona [Ste82]. ARMA [Ben85, BM85c]. Arrangements [Tel89, Chu84]. array [Sch85]. arrays [CP87, SFK87]. art [Bér89a, Mak86a, McW89, MM86a, Sch86a]. articles [Jay92, Los92]. Artificial [RLM^{+}87, GB85, (Co85, MC85, Ste87]. arts [Róz86, Yu89]. Asada [Gra87]. ASE [Ano88f]. Aspects [SL85b, CBC89, RR89, ST88]. assembly [Pou89]. Assessing [McL86]. asset [Cer89b]. assignment [Eva81b, SB88]. assistance [BN81]. Associated [Rei81, Blo85, FN86, Web83]. associations [Vai86]. asymmetric [AE87, JM89, Mei86b]. asymmetries [San86]. Asymmetry [Ács89, Her86, Don86, PT86]. Asymptotic [Bla83, Bro87, FMWW86, Ara83, BN81, BC86, Mei86a, Mór81b]. asynchronously [Wit83c]. atmospheres [KESC89]. atmospheric [CS88b, TR88]. atom [Pau89]. Atomatics [Tuz88]. atomic [Asl88]. attractivity [KLS89]. authentication [HHL89]. Author [Ano85b, Ano85c, Ano89b, Ano86i, Ano86j, Ano87y, Ano88-30, Ano88-31, Ano89q, Ano85d, Ano86b, Ano87d, Ano87e, Ano88g, Ano89c]. authors [Ano84d]. automata [Pic89b]. Automated [LRG89, Pou89]. Automatic [Dal85, GR88, KKS86, KT86, Mur87, VO85, Omo82, Wac85]. autooscillations [SM80]. autoregressive [SSB88, WAYU88]. auxiliary [Sza87a]. available [Bur82]. avalanche [PPR^{+}89]. Avar [Bér89a]. average [Dev81b, Dev81a]. avoidance [CLS87, CL89c, MP89]. axis [AMR89]. Ay [SE88b].

B [Mea86a]. Babel [Cho89]. backtracking [Lou83]. backward [Cas83a, Ver86]. bacteriophage [RTF^{+}88]. bags [BPXY88]. Bairstow [AV82]. balanced [CS88b]. balsam [ByMS86]. Banach [GP85, TW80]. Banyan [Jul88]. bargaining [ERH89]. Barr [Ste87]. Barrier [SS87, PM87]. Barting [Sch86a]. base [HF86, WJMM83]. based [AAM86, CA86, Hag89, Hav83, KO88, LL88b, Lip84, LV86, LO82, SDS81, Sen88a, Wer86]. Basel [Mea86b, Web86b]. basic [STY83]. Basin [Bér89a]. basis [VO85, vNvN89]. Basset [BT89]. bathymetry [BBPW85]. batik [Haa89]. beam [Ber82, Ver86]. bearing [HLHG89]. beauty [Pic89b]. bed [PAC85]. behavior [Bla83, BC88, CO83, FMWW86, Lay85, Pou88]. behaviour [Bro87, Ira87, Prü83]. Behrens [Cho85]. being [Rod88a]. Bellman [Ano86l, KK86, Pan86]. Benchmark [TT86, Ano85i]. benefit [SI89]. benefits [SDB89]. benzenoids [CBC89]. Berlin [Kam87]. Bessel [GS81, PB82, Pie84]. Better [Tur88]. between [AG80, CK88a, GHN89, MT85, Sla86]. Beyond [Cho89]. Bézier [Scu89]. bias [But85, SDB89]. bibliography [Rod87d, Rod89a]. Bidiagonalization [Hag87]. bidimensional [OP86]. Bifurcation [Cus83, MPL86, Qin86, Swa88]. biharmonic [Lot87]. Bilateral [San86]. binary [AR87a, Pur83a, Vla86, Wal85]. binomial [HP89b]. biochemical [FS83]. bioengineering [Nic87]. bioingineriei [Nic87]. biological [Gre89, Vai86]. biology [Ras85]. Biomathematics [Gro87]. biomechanics [MS88d]. biomedical [Koh89]. biopolymers [LRBK89]. biquintic [WA89]. birth [Let87]. bisection [GM88]. bisexual [Kra87]. bit [Cod88, Pan81a, Pan82a, Pan85c]. bit-cost [Cod88]. bit-operation [Pan81a, Pan82a]. black [Sch86b]. blade [IR89]. blending [Sch86c]. Block [Cas85b, STY83, TSS88, SGN87]. blocking [Lau80]. blood [MS88b, SvDHH89]. board [Ano85f, Ano80b, Ano81b, Ano82b, Ano83b, Ano84b, Ano85e, Ano86d, Ano87j, Ano87k, Ano88u, Ano88v, Ano89j, Ano89k]. bodies [Pet89]. body [MS88d]. boll [FC83]. Boltzmann [TB87]. bonding [Mak86b]. bone [MR89b]. Book [Ano86c, Ano87f, Ano87g, Ano87h, Ano87i, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Col82, Con87a, Eld87, Gra87, Gre87a, Gro87, Kam87, Lev82, Lov82, Mea86a, Nic87, O’C87b, O’C87a, Rod84a, Rod84b, Rod84c, Rod85a, Rod85b, Sha82, Ste87, Ste81a, Ste81b, Ste81c, Ste81d, Ste82, Wal87]. bookbindings [RR89]. Books [Ano81a, Ano82a]. Boolean [HWYL88]. Both [Ste81c]. bottom [HKW89]. boundaries [IM89a]. Boundary [Aga88, Ado83, AC86, AW89, AAM86, Ano87w, BE85, BP85, BN86, Cas83b, Cas84, Cas85a, Cas86, DD89, Dey88, ES87, Eva88, Fit83, FGP83, HK85, JT89, KKRS84, KKK88, Kha86, Lay83, LE81, Lip84, McC85, NP81, Roh83, SP89, STM87a, TR88, UW80a, UW80b, Usm85, VP85]. boundary-value [AC86, AAM86, Cas86, HK85, Lip84, Usm85]. bounded [LM86a]. boundedness [LL88a]. bounding [Gip85]. Bounds [AWYU82, Ara83, Avi81, HS81, JáJ85, Kra87, Pet87, PM84, Tro80]. Boussinesq [IM89b]. Box [Ste81c]. bracketing [Hag89]. branch [EW86]. branches [Cus85]. branching [Sho80, TSJ89]. break [AB81b]. breakdown [PPR^{+}89]. breaking [MM86a, PPR^{+}89]. breast [TO87]. brickwork [Mak89]. bridging [XS87]. Brief [Zel92]. brine [CW85]. Brockington [Ste81b]. broken [Man86]. brown [SMH87]. Brownian [Lou84]. bucketing [Dev81b]. Buckling [Tar89]. buckminsterfullerene [Kro89]. Bucuresti [Nic87]. budworm [ByMS86]. budworm-balsam [ByMS86]. Building [Ste82]. built [AB87a]. built-in [AB87a]. Buoyancy [Wan88, CW89, Wan87, WW88]. Buoyancy-induced [Wan88]. Burgers [OP86]. Byrnes [Sha82].

C [Col82, Mac86a, Sha82, Sni88, Kro89]. C-programming [Sni88]. C. [Jay92, Zel92]. Cagniard [MU82]. calculating [Sey87]. calculation [BM84a, Dal85, Wac85, YK87]. calculations [Ede80]. calculators [Lev82]. calculus [BEW86a, BEW86b, Ede80, HP87]. Calibration [Har86d]. Calif [Ste87]. California [HC89]. Call [Ano88r, Ano88p, Ano88q, Ano83a]. called [Pau89]. Cambridge [Lov82]. Can [Sza89]. canceling [Pan82c]. cancer [GGR^{+}87, Per87, Rig87, SDB89, TO87]. cancer-therapy [Rig87]. cannibalism [Lev83]. canonical [FHS84]. capability [Ano87-35]. capable [Big88]. capacity [Pou89]. capillarity [Gre88a]. Carbon [Bal89]. Carleman [Fit83]. Carlo [CA86]. Carmichael [WH82]. Carpathian [Bér89a]. carry [Pur83b]. carry-free [Pur83b]. cars [YdV89, YV89]. cascade [CF88]. case [Avi81, BK87, KW89, Mar89, MdlS86, MS88f, PB81a, PB81b, Tel89, TK89b, YdV89, YV89]. catalog [Cha89]. categorical [Krz86]. category [WLL88]. cattle [HC89]. Cauchy [DS85, Ger82, Ger86, GS87b, Rab86]. Cauchy-type [DS85]. causal [Ano89n, Rob87a, Rob87b]. causality [dM84]. cavity [Gre87b]. cell [BGG^{+}87, BM89b, BM90, DHT86, FWPT87, KA89a, Pou89, Rig87, SvDHH89, Web89]. cell-size [DHT86]. cells [AK89b, DV87, Rig87]. Cellular [AA85, Her86, Mur83, Wit83c]. Century [Bér89a]. certain [FMWW86, OB88]. chains [Chu88b]. Champaign [Lev82]. channel [Gre88b]. channels [LM86b]. Chaos [Ano88s, Bla86, CSUZ89]. characteristic [Dev81c]. characteristics [CLF86, KS85, LCG86, MS88c, Zem85]. characterization [SA87]. characters [dW88]. Charles [Lev82]. Chebyshev [GR88, Ger86, GS87b, HK85]. check [Bos87]. checkers [JA87]. checking [JA87]. Chemical [Len89, Bal86a, MLDR87, MR88a, Pea86, Sla86]. chemist [Ács89]. chemistry [Sla86, SU88]. chemotherapy [GGR^{+}87]. Chi [Har89b, MC84]. chi-squares [MC84]. child [SvDHH89]. children [Ha89]. Chinese [Ha89, Yu89]. chips [XS87]. Chirality [Sok86]. Choice [EC86, CF84a, Sod85, Tho86b]. Christian [Ker89]. chromatic [HS81]. chromatographic [PAC85]. Chuan [Har89b]. circle [Eng88b]. circuit [Lir89]. circuits [AR85, AB87a, JáJ85]. circular [KTW89, San86, SRP87]. circulation [STP88, SM88a]. circulation-preserving [STP88]. citation [Ano87-30]. Clarendon [Ste81b]. class [Ami89b, BR86, Blo87, Blo88b, CB88, Fra80, GS87c, Han86, HL89, KSTK84, PR85, STY83, SGN87]. classes [KS85, Sha86a]. classical [AB81a, Blo85, MWBM89, MRV85]. Classification [Jen86, Cho86, EC86, Gup86, Kno86, Mak89, O’N86]. classifier [Han86]. Cliffs [Col82]. clinical [Gre87a]. Clonal [AK89b]. closed [Jär87]. closed-loop [Jär87]. closest [Tan89]. Closure [Rei81]. coastlines [Kap86b]. code [Fel85b, JV88, SDS81]. Codes [Rao84, Ano87z, BN81, Bos87, Her88, SBB83]. coding [O’N89]. coefficients [BB84, BG88, HP89b, Kar87, MM84, OS84, SSB88, Sod87]. cold [CW89, Col85, WW88]. collection [Ano85i]. collision [LM86b, MP89]. collision-avoidance [MP89]. collocation [AW86, BT89, Hay80, Yan87, Yan89]. colony [Mur83]. Color [Sen88b, CK88a, Zou89]. Colored [eZ86, Sch86b, Was86]. coloured [Mak86a]. columnar [Gra86]. combat [GP89, Jär89, LRG89, MPW87, Raj89, RLM^{+}87]. combination [MC84]. combinations [Lin89, Pan81b]. combined [KLT80]. combustion [Sod85]. comment [Ado88a, Zel92]. Commentary [Jay92]. Comments [Abe84, Los92, Wit83b]. common [Che87a, Los89b, Pur83b]. communication [Lau80]. communications [AWYU82, LM86b]. commutative [BL81]. companion [STY83, SGN87]. Company [Mea86a]. Comparison [LL88b, MLDR87, Zla82, Cas86, Gyö89, HS81, KW81, MSH80, MM86b]. Comparisons [HTY87]. compartmental [BM89b, BM90]. Competing [Smi89]. complete [Ver89]. completion [Che87a]. Complex [TSY88, EL88, LL88c, McC86]. Complexities [Tak88]. complexity [Dev81b, Dev81a, Pan81a, Pan82a, Pan85c, Pan85b, Pan87a, Pan87b, TW80, Tro80]. components [Los89b]. composed [YK87]. Composite [MRV85, BD85, MDR85]. Composites [BM85a]. compressible [FS85]. Computation [Die84, Fra80, Gin80, Hag88, MS81, PB86, BB87, BD85, CH89, ERH89, MP86, PB82, Pie84, VO85, Wer84, Col82].

Computational

[Ano84a, Fle88, Ano88-43, Ano88-33, Ano88-44, Ata85, BN81, DeS85, Dev88, Dun80, Gli85, PP89, Sla86, WFW81, Zad83]. computations [Kan88, Pan85c, PR86a, SL85b]. Compute [Fra89]. Computer [Ano88-34, BE88, BN81, Boy86, Kam87, KSTK84, Koh89, Rei81, SMH87, Ano85i, Ano87z, Asz80, BGG^{+}87, Dev81c, Dev82, Ede80, Exo88, FBS88, Har86c, IE88, Iwa87, Lau80, ML83, Pie84, RL88, RW87, RR88a, Tuz88, Con87a, Ste81b]. Computer-Aided [Rei81]. computer-analyzed [BGG^{+}87]. Computerized [PB81a, PB81b, BCC^{+}86]. Computers [BM90, Che89, Kou90, Ric88, Zad92, Eso88, GB88, NT85, RR88b, Ste81c, Wil82, Wit87b, Ano88r]. Computing [Bur82, TSY88, Eld83, EW86, FHS84, Pan87a, Ral87, TSS88, Usm85, WW89a]. concealability [KL84]. concepts [Mye87, Sla86]. conceptual [Sow83]. concerning [Dem82, Sza87a]. condensations [MP88]. condition [BE85, Blo87, CK89, KKK88]. conditional [Sza89]. conditioned [Ber88]. Conditions [Che84b, Ado83, Ano87w, BP85, Dey88, ES87, FGP83, GS89, Lay83, Qin86, VP85, Yan89]. conducting [MB89]. conduction [Blo86, BD85, MR88d, Mur88, Mur89]. Conference [Ano88-34, Ano84a]. Confluent [Sal84a]. conforming [GNP82]. confounding [Ano89o, RM87]. conjugate [Hag89, SvR88, ZN88]. conjunct [Omo82]. conjunctive [DP88]. Connectionist [Par88]. connectivity [Abe84, Gar82, Sha81a]. conquer [AS89]. Conservation [Sha86b, HHHK88]. conservative [HL86]. conserving [FS85]. considerations [Mar83]. considering [KKK88]. consistence [Tho80]. Constable [Lov82]. constant [Gei81]. constants [BP86]. constrained [HRA89, Pra80, Rus89]. constraint [Mar85b]. constraints [AF87, CF84a, VMO89]. constructing [AB81b]. Construction [Lin86, AHL83, PM87, Ser80]. constructions [Fle88]. containing [Ano88-37]. containment [YdV89]. content [Che87b]. Contents [Ano85h, Ano86i, Ano86j, Ano87y, Ano88-30, Ano88-31, Ano88-37, Ano89p, Ano89q]. contexts [Was86]. Contingency [SCH87]. continued [CdC85]. Continuous [LL88a, Sar84, Ber89b, CW86, DOS88, Krz86, Li88, Mar86c, MB89, Sar88, SLPC87, TT86, Vla86]. continuous-time [SLPC87]. continuously [ODS86, Tho87]. Contopoulos [PB81a, PB81b]. contractions [Riv86]. contributions [Ano86l]. control [ÅH86, Ano89n, Ant89, Asz80, Chu88a, CLS87, IE88, LL88a, LK83, Mar85b, McD83, Med88a, Med88b, MdlS86, Pan86, Pou89, Qin84, Rob87a, Rob87b, SFK87, SE88b, Sen88a, STY83, SGN87, Str89, YV81, dlS85, dlSGA87, Gra87]. Controllability [KW89, Qin84]. controlled [FN86]. controllers [CL89c]. controls [Jär87, YJ81]. convection [JM84]. convention [EA85]. Convergence [SZ89, TW80, WO87, And86, CK89, Chu88b, KW81, LM86a, MdlS86, SB88]. convergent [Rus89, Sny89]. converging [EW86]. Conversion [Rao84]. conversions [SLPC87]. convex [ATB82, Dev81a, Dev82, DD83, MT85, OP87, Ubh83, Ubh84, Ubh87]. conveyance [EL88]. convolution [FMWW86, JM82]. Cooperation [Tar88]. Cooperative [Mar86b, Tan89]. coordinates [SD87, Ste88]. coreflectors [WLL88]. corner [CF84b]. coronoids [CBC89]. correction [Cas83b, KLT80]. corrections [Cas86]. corrector [DD83, Kha89]. correlation [SS88]. correlations [Web89]. corresponding [Ber82]. corridor [dPLBA87]. Corrigenda [Ado88a]. Corrigendum [Ano88t]. cortex [GT88]. Cost [SI89, Che84a, Cod88, YdV89]. Cost-benefit [SI89]. cotton [FC83]. Could [San86]. counter [PAC85]. counter-current [PAC85]. counterexamples [Hua89]. counterpoint [MWBM89]. country [Loe86b]. coupled [Gro88, KTPB85]. coupling [BD85]. course [Kap86a]. court [Loe86b]. cover [Lev82, Ste81c]. Covering [GH88]. crack [MR89b]. crafts [Yu89]. cranial [MR88b]. create [Sch86b]. criterion [JT89]. crop [FC83]. cryptography [HHL89]. cryptosystem [AB81b, Web88]. cryptosystems [AB81b]. Crystal [Asl88, GK88, Bok88, Kop88b, Pau89, Sen86, SU88, Zvy88]. crystallization [SBCC86]. Crystallographic [Mam86, Hau88, Mol88]. crystallography [Eng88a, Fic86, Hau88, Mac86c, Rud88, Vai88]. crystals [Roc88, TT89]. CSRL [BMC85]. CT [LV86]. cube [BCC^{+}86, Bur82, HWYL88]. cubic [BE88, Che84b, PR86b, Pau89]. Cubical [Har88a]. cultural [Was86]. Curie [Bra86, Shu88b]. current [PPR^{+}89, PAC85]. currents [CLF86, CARG89]. curve [SM80]. curved [BM84a]. curves [Scu89, Web89]. cycle [BM89b, BM90, KA89a, Rig87]. cycles [BE88]. cyclic [CS89, GHN89]. cylinder [SRP87]. cylinders [OB88]. Cylindrically [Kup87]. cytofluorimetric [Ros87]. cytometric [BGG^{+}87].

D [Sha82, Wal87, Fol87, MS88f, dVWY87]. D.E. [Yac82]. Dalton [Wac85]. damage [KO88]. damped [Ser80]. dance [Loe86b]. Dante [dF89]. data [Ano88-37, Ano88-35, Ano88-36, Ano89u, Ber89b, CW85, Die84, Fol87, Fra82, FBG88, HTY87, Har86d, HF86, Krz86, Los89a, Ros87, STY87, Sha81a, TO87, WJMM83]. databases [Bai89]. date [Che87a, Che87b, CL89b]. daughter [Web89]. day [Fel85b]. death [Let87]. decagonal [Pau89]. Decell [KTW83]. deception [Yav87a]. decision [AR87a, BL81, LRG89, Oda88]. decomposition [Ado88b, Ger86, HF87, Mao86, TM87]. deconvolution [DT88, Mor86, Mor89]. decoration [Cam89]. deduction [Omo82]. default [RC83]. defect [Pic89b]. defense [Squ81]. deferred [Cas83b, Cas86]. deficiency [HP89a]. Definite [Fra89]. Definition [dW88, Her88]. deflection [UW80b]. degree [KSTK84]. degrees [PB81a, PB81b]. Dekker [Eld87, Mea86b, Web86b]. delay [AM81, CW86, Gyö89, LL88a, Mór81b]. delays [dPLBA87]. Delicate [Pac89]. Dense [PR89]. Density [Mur83, BR86, Kri85, MM86b]. dependent [Bra83, BN86, Chu88b, Eld83, Eld87, Kar87, KSW85, LM80, Lev83, Mar83, Mar85a, Mar86a, Men86, Mur83, Sho80, Swa88, Sza87a, Sza87b, Ven87, Web83]. depth [BS89, MWBM89]. depths [JáJ85]. Derivation [WSR87, Bok88, Fer82, Kri85]. derivative [BE85, BSPSM81, Hag89, SDS81, Sal84a]. derivative-based [Hag89]. derivatives [BHO82, BEW86a, BEW86b, CS88a, EC83, KT86, MSH80, VO85, Zad89, Zad92]. derive [San86]. derived [Pic89b]. Descartes [HP89a]. describing [Mar85a]. Description [McD83, EGZK87, SE81]. Design [CKB83, HP87, Ste82, BK87, CK87, Cho89, Gre87a, Kap86a, MC85, McW89, Pan85d, SFK87, dlS87, dlSGA87]. detect [MHM87]. detecting [JV88, XS87]. Detection [Abe84, Gar82, LLM84, Tur88]. determinant [HWL89]. Determination [BC88, Che87b, CL89b, EO82, Orl85, OP89, SLHC89]. determine [KTW83]. Determining [CARG89]. deterministic [SL88]. development [Kra82, Lum87, Mei86b, ST88, STP88, TSJ89, Wil82]. developmental [DHT86]. deviating [AC86]. deviation [Che87a]. Dfl [Ste81d]. diagnosable [CK87]. Diagnosis [SA87, BMC85]. diagnostic [Lir89]. diagonalization [Hag87]. Diagonally [Cas84]. diagrams [AR87a]. diameter [KK87]. dictionary [Sod87]. Didday [Ste81c]. dielectric [Blo86, Blo88a]. Difference [CSL85a, AC86, Ano87c, Ano87u, DD83, Dey88, EO82, Gip87, HL86, Kov80, LBC82, Lay85, LM86a, Luk81, LTW83, McD85, Mil86, PO85, PR85, Sal84a, STM87a, St.85, Usm85]. differences [Mur89]. differentiable [Che85, Sza89]. Differential [And86, Mea86b, Rei81, Web86b, Ado83, Ado87a, Ado88a, AC86, AG87, Aga88, AW86, AB87b, AM81, Ano84c, Ano88r, Bel83, BZ89, Bru87, CL89a, Chu88a, Cuy82, DD83, Dey85, Eva81a, Ewi80, Fat84, Fat85, Gai86, GS87a, GZ89, GP89, GS87c, Gyö88a, Gyö88b, Gyö89, HD83, HRA89, Ira87, Iwa87, Jär89, KLT80, Kha89, KA89a, Kos88, LP85, Lam86, LO86, MN88, MR89a, MP89, MR88c, MS88f, MU82, Neu80, OS84, OP86, PM87, PR85, PM84, RZ86, Rok85, Sar84, Sar88, Sat85, SZ88, SS89, St.85, Tho86a, Tho86b, UM85, UA85, Wit83b, Wit83c, Yav84, Yav87a, Yav87b, ZL80, Zed87, Zem85, ZWS85, dVWY87]. differential-integral [Kos88]. differentiating [VO85]. differentiation [Cas83a, KT86, Lum87, Mur87, Sha81b]. differentiators [HS87, HS88]. diffuion [Gip87]. diffusion [BM89b, BM90, EA85, KSW85, Mac83, MS88f, Roh83, Sod85, Sod87]. diffusion-convention [EA85]. diffusion/drift [Roh83]. diffusions [FN86]. diffusive [LM86a]. digital [dlS85]. digraph [MP88]. dihedral [WD86]. Dilithium [Ste81c]. dilute [MLDR87]. dimension [HHHK88, Zam89]. dimensional [AR87c, BBPW85, BJ87, ByMS86, Bos87, Bro89, CLF86, CP87, CK88b, DV87, FS85, HF86, LCG86, LM86a, Lin86, MdlS86, Mor81a, Mor89, ND88, SL85b, Wai81, Wai82, Wal85, ZA88]. dimensionality [LRBK89, NS88]. dimensions [GS86, Har88a, St.85]. direct [Hau88, Lip84]. directed [McD83]. direction [CS88a, Eva81a, Hay80, SvR88]. Dirk [AR87b]. disc [AG80]. discontinuities [Blo88b]. discontinuous [He87, Kob89a, Kob89b]. discounted [LT87]. discourse [Sid83]. discrete [Bal86b, Bal88, Ber89b, BCP87, BDK86, BC88, DS81, KW89, Mur87, MH89, Sen88a, SLPC87, Ubh84, Wu88, Yan87, Yan89, dlS87, dlSGA87]. discrete-time [SLPC87]. discretisation [AM81, PM84]. Discretization [Kup86]. discriminant [CMN88, CNM88, Krz86, MM86b, Vla86]. Discrimination [Cho86, AL86, CA86, NR86]. diseases [SCH87]. disjoint [MT85]. disk [KTW89]. disorder [Fic88]. dispersion [DGL88, ST88]. displacement [BS80a]. displaying [STY87]. dissipation [Ame88]. dissipative [Mar86a]. distance [MT85]. distances [ATB82]. Distributed [LM86b, Ste81c, AB87b, CL89a, Dev80, Gri83, MB89, Pou89]. Distribution [NR86, DHT86, Mar86c, MC84, SB88, SRP87]. Distribution-free [NR86]. distributions [CNM88, DBH88]. disturbances [CSD88, Sza87a, Sza87b]. diverging [EW86]. divide [AS89]. divide-and-conquer [AS89]. dividing [Wit83c]. Division [LLM84, BP86]. divisor [Pur83b]. DNA [Rig87]. docking [Dav86]. domain [EC83, MK85, SCSS85]. domains [LM86a, SB85]. Dordrecht [Sha82]. doses [Rig87]. dot [BS89]. dot-depth [BS89]. drawing [Col85]. drift [Roh83]. driven [CARG89, HKW89]. Dual [ÅH86, TM87]. Duality [BM81, HP89a]. due [CSD88, Che87a, Che87b, CL89b]. due-date [Che87a, Che87b, CL89b]. duration [TH80]. during [Mei86b, PPR^{+}89, San86]. dyadic [BEW86a, BEW86b]. Dynamic [AB81a, Loe89, NS88, Ami89b, Ata85, Bel83, BKL88, CM88, CF84a, Die89, EW86, EW88, EL88, LT87, LK85, Sen89b, STM88, Sny89, Tan89, Zad89, Zad92, dPLBA87]. dynamical [Kra87, PB81a, PB81b, ZA88]. Dynamics [Web87, AMR89, Ano88-43, Cus83, Eld83, Eld87, GS87a, HD83, Kov80, KLS89, KSW85, Kup86, Kup87, Mar83, MPL86, Web83, Web89, WSR87, dlSGA87].

E-M [Ros87]. E. [Los92]. early [Cum86, San86]. easy [Was84]. econometric [Chr89]. economic [AR89, CB89, KW89, YNA89]. economics [O’N89]. economy [Tót86]. ecosystem [ByMS86]. ecosystems [Ant89]. ectotherms [ST88]. Ed. [Web86b]. edge [LB86]. Edited [Gro87, Sha82, Ste81d]. editor [Rod87c, Rod88b, Wac85, Wit87b]. Editorial [Ano85f, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano88w, Ano88x, Ano88y, Ano88z, Ano88-27, Ano88-28, Ano89l, Ano89m, Ano80b, Ano81b, Ano82b, Ano83b, Ano84b, Ano85e, Ano86d, Ano87j, Ano87k, Ano88u, Ano88v, Ano89j, Ano89k]. Editura [Nic87]. edn [O’C87a]. Eds. [Mea86a, Mea86b]. Education [Ste82, GG87, (Co85]. EEG [MWBM89]. Effect [MB89, Ano89n, Rob87a, Rob87b, Tho86a]. effective [Gip87]. Effectiveness [MHM87]. Effects [Dey88, ByMS86, MDR85]. efficiency [DS81, EC86, RT81, Sen89b]. Efficient [LY88, PR89, SJ85, SZ88, WW89a, AB81b, Ewi80, Fel85a, FS80, PR86a, Pan89, SB88]. efficiently [JA87]. eigenlength [MN83]. eigensystems [Ger80]. eigenvalue [LO86, NS88, Pan83]. Eigenvalues [Kob89b, EO82, Usm85]. elastic [CSD88, KTPB85, SP89]. elasticity [KTPB85, YK87]. elastohydrodynamic [WO87]. elastostatic [BS80a]. elastostatics [OK82]. electrical [LdT85]. electromagnatic [Blo85]. electromagnetofluid [STM88]. electromagnetofluid-dynamic [STM88]. Electronic [Lir89]. electrostatics [UM85]. element [BJ87, BB84, BE85, BG88, CN89, Dal85, DGL88, FLL82, FGP83, GNP82, Hay80, KKK88, LB86, Lay85, Lay86, LL88c, Man81, Man87, MR88a, NP81, OK82, Ste88, TR88, Tho80, TG88, Vic85, Wac85, WO87, Zam89]. element-spectral [Ste88]. elementarity [Gav89]. elementary [BHM88, Man86]. elements [BM84a, Ber82, BN86, Che84b, FS85, MRV85, OP86, Pau89, Tom85, Col82]. Elliptic [Ado88b, BB84, BE85, BG88, Eva81a, LY88, LE81, Lip84, MN88, Man81, PR85, Rud88]. Elsevier [Mea86a]. embedded [AR87b, Cas85b]. embryogenesis [San86]. enciclopedica [Nic87]. enclosures [BD85]. encryption [KL84]. energy [Gav89, McD85]. Engineering [Mea86b, Ano88f, Ano88e, Ano88d, MR88a, RW87, RR88a, RR88b]. England [Ano88e, Ste81b]. Englewood [Col82]. Engquist [Mea86a]. Enright [AR84]. Enskog [TB87]. entropy [BR86]. Enumerating [BM84b]. Enumeration [AHL83]. envelope [GS89]. environment [CLS87, CL89c]. environmental [RW87]. environments [SJ85]. EPI [Rig87]. EPI-genetics [Rig87]. epidemic [BC86, Tho89]. epidemiology [Ano89o, RM87]. epidermis [FWPT87]. EPR [Roc88]. epsilon [VMO89]. epsilon-constraints [VMO89]. Equation [BGT85, Rei81, Ado86, Ado87a, Ado88a, Ame85, BE85, BdSM88, BDK86, BT89, CSL85a, CSL85b, CK89, EA85, Fat84, Fit83, Gyö89, HS87, HS88, IM89b, Kos88, Kri85, LK85, LM86a, Lev83, LTW83, MS85, McD85, MK85, Mor81a, MS88f, OP86, PO85, SE88a, SSB88, STB88, SL85b, Sin83, Soc88, St.85, Tho86a, Tho86b, TB87, Wit83c]. Equations [Mea86b, AE87, Ado83, Ado87b, Ado88b, Ado88c, AC86, AG87, Aga88, AW86, AM81, AMR89, Ano84c, Ano88r, AFM88, BJ87, BB84, Bal86b, Bal88, BCG87, BP85, Bel83, BG88, BZ89, BS89, Blo85, Bru87, Cah88, CN89, CLF86, Chu88a, CW86, CS88a, Cus83, Cus85, Cus86, Cuy82, CvdC83, DD83, Dey85, Dey88, DS81, DS85, Duf83, Eva81a, ES88a, Ewi80, Fat85, Ger82, GS87b, GS87c, GP85, Gyö88a, Gyö88b, Gyö89, HD83, HGP86, Ira87, IJ80, Iwa87, JJS83, JM82, KKS86, Kar86, Kha89, KA89a, Kov80, LP85, Lam86, Lam88, LCG86, Let87, LO86, Luk81, MN88, Man81, Mar86a, Mar85b, MR89a, MB80, Med88a, MS88a, Men86, MR88c, MS88f, Muk81, ND88, Neu80, OS84, OP86, PR86b, PR85, PM84, Pur83a, RZ86, Rok85, Sar84]. equations [Sar88, Sat85, SZ88, SP89, UA85, Vic85, Wai81, Wai82, Wer86, Wit83b, WFW81, Yee86, YE86, YE87b, ZL80, Zed87, Zem85, ZWS85, dM84, Ger86, Web86b]. equations-1 [MS88a]. equilibrium [BG88, Cus85, Gai86, Har88c, Mar86b, Str89, TH80]. equivalent [Che85]. Ergodic [FBS88]. Errata [Ano89n, Ano89o, BM90, Kou90, Rob87b, Zad92]. Erratum [Ano86e]. Error [AW86, Bal88, BCG87, Gin80, HF87, Jur83, Ara83, CMN88, CNM88, CdC85, DLMP89, Han87, JV88, Kno86, Mor89, O’N86, Oza81, Pet87, PM84, Sha86c, Tho80, WA89, WO87]. error-detecting [JV88]. error-free [Mor89]. errors [Ali87, Gip85, LM86b]. escape [GS89]. Escher [Mac86a]. essay [Doc86]. essential [FGP83]. estimate [RZ86]. estimates [Bal88, BCG87, WA89, WO87]. Estimating [Han86, TH80, Kra87]. Estimation [CMN88, Oza81, AW86, BSPSM81, CNM88, DBH88, DLMP89, HF87, Kno86, Sha86c, Vla86, Zad89, Zad92]. estimator [Han87]. Euclidean [AS88, Avi81, Pet89]. Euler [BP85, Lam86, Yee86]. Eulerian [Kup86].

Evaluation [AJMT81, MTP82, BSPSM81, GS81, LY88, Pan82a, Pan85d, Pan87b, Pan89, Rig87, WBZB88]. evasion [AR87c, CN87, CLS87, CL89c, GS87a, GZ89, GEG87, Jär87, LO89, Mer87, MP89, MPW87, Pac87, PR87, Rod87d, Rod89a, SGG89, SS87, Yav87a, Yav87b, dVWY87]. evasive [IM89a]. event [HRA89]. event-constrained [HRA89]. events [LL88b]. evolution [BR86, dM84]. Exact [DT88, PB81a]. example [BP85, Los89a, Ng81]. Examples [Ger89, Ano87-31, CL89c]. excursion [Lou84]. exhaustive [Ake87]. Existence [Cah88, Blo87, Blo88a, PP89, TB87, Yac82]. exogenous [OvD89]. expansion [CL81, CF86, HK85]. expansions [Mei86a]. expectations [Bas89, Chr89, Die89, KW89]. expected [Dev88, Vla86]. experience [Lan89]. experimental [KLT80]. experimentally [Ber88]. Experiments [LO82, KKRS84, SPF86, Ver82, VP85, Gre87a]. Expert [GGR^{+}87, BMC85, Lir89]. Explicit [DD83, WA89, Cas85b, CSL85b, EA85, ES88a, JT89, Lip84, PO85, SE88a, Sha84, YE87b]. exploitation [AR89]. Exploiting [BZ89, ZWS85]. Exponential [CLOT86, CLM87, Sal84a, Sal84b]. exposure [Ano89n, Rob87a, Rob87b]. Extended [WJMM83, AS88, BB87, Cas83a, Li89, Tho80, UA85]. Extending [HP89b]. extension [Akr81, Bra86, BEW86a, BEW86b, Ger89, WCB85]. Extensions [Chr89, CL89b, Zen83]. exterior [Ede80, MR89b]. external [SF88]. extinction [Li88]. extra [dlS87]. extrapolation [SBB83]. eye [Low89]. eyes [Ács89].

F [Eld87, Sha82]. Faber [Ell86]. factor [MB80, Osb86]. factorial [Sal84a]. factorization [NT85]. factors [KSTK84, Los89b, TSS88]. fallible [Han86]. Families [Sar88]. family [Lou83]. Fast [BP86, BHO82, CH89, Pan82b, Pan85a, PR86a, Pan89, PR89, SLPC87, TSY88, Bur82, Jul88, SLHC89, Fer82, MS88f]. Fastback [Con87b]. Fasteners [MDR85]. Fault [KK87, MHM87, SA87]. faults [MHM87, XS87]. feasibility [Mar88]. Fedorov [CK88a]. feedback [JM89, LL88a]. Feigenbaum [Ste87]. FEM [FS80, Soc88]. fermat [Mor89, Mor86]. few [Wit83b]. field [BP86, FS80, HWL89, KTW89, MS88b, SRP87, WCB85]. fields [Mar86a]. Fighter [IM89a]. figures [Lev82]. file [RPL87]. films [Fom86]. filter [Ben85]. Filtering [KA89b, BS80b, BC88]. filtration [BM85c]. Finding [Lam88, MT85, Dev81a, Pet87, Pur83b, WFW81]. finds [Bér89a]. Finite [AC86, Ano87u, KKK88, LBC82, Man87, STM87a, Ano87c, Asz80, BM84a, BJ87, BB84, BE85, BG88, CN89, Dal85, DBH88, DD83, DGL88, FLL82, FGP83, GNP82, Hay80, HWL89, KTW83, LB86, Lay85, Lay86, LL88c, LM86a, Man81, McD85, MR88a, MRV85, Mur89, NP81, OK82, Pan85a, PO85, PR85, Pou89, Ric81, St.85, Ste88, TR88, Tho80, TG88, Usm85, Vic85, Wac85, WO87, ZA88]. Finite-difference [AC86, STM87a, LM86a, McD85, PO85, St.85]. finite-dimensional [ZA88]. finite-element [Lay86, Vic85]. finitely [BL81]. fir [ByMS86]. First [Kam87, BCG87, BM85c, BSPSM81, DGL88, EC83, HHHK88, Lay85, MD86, PR85, Sza89]. first-order [BCG87, PR85]. fish [Bla83]. Fisher [Cho85]. fishery [TK89a, TK89b]. Fixed [AM81, BE85, Wer84]. Fleiss [Gre87a]. flexible [KT89]. Floating [Gin80]. Floating-Point [Gin80]. Flow [MS88b, AG80, Ber89b, BGG^{+}87, Bro87, CF88, CS88b, Gre87b, GNP82, HL86, JM84, Kha86, KTW89, Man87, Neu80, SM80, Sha81a, SRP87, TT86, WLR87]. flows [CW89, FS85, GS80, HKW89, Kup87, LBC82, LL88c, Ros80, SF88, STP88, STM88, Wan87, Wan88, WW88, dPLBA87]. fluctuations [BR86]. fluid [Ano88-43, JM84, Kup86, Kup87, SE81, TT86]. fluids [GS80]. Foerster [Wit83c]. follow [Cho84]. follow-up [Cho84]. force [Mar86a]. forced [MR88b]. forcing [MH89]. forecast [Ste88]. forecasting [Ber89b, SM88b, YNA89]. Foreign [Ste82]. Forest [Ste81c]. Foreword [Ano86f, Ano86g, Ano87v, Ano88-29, Reg87, Rod86, Rod87a, Rod87b, SL85a, Wit83a, Wit85a]. forgotten [Akr81]. form [WFW81, Yee86]. Formal [ML83, PB81a, PB81b, Sha81b]. formalism [Lie86]. formalisms [Isr83]. Formation [Blo88b, Mei86b, Blo86, MN83, PPR^{+}89]. forms [FHS84, STY83, SGN87]. formula [CF86, ES88b, SE88b]. formulae [AE87, Cas83a, Cas84, Cas86]. formulas [Fat85, SDS81]. Formulation [SP89, Ber82, BS80a, Dem82, FC83, Mor81a, YK87]. formulations [Bas89, BCP87, FS80, Gip85]. forward [CA86, CdC85]. foundation [Yak89]. foundations [Ano89o, RM87, Ros89, SDD88]. Four [Wai82, PG87, STY83]. Fourier [Fer82, MS88f, ES87, Jul88, KW81, Ng81, PB86]. fourth [BB84, BG88, Kha89, SE88b, Zen89]. fractals [Kap86b]. fraction [CL81, CF86, CdC85]. fractional [BEW86a, BEW86b]. fractionation [Rig87]. fracture [Gra86]. Free [MR88b, BE85, BM84b, Fab89, Kha86, Mor89, NR86, Pur83b]. free-boundary [Kha86]. freedom [PB81a, PB81b]. frequently [HWL89]. friction [MDR85]. frontier [Sen89b]. fullerenes [Kro89]. Fully [BDK86, Bal86b, Bal88, DS81]. Function [CKB83, TSY88, BR86, Ber82, Boz81, CN89, Dev81c, DS85, LB86, LTW83, Mar86c, MC84, ND88, Pic89a, Rob86, TM87, Wu88, YdV89]. Functional [AR85, AR87a, Cah88, Gyö88b]. Functions [AJMT81, CLOT86, MTP82, BS88, BM84b, Ber88, Che84a, Che85, CMN88, CNM88, Coh81, Dal85, EC83, FMWW86, GS81, JM89, KT86, LY88, MN88, MSH80, Mei86a, MR88c, OB88, PB82, Pie84, PB86, Rei87, Ubh84, Wac85, vNvN89]. Fundamental [Mye87, SU88, Mol88]. fundamentals [CB89]. Further [BEW86a, Wit83c]. furthest [Orl85, Tou83]. Fuzzy [BPXY88, McA88, BM84b, LL88b, LL89, ML88, WLL88, Zad83].

G [Cox86, Eld87, Wac85, Pou88]. G/M/1 [Pou88]. gain [ÅH86]. Galerkin [BDK86, CK89, DS81, Eva88, Kar86, Lay83, Ste88, vNvN89]. game [BCP87, Jär89, LT87, LO89, SD87, SGG89, SS87, Tan89, Yav87b, YdV89, YV89, dVWY87]. Games [BS89, AB87b, Ami89b, CN87, CL89a, Gai86, GS87a, GZ89, GP89, GEG87, HRA89, MP89, PM87, Raj89, RLM^{+}87, SS89, Yav84, Yav87a]. gapon [Lie86]. gas [Kov80, MPL86, Tur88]. Gauss [Ger86, GS87b, LTW83, Rab86]. Gaussian [CMN88, JJS83, Kit89]. GCD [BB87]. General [BEW86b, MM88, Ros87, Ado87a, Ado88a, AMR89, BC86, Bru87, Fer82, LB86, Sal84a, WSR87]. general-factorial [Sal84a]. Generalised [Mac86c, Fra80]. generalization [MD86, Sal84b]. Generalized [BM88, CLOT86, Kop88b, LP81, PS84, Sza87a, Wac84a, Zam88, CM88, CLM87, Cox86, HWYL88, KTW83, LJP84, Mak86a, Wal85, WD86]. generally [Pou89]. generated [Ber89b, Exo88, SBCC86]. Generating [Dev80, Sow83, Lin89, Lou83, Wu88]. generation [AR87a, BS88, Dev81c, Dev82, Han88, KSTK84, McD83, SB88]. generators [Boz81]. genetics [Rig87]. Geometric [BS85, PM87]. geometrical [Cum86, Lan89, TT89, Sha82]. geometries [Pet89]. Geometry [Sen86, Ali87, Ata85, Cam89, Dev88, Fom89, HC86, Kap86b, Rud88]. GF [HWL89]. Give [Fel85b]. given [Bro87, Dev81c]. Givens [HGP86]. Glassman [Fer82]. Global [BC86, Cus85, DLMP89, KLS89, Sha86c, Ber88, Blo87, Blo88a, GZ89, SB88, Sin83, TB87]. Globally [Ami89a, Eld83]. Go [Yak89]. Go-Moku [Yak89]. golden [Sta89]. Golem [Mil82]. Goursat [Ira87]. governed [Med88a]. graceful [HH88]. gradient [Hag89, SvR88, ZN88]. gradients [Fle88]. grammar [MK86]. Graph [Esp88, Her88, SDD88, Cox86, Exo88, HS81, Lou83]. graphic [Sch86a]. graphical [HTY87]. graphics [Ano88s, Con87a, Esp88, GG87, Har86c, Mye87]. graphs [BE88, Dev88, GH88, Har88a, HH88, HHW88, Hos86, KSTK84, McA88, Par88, STY87, Sow83]. Grashof [STM87a]. Great [Eng88b]. greatest [Pur83b]. grid [LM80]. grids [Die84, Kan88]. Group [ES88a, SE88a, WD86, AL86, Asm87, Che89, Fic86, Kop88a, SC88, YE87b]. group-theoretical [Che89]. groupoids [Bro89]. groups [Bok88, CK88a, CK88b, Mol88, Ric88, Scu89, WD86]. Grove [Ste81c]. growth [Ami89a, Bra83, Cus85, Cus86, Gri83, Mar85a, Mur83, Web86a]. GRP [LdT85]. guaranteed [Ber88]. guaranteeing [LL88a]. guidance [Mer87, SGG89]. guided [Fel85a]. guidelines [Ano85d, Ano86b, Ano87d, Ano87e, Ano88g, Ano89c]. Gulf [CARG89, LBC82]. gyrotropic [GK88].

H [Col82]. H. [Cro86, Van87]. half [Pac87]. Hall [Col82]. Halley [Pic89a]. Hambidge [McW89]. handbook [Ste87]. Hankel [PB82]. hard [Lev82]. harder [AB81b]. hardware [AB87a, Ste81c]. Harmonic [Ran88, NP81]. Haruhiko [Gra87]. harvesting [Bra83]. Harvill [Ste82]. Hash [CKB83]. hate [Wil86]. having [AB87a, FGP83, IJ80, Pan89]. healthy [Ano89n, Rob87a, Rob87b]. heart [MS88e]. heat [BdSM88, BD85, CSD88, CK89, CO83, MB89, MR88d, Mur88, Mur89, STM87b]. Helgason [Ste81a]. helicopter [IR89]. Helmholtz [BGT85]. Hensel [Rao84]. heterogeneity [AK89b]. heterogeneous [DV87, Pou87]. heuristic [Eva81b]. hexadecahedrane [Kou90]. hexcadecahedrane [Kou89]. hidden [Fom89]. hierarchy [BS89, MP88]. High [Lay86, VP85, CL81, DGL88, EW86, EW88, Gav89, Ser80, STM87a, Yan89]. High-accuracy [Lay86]. high-level [CL81, EW86]. high-order [Ser80]. Higher [Bal86b, DOS88, Ber89b, Bro89, Mei86b, Mor81a, PMS86]. higher-dimensional [Bro89]. Higher-order [Bal86b, PMS86]. histograms [BGG^{+}87]. history [Len89, SvDHH89]. hitting [YJ81]. holds [ZA88]. Holland [Ste81d]. holographic [HF86]. holography [HF86]. Home [Ste81c]. homosexual [Kra87]. honeycomb [Tar89]. Hopf [Swa88]. horizontal [CW89]. horizontally [SJ85]. Horst [Sch86a]. hovering [IR89]. HTdR [FWPT87]. HTLV [Kra87]. hull [Dev81a]. hulls [Dev82]. human [Ano89r, MS88c]. Hungarians [Bér89a]. hunter [BCP87]. Huygens [AFM88]. huyperbolic [Mar85a]. hybrid [Fat85, FLL82, Pra82, SA87, Ste88, VS88]. hybridization [Mak86b]. hybridized [HS87, HS88]. Hyperbolic [Ano84c, BP85, Dun86, McC85, Wit83b, AE87, Bal86b, Bal88, Bel83, BM85b, Blo85, BN86, DD83, Dey85, Dey88, DS81, DGL88, ES88a, Fit87, HD83, Ira87, KL83, Kar86, Kar87, Kos88, LP85, Lay83, Lay85, Mar86a, McC86, Ras85, Vic85, WW89b, Wit83c]. hypercube [HHW88, HL88, MR89a, YWLB87]. hypercubes [Bha82, GH88]. hypergeometric [LTW83]. hypotheses [KW89].

I/O [CK87, XS87]. Ia [Fle89a]. Ib [Fle89b]. IBM [Con87a]. ice [Wan87, WW88]. icosahedral [Pau89]. icospiral [Kro89]. ideal [Kup86]. ideas [McW89, Vai88]. identically [Dev80]. Identification [Los89a, CG89, Møl86, MH89, Sza88]. identify [Sza89]. identity [Ric81]. IEEE [Kam87]. II [Cha89, DHT86, GE87, MW89, PB81b, Ros80, Sza87a, TK89b]. III [Kra87]. III/LAV [Kra87]. Illinois [Lev82]. IMACS [Ano88-33]. Image [HF86, Kam87, UM85, O’C87b]. images [Mac86a]. imagination [Cum86]. Imaging [Kam87, Ali87]. imbedded [ODS86, Tho87]. imbedding [Den86b]. immigration [Sho80]. immunologic [WAYU88]. Immunology [Gro87]. imperfections [Bar86]. implementation [Ede80, Jul88, LJP84, WW89a]. implementing [JA87]. implicatii [Nic87]. implications [McD83, Nic87]. Implicit [Bru87, MR88c, Ano87c, Cas84, CS88a, Dey85, Dey88, Eva81a, Kov80, LE81, Pan82c, St.85, Yee86]. implying [Blo87]. importance [Rod88a]. impossible [Mac86a]. imprecise [Met80]. improved [LJP84]. improvement [ZL80]. improvements [BS85, Dun80]. incircle [OP87]. includes [Kri85]. including [CBC89, MK85]. Incomplete [DP88, Die84, Har86d, LY88, Mar86c]. incompressible [CF88, GNP82]. increasing [OP89]. independent [Dev80, Lou83, Neu80]. Index [Ano85g, Ano86h, Ano85b, Ano85c, Ano85l, Ano86i, Ano86j, Ano87y, Ano87-30, Ano87-34, Ano88-30, Ano88-31, Ano89b, Ano89q, Tro80]. indexed [Dal85, Wac85]. India [Ger89]. indirect [Sen88a]. induced [CLF86, CW89, SM88a, Tót86, Wan87, Wan88, WW88]. inductive [Big88]. inelastic [IJ80]. inequalities [And86, BM81, Net80, Pan85a]. inexact [Los89a]. infectious [SCH87]. infectivity [Kra87]. inference [Ano89n, Big88, Rob87a]. inferencing [LK83]. infinite [MB89]. infinitum [Zvi89]. influence [Blo86, FWPT87, KTPB85, Wil82]. Influences [McW89]. Information [GT88, Oda88, DP88, Ran87, TC89, Tuz88]. infusion [WLR87]. Inherited [Rig87, Web87]. inhomogeneous [FGP83]. Initial [Ano87w, Fit83, Cas83a, CK89, DOS88, ES87, ES88b, Far87, Fat82, Fat86, FBG88, Ger80, KKRS84, KS85, Mil86, VS88, Zem85, vN87, vN88]. initial-value [Far87, FBG88, KKRS84, Mil86, Zem85]. innovation [BM88]. inorganic [MK86]. input [OvD89, TH80]. inputs [OvD89]. insects [San86]. insensitive [SDS81]. inspired [Sha81b]. installation [MS81]. instantaneous [MB89]. Institute [Tar88]. Instructive [Ver82]. instruments [Wal87]. insulated [HKW89]. insulators [LdT85]. intefence [Rob87b]. Integer [Ric81, Kar81, Pur83a, TD88]. integers [Lam88, Pur83b, Zen89]. Integral [GZ89, Ger86, Ado83, Cah88, Cuy82, DS85, Ger82, GS87b, HK85, JJS83, JM82, KKS86, Kos88, MB80, Med88a, MS88a, MR88d, PR86b, Pas82, STB88, SP89, WFW81]. integrals [CLM87, LY88, Rab86, Rok85]. integrated [But85, TK89a, TK89b]. integration [AR87b, Cas83a, Cas84, Cas86, SZ88, UA85]. integrator [ÅH86]. integro [Bru87]. integro-differential [Bru87]. intelligence [GB85, (Co85, MC85, RLM^{+}87, Ste87]. intense [Blo86]. interacting [Ven87]. interaction [Blo86, Blo88a, IJ80]. interactions [Asl88, GS80, Prü83, WJMM83]. interactive [VMO89]. interactively [FL85]. interarrival [Gei81]. interconnection [KK87]. interconnections [Eso88]. interface [EG85, MRV85, STM87b]. interfaces [Mea86a]. interfacial [MDR85]. Interior [OK82, Man87]. internal [SF88]. International [Ano88-34, Kam87, Ste82, Ano88-33]. interplay [Sla86]. Interpolation [Fol87, LB86, Sha87, BdSM88, Ber88, Che84b, Dun80, Fra82, MM88, Sal84a, Sal84b, Sch86c, WA89]. interpolatory [PR86b, TW80]. interpolatory-newton [TW80]. Interpretation [Pau89, Big88, Kam87]. Interpreting [Isr83]. interrupted [YV81, Yav87a]. Interscience [Gra87, Ste81a]. intersection [McA88]. interval [Pet87]. intervals [Lay85]. intrapoint [Mal89]. Introduction [Ano87x, NN88, Win85, O’C87a, Ral87, Wac85]. Invariance [Roc88]. Invariant [Ame85, Den86b, Kup87]. inventories [HC89]. inventory [MS81, Oda88]. inverse [AMR89, Fra80, Kob89a, Lip84, Mor81a, MR88d, Mur88, Mur89]. inverses [KTW83]. investigation [GE87]. investigations [MWBM89]. inviscid [CF88]. Irreducible [Wim83]. irregular [EG85]. ISBN [Lev82]. ISBS [Ste81c]. Islamic [Cho89]. ISMIII’82 [Kam87]. isomerism [Sla86]. isomorphism [Kar81]. isoparametric [Tom85]. isothermal [JM84]. isotropic [LO89, YK87]. isovector [Ede80]. issue [Ano85a, Ano86a, Ano87a, Ano87b, Ano88a, Ano88r, Ano88-37, Ano89a]. issues [MC85, RC83]. iterated [BB83, Cas86]. iteration [Lip84, MN88, Rob86, TW80, Wac84b]. Iterations [Gar80, PS84, Pet87, Tro80]. Iterative [CF87a, Wer86, CP87, CF87b, CF87c, Neu80, Pas82, YE86, ZL80].

J [Cro86, Lev82, Lov82, Mea86b, O’C87b, Wal87]. Jacobi [CF87b, GHN89, Lip84, MP86, Smi83]. Jacque [Gra87]. Javanese [Haa89]. Jay [McW89]. Jaynes [Los92]. Jean [Gra87]. Jean-Jacque [Gra87]. Jeff [Ste81a]. Jersey [Col82]. Jitterbug [Ver89]. job [Che87a]. joint [Gra86, TA89]. JOR [CF87b]. Joseph [Gre87a]. joy [Har89a].

k-nearest [EC86]. Kalinowski [Mea86b]. Kane [AMR89]. Kaufman [Ste87]. Kekulé [CG86]. Kennington [Ste81a]. kernel [MM86b]. key [HHL89, KL84, Web88]. Kiguradze [KS85]. kind [SGG89]. Kinetics [Rig87, BGG^{+}87, FWPT87, Len89]. Kleene [BM84b]. Knowledge [SB85, Hav83, McC83, Tuz88]. known [EC83]. Kolmogorov [Let87]. Korteweg [BDK86]. Kronrod [Rab86]. Kutta [Bru87, DP87, Kar87, Ver82, Cas84, Cas85a, Cas85b, DP86, DLMP89, Gip85, JT89, ODS86, SE88b, Sha84, Sha87, SPF86, Tho87].

L [Gre87a, Lov82, Mea86b, Ort86, Ste81a]. laboratory [HP87, Wal87]. laceflowers [CBC89]. Ladany [Ste81d]. Lagrangian [IE88, Soc88, TM87]. Lakshmikantham [Web86b]. Language [CKB83, BMC85, CL81, DW83, Had85, LK83, McC83, McD83, Ral87, Sow83, WJMM83]. languages [GB85, TT86, Zad83]. Laplace [DD89, Duf83, PB86]. large [Chr89, EL88, He87, Mar88, SB88, Sny89, Sza87a, ZB88, ZN88]. large-scale [He87, Mar88, SB88, ZB88]. Larger [WH82]. largest [OP87]. Latin [AHL83]. lattice [Ami89b, dW88]. lattice-theoretic [Ami89b]. launch [GS89]. LAV [Kra87]. law [SS88, SGG89, SU88]. Lawes [Wal87]. laws [HHHK88, Mer87, Sha86b, Sza87a]. layer [TR88]. layers [Cas85a]. Layman [Sal84b]. leak [Tur88]. learning [LRG89, Yak89]. least [BM88, CF84b, Die84, KT89, Los89b, O’N89, PR86a, Ubh87, ZN88]. least-squares [Die84, O’N89, PR86a, ZN88]. leaving [Han87]. leaving-one-out [Han87]. Lebesgue [GB88]. Lecture [Gro87, Mea86b, Web86b]. left [MS88e]. leg [Fat84, Fat85]. lei [Nic87]. Leningrad [Har88b]. Letter [Rod88b]. Letters [Rod87c]. leukemia [SMH87, SvDHH89]. level [CL81, EW86, EW88, Wac84b]. Levin [Gro87]. Lewis [Col82]. Lexicon [CKB83]. Lidstone [AW89]. Lie [Rei81]. lifting [IR89, RI89]. Lightning [Con87c]. like [GB88, MR88c, Wer84]. Limits [Har86a]. line [Mar89, SD87, Ung82]. line-of-sight [SD87]. Linear [AG87, AB87b, OP87, PR89, Ubh83, Abe84, AW86, Ake87, BZ89, Cas86, CN87, CMN88, CNM88, Chu88a, Chu84, CF87a, CF87c, Cod88, CW86, Die89, Gar82, GP85, GEG87, HGP86, HWL89, KT89, Lam88, LL88a, LE81, Los89a, Mar88, MT85, Mil86, MC84, Pan85c, Pan85a, PR86a, Rob86, Sha82, SP89, Squ81, Swa88, TC89, TD88, Wai81, Wai82, Wil87, WFW81, YE86, YE87b, Zad89, Zad92, ZWS85, ZN88, ZA88, vN87]. linearisation [Chr89]. Linearized [Yee86]. lines [Blo87, Blo88b, Gyö88a, LO86]. linguistic [Big88, WBZB88]. linked [Omo82]. Liouville [Kob89a, Kob89b]. liquids [Fin89, Kha86, Roc88]. List [Ano85h, Ano86i, Ano86j, Ano87y, Ano88-30, Ano88-31, Ano89p, Ano89q]. Listing [Ano87z]. literature [PT86]. living [Pet89]. LMM [SZ89]. Lobatto [Ger86]. Local [Fin89, Gro88, Sin83, Fra82, RPL87, RL88]. local-area [RL88]. localization [HLHG89]. Locally [CF88, SF88]. Locating [FL85, XS87]. location [Che84a, KC80, SLHC89]. Logarithmic [CLOT86]. logic [AR85, CP87, SFK87, Lov82]. Logistic [Web86a, AL86, Gyö89, LM86a]. long [Lay85]. longest [OP89]. look [CL81]. looking [Ali87]. loop [Jär87, RPL87]. looped [EW88]. loosely [Gro88]. loosely-coupled [Gro88]. loss [JM89, MP89]. lost [Ger89]. Lotockiı [Sal84b]. love [Wil86, STB88]. Low [CF87b, DGL88]. Lower [JáJ85, Kra87]. LSI [XS87]. LSI/VLSI [XS87]. lubrication [WO87]. Luke [Ort86]. lung [SDB89]. lymphoblastic [SvDHH89]. Lymphocyte [Lum87].

M [Pou88, Pou88, Ros87]. M. [Mac86a].

M/M/1/N [Pou88]. Machine [Rob86, Yak89]. machines [BS80b]. Machol [Ste81d]. macroeconomic [Har86d]. macromolecules [Vai86]. MAGE [YE87b]. Magic [Sha86a, Loe89]. magnetic [KTW89]. magnetized [STM87c]. Magneto [CSD88, MB89]. Magneto-thermo-elastic [CSD88]. magneto-thermo-viscoelastic [MB89]. magnetohydrodynamic [KESC89, KTW89]. magnetostatics [Pas82]. mail [Ste82]. making [MM86a]. Malcolme [Wal87]. Malcolme-Lawes [Wal87]. male [Kra87]. malignant [SvDHH89]. man [WAYU88]. management [TK89a, TK89b]. manipulators [AMR89]. mapping [Han88, STP88]. maps [Pic89a]. Marcel [Mea86b, Web86b]. markets [Fab89]. Markov [Chu88b]. Markovian [Pou87]. Martin [Sha82]. Mass [Lov82, FS85, Fle88, MLDR87]. mass-conserving [FS85]. Matching [Hos86, Avi81]. material [KO88]. materials [KO88, YK87]. Math [BM90, Kou90, Wit87a]. Mathematical [Ali87, Ano88r, Ano88-34, Eng88a, GE87, MR88a, Wit87b, Ame88, BM85a, BR86, Bir85, Bla86, Joh86, Lum87, Mar83, MWBM89, MLDR87, MR88b, MS88b, MS88c, MS88d, O’C87b, Tuz88]. Mathematics [Eso88, Fus87, Mea86b, Pic89b, Web86b, Zad92, Eld87, HP86, Kap86a]. Mathl [Rob87b]. matrices [BM84a, EO82, Fra80, GHN89, Kar81, WD86]. Matrix [CP88, Fra89, Lev82, TSY88, AS88, AS89, CF86, CS89, GHN89, KTW83, Lip84, Mao86, MSH80, Pan81a, Pan81b, Pan82a, Pan82b, Pan82c, Pan85c, Pug89, Tak88, TSS88, WCB85, WSR87]. max [Med88a, Med88b, Rus89]. maximal [Lou83]. maximize [YJ81]. maximum [Che87a, Dev80, Yan87, Yan89]. Maxwell [Rod88a]. McKendrick [Cus83, Cus85, Cus86, Lev83]. MDX [SCSS85]. mean [Bro87, Oza81, Sza89]. means [Cho84, Sid83]. measure [GB88, LL88b]. measurement [Sta89]. measurements [Gup86, HLHG89]. measures [Kra87]. mechanical [LdT85]. Mechanics [Ste82, Ano88-43, Ano88-44, MS88e, SE81]. Mechanisms [Gre88a, Hav83]. media [CW89, GK88, Kha86, MS85, Ros80, Soc88, Wan87, Wan88]. Medical [Kam87]. medicine [Wit87b]. medium [KTPB85, MPW87]. medium-range [MPW87]. megakaryocyte [TSJ89]. melting [Wan87, WW88]. memoriam [Wit85b]. memory [dM84]. Merlin [Loe89]. mesh [BE85, Gro88, Han88, Kan88]. mesoscale [LL88c]. messages [KL84]. metaphors [Ben89]. method [AV82, AAM86, Aok89, BJ87, BB84, BdSM88, BG88, CLF86, CLM87, DGL88, Dun80, ES87, EA85, Eva88, Fat84, FLL82, FS80, GZ89, GS87b, Gyö88a, Hag89, Han88, HK85, Hay80, HHHK88, IR89, KKRS84, KKK88, LCG86, LO86, Luk81, LTW83, Mar85b, Mar88, MS88a, MS88f, MR88d, Mur89, MU82, NP81, Neu80, Omo82, OS84, OP86, Pan83, Pas82, Pra82, RZ86, SF88, Sha88, Sha81b, SLHC89, Sni88, SvR88, Sny89, Sod85, SPF86, SL88, TR88, Tho80, TSS88, Ver86, Wai81, Wai82, Zed87, vNvN89]. methodology [WSR87]. methods [AC86, AG87, AM81, Ano87c, BM81, BDK86, Bru87, BT89, Cas85a, Cas85b, Cer89b, Che89, CF87a, CF87b, CF87c, CF84b, CS89, DS81, Ell86, Eva81a, ES88a, Ewi80, Fat82, FG80, Ger86, Ger80, GNP82, HTY87, Hau88, JT89, JM82, Kar86, Kar87, KW81, Lay86, LP81, LJP84, LE81, Lip84, Man81, McD85, MLDR87, MdlS86, MR88c, MM86b, OK82, Pan81b, Pan85a, PMS86, PM84, Pra80, RT81, Ros80, Ros87, SE88a, SC88, Sey87, Sha84, Sha86c, Sha87, Sha82, STM87a, Tho87, Usm85, Ver82, VS88, WW89a, Wer84, Wu88, Yan87, Yan89, YE87a, ZA88, vN87, vN88]. MHD [STP88]. Michael [Lov82]. Micro [RW87, RR88a, Gri83, RR88b]. Micro-computer [RW87, RR88a]. micro-computers [RR88b]. micro-organisms [Gri83]. Microcomputer [GG87, Fus87, HP87, MRV85, Mye87, Wil87]. Microcomputers [AJMT81, MTP82, Wal87]. micropolar [AG80, GS80, JM84]. microprocessors [SSS87]. mid [LB86]. mid-edge [LB86]. migration [Lum87]. mild [Cas85a]. Milton [Joh86]. min [Med88a, Med88b, Rus89]. min-max [Med88a, Med88b, Rus89]. Mind [Low89]. mineral [SS88]. Minimal [CKB83, Mol88, SFK87, Tel89]. minimax [Che85]. minimization [Hag89, Pra80, SvR88, Sny89]. Minimizing [Che87a]. minimum [MT85, Tan89]. Minnigerode [Bra86]. misclassification [Vla86]. mismanagement [Tho89]. Mispacking [GH88]. missiles [IM89a]. missing [PG87]. Mixed [JM84, BB84, CS88a, ES87, GNP82, Krz86, LK85, Man81, Med88a, Med88b, TD88, Vla86]. mixture [DBH88, HF87]. Mizil [Nic87]. Mobile [MP86]. modal [SB88]. Model [EGZK87, AR89, Ami89a, BBPW85, ByMS86, BM89b, BM90, CB89, Duf83, FC83, Fit87, Gri83, GE87, HC86, Kra87, LBC82, LL88c, MWBM89, MLDR87, MR88b, MS88d, MN83, Mór81b, Rig87, SLPC87, SM88b, Ste88, TC89, TK89a, TK89b, TO87, TG88, WLR87, dPLBA87]. Modeling [TSJ89, Aok89, BKL88, Gre87b, Gre88b, KA89a, Koh89, Roh83, TR88, Wit83c, Wit87b]. Modelling [Ano88-34, EGZK87, FWPT87, GE87, Rob87b, Ali87, CS88b, DV87, Fle88, MR88a, ST88, Ste81b, Tuz88, WAYU88]. Models [Lev83, AA85, Ame88, Bas89, Chr89, Die89, Eld83, Har86d, Jul88, KA89b, Li88, Lum87, Mac83, Mit89a, O’N89, RR88b, Rus89, Swa88, Web86a, Zad89, Zad92, ZB88]. Modern [Hil89, Eng88a, Vai88]. modification [Sha88]. modifications [CF87b]. modified [Cas83a, GS81, LL88c, PB86, YE87b]. Modular [BB83, Pan82a]. modularity [Loe86a]. Modules [Mea86a]. modulo [MM84]. Moirés [Gig86, Wit86]. Moku [Yak89]. molecular [HC86, Sla86]. molecular-structure [Sla86]. molecules [Mak86b]. mollification [Mur87, Mur89, MH89]. MOLP [VMO89]. moments [PB86, Sza89]. Monitoring [Gin80]. Monographs [Eld87]. monotone [JáJ85, Ubh83]. monotonic [FBS88]. Monte [CA86]. Moorish [GGS86]. morphology [O’C87b]. mortality [Ano89n, Rob87a, Rob87b]. MOS [JA87]. mosaic [Cro86]. mosaics [PG87]. Moscow [Har88b]. most [Mar85b]. mother [Web89]. motion [HLHG89, LK85, Pac87, Ver89]. mouse [FWPT87]. movements [CARG89, Har89b]. moving [Kan88, PAC85, Roh83, YJ81, Yav87b]. moving-bed [PAC85]. MRAC [dlS87]. multi [LM80, MS81, OvD89]. multi-grid [LM80]. multi-input [OvD89]. multi-installation [MS81]. multi-output [OvD89]. multicommodity [Eva81b]. Multicriterion [GP89]. multidimensional [CK88a, DD89, Tel89, Yee86]. multilevel [VO85]. multimodality [ATB82]. multinomial [MM84]. Multiple [AL86, CW89, Gar80, Krz86, Sch85, Wan87, WW88, CF84a, FL85, Gup86, Lip84, MHM87, Ng81, PMS86, YK87, Zem85]. multiple-choice [CF84a]. multiplication [CP88, Pan81a, Pan82b, Pan82c, Pan85c, Pug89, Tak88]. multiplications [Pan81b]. Multirate [dlS85]. multiscale [Kri85]. multistep [Fat82, Fat84, JM82]. multivariable [STY83, SGN87]. Multivariate [Mit89a, HTY87, O’C87a, Riv86, STY87, WAYU88]. music [Apa89, Don86, Wil86]. Muslim [Róz86]. mutualists [Smi89]. my [Los92]. mycin [SCSS85, SCSS85]. myelocytic [SMH87].

N [Ste81b, Pou88]. Nash [ERH89, Gai86]. Nash-equilibrium [Gai86]. national [Koh89]. Natural [CKB83, Gra86, Had85, LK83, McC83, McD83, Pie85, WJMM83, Zad83]. nature [Sen89c]. Naval [Ste82]. Navier [Ado86, BJ87, CN89, Yee86]. navigation [YV89]. Navy [BKL88]. NBS [Ano85i]. near [PB81b, STM87b, Ung82]. near-resonance [PB81b]. nearest [Dun80, EC86]. need [Lum87]. neighbor [EC86, Orl85, Tou83]. neighbour [Dun80]. neighbourhood [MR89b]. net [Ste81b]. nets [Bal89]. network [Eva81b, HWYL88, IE88, Isr83, Lau80, ML88, Mór81b, RPL87, Ste81a, YWLB87]. networks [BM89a, Chu88b, Gra86, KK87, Par88, RL88, SB88, SDD88, WAYU88]. Neumann [Bra86]. neural [Fit87]. neurophysiological [MWBM89]. neurosurgery [LV86]. neutral [Chu88a, Gyö89, SM80]. newton [TW80, Wer84]. Newton-like [Wer84]. Niculescu [Nic87]. Niculescu-Mizil [Nic87]. Niggli [dW88]. No [Zad92, GS89]. no-escape [GS89]. nodal [Lau80]. Node [HH88]. Node-graceful [HH88]. Noise [Mur88, Møl86]. Non [BD85, CMN88, Fic86, Kit89, Pet89, vN87, Cas86, LE81, Ras85, Soc88, Swa88]. Non-Euclidean [Pet89]. Non-Gaussian [CMN88, Kit89]. Non-linear [vN87, Cas86, LE81, Swa88]. non-oscillatory [Soc88]. Non-space-group [Fic86]. Non-stationary [BD85]. non-strictly [Ras85]. nonclassical [Ami89a]. nonconforming [CN89]. nonconservative [Mar85a]. nonconvex [TM87]. noncooperative [LT87]. nonequilibrium [CO83]. Nonlinear [Bel83, Bra83, BS80b, Fat82, FN86, HD83, KSW85, Rei81, Web86b, AR89, Ado88c, AAM86, Bal86b, Bal88, BR86, Blo85, Blo86, Blo87, Blo88a, Blo88b, BT89, Cas83b, Cas84, Cas85a, CB88, Cuy82, CvdC83, Dey85, Eld87, ES88b, Ewi80, GS87c, Gyö88a, Had81, HHHK88, KKRS84, KKS86, KL83, KESC89, Mar83, Mar85a, Mar85b, Muk81, Neu80, OP86, Pra82, Qin84, Ran88, Ras85, SM80, Sen88a, Sin83, Sni88, Sod87, Was84, Web83, Web88, WW89b, Wer86, Mea86b]. nonlinearity [MK85]. Nonlinearly [Ven87]. Nonlocal [SE81]. nonserial [EW86, EW88, EL88]. Nonsingular [JM82]. nonstandard [BJ87]. Nonstationary [CG89]. nonuniform [MLDR87]. Normal [Rei87, AK89b, BHM88, Coh81]. normality [BHM88]. Normalized [LE81]. North [Ste81d]. North-Holland [Ste81d]. Norway [SMH87]. notated [Har86c]. notation [Har86c]. Note [Van87, AR84, Coh81, KO88, KTPB85, RT81, Ung82, Zam89]. Notes [Den86a, Gro87, Mea86b, Web86b]. nother [Ács89]. notice [Ano84d]. notion [Gav89]. novel [BS80a, DW83]. nozzle [HL86]. nuclear [OK88]. null [WCB85]. null-field [WCB85]. number [DT88, HS81, HWYL88, LLM84, Mor86, Mor89, STM87a, Wil82, dF89]. number-theoretic [DT88]. numbers [BS88, Hag88, LL88b, Met80, Rao84, Sza87a, WH82]. Numerical [AG80, BSPSM81, Cas86, CS88b, Fat86, GS81, Ger80, Had81, IJ80, IM89b, KKRS84, Kos88, LCG86, LO86, MS88a, MH89, OS84, PS84, Roh83, Ros80, STB88, Spa80, SPF86, WLR87, ZB88, AR87b, BM85a, BGT85, Bir85, BD85, Cas83b, Cas84, CLM87, Dey85, Dey88, DS85, Ell86, FC83, FG80, FBG88, Gei81, HHHK88, HS87, HS88, JM82, Kha86, KO88, Kov80, Lay83, LM80, LK85, Let87, Lou84, MS85, MB80, MdlS86, Mur87, Mur89, OP86, Pan85d, SE88a, SSB88, SJ85, SF88, ST88, Sha86b, SRP87, SL88, VP85, Wan87, WCB85, Yav84, ZA88]. numerical-stabilization [ZA88]. Numerically [Fel85a, But85]. Numerov [Sha88]. nutritional [TA89]. Nystrom [DP87, Bru87, Kar87].

O [Con87a, CK87, XS87]. O.D.E. [HK85]. Obituary [Ano81c]. object [Low86]. objectives [SS89]. oblique [BE85]. observable [CN87]. observation [FN86, Rot86, Yav87a]. observations [BE88, CL89b, YV81]. observed [Ben85, BM85c, FBG88]. observer [HLHG89]. occupancy [AB81a]. occurring [Qin86]. Ocean [Ano87-27, BBPW85, BS85, Bir85, CSL85a, CSL85b, DeS85, EG85, Fel85a, LK85, PO85, Pie85, Sch85]. ODE [SDS81]. ODEs [AR87b, Cas83a, Sha86b, Wu88]. O’Donnell [Lov82]. off [Oza81]. Office [Ste82]. oil [CARG89]. old [Bér89a]. One [Fat84, Fat85, Smi83, Cas86, Han87, HHHK88, MdlS86, Raj89, Sha86c, Tro80, Zam89, vN87, vN88]. one-dimension [HHHK88]. one-dimensional [MdlS86]. One-parameter [Smi83]. one-point [Tro80]. one-step [Cas86, Sha86c, vN87, vN88]. only [Pan89, Ste82]. Onogurians [Bér89a]. open [KKK88]. opening [Wit83b]. operation [Pan81a, Pan82a]. operational [OS84, SM88b]. operator [Cuy82, Fle88]. operators [FMWW86, Sal84a]. optics [GK88]. Optimal [Asz80, Ber89b, Che87b, EL88, GS89, HLHG89, HRA89, JM89, KC80, Mao86, Mil86, RPL87, Riv86, Sza88, YJ81, YV81, dVWY87, Ami89a, BM89a, BSPSM81, CL89b, Chu84, Far87, FN86, Oda88, SGG89, Ste81d]. Optimization [Ant89, CB88, GZ89, Per87, Sni88, KK86]. Optimum [Ste82]. orbits [Boy86, FBS88]. Order [Fic88, AE87, Aga88, AHL83, BB84, Bal86b, Bal88, BCG87, BM85c, Ber89b, BG88, BSPSM81, Bru87, BEW86a, BEW86b, Cas83b, DOS88, DS81, DGL88, Fat84, Fat85, GS87c, HHHK88, Kar86, Kha89, Lay85, Luk81, MD86, ODS86, PMS86, PR85, PM84, SE88b, Ser80, Sha87, St.85, VP85, Yan89, ZA88]. Order-disorder [Fic88]. ordinary [AW86, BZ89, GS87c, Iwa87, MR89a, MR88c, O’N89, PM84, Sar84, Sar88, SZ88, Tho86a, Tho86b, UA85, ZL80, ZWS85]. organisms [Gri83, Mei86b]. organization [Her86, McC83]. oriented [MWBM89]. Origin [Kar88, Gra86]. Ornamental [Mak89, Bér89a]. ornaments [GGS86]. Orr [Ger80]. orthogonal [BHO82, Mar88]. orthonormalization [Mao86]. orthoscheme [Cox89]. Oscillation [Gyö89]. oscillator [Spa80]. oscillatory [Soc88]. other [Ano87c, AB81b, GGS86, Kro89, Pan85d, SBCC86]. outline [Sat85]. Output [SBB83, OvD89]. outs [MS81]. Overall [EC83]. overdetermined [JJS83]. overhead [SFK87]. overrelaxation [GHN89]. Overview [Cho86, All83, CM88]. Oxford [Ste81b].

P [Ste81d]. P.O [Ste81c]. packet [Lau80, Mór81b]. packing [BM89a, GH88, WBZB88]. padé [AS88, AS89, Cuy82, CvdC83]. Padé-approximants [Cuy82, CvdC83]. paid [Ste82]. pair [Pan81a]. pairs [Orl85]. Pall [Zen83]. pandiagonal [AHL83]. Papadakis [Mea86b]. Papadimitriou [Col82]. paper [Van87]. Papers [Ano86k, Ano89r, Ano83a, Ano88r, Ano88p, Ano88q, Zel92]. parabolic [BN86, CSL85a, CS88a, Kha89, Kri85, McD85, MK85, PO85, Pie85, Sch86c, SL85b, SRP87, Ver86]. paradigm [SCSS85]. Parallel [CP88, Gar80, HGP86, Muk81, PR89, Pug89, Ano85i, BP86, CF87a, EW84, GM88, Gro88, Jul88, Lin89, Lip84, Lot87, MR89a, MP86, NT85, Pan85d, Pan87b, Pan89, SHPT87, WW89a]. Parameter [DBH88, AB87b, AR87b, CL89a, CS89, EC86, Gri83, Smi83, Ver86, dlS87]. Parametric [MM86b, Sni88]. parasite [HD83]. parity [MHM87]. Part [Ano86k, Ano89r]. Partial [Ado83, Mea86b, Rei81, Sat85, Ado87a, Ado88a, Ano84c, Ano88r, Bel83, CL81, CF86, DD83, Dey85, Ewi80, FN86, Gyö88a, HD83, Kha89, LP85, Lam88, LO86, MN88, MS88f, Neu80, NR86, OS84, OP86, PR85, Rok85, St.85, Tho86a, Wit83b, Wit83c, Zem85]. Partially [CN87, Ben85, BM85c]. Particle [Gre89]. particles [Man86]. Pascal [BS88, Ral87]. Pascal-SC [Ral87]. passivity [Sen88a]. Password [HHL89]. patch [MN83]. path [Pan81a]. pathfinder [SDD88]. paths [Ami89a]. Pattern [Was86, Big88, Cum86, Gra86, HP89b, PPR^{+}89]. patterns [Cro86, Haa89, Mak89, Mam86, Sch86b, SBCC86, Tar89]. PDE [Mea86a, FC83, MR88a]. PDE/PROTRAN [MR88a]. PDEs [McC86]. peak [dPLBA87]. Pearson [CNM88]. Penalty [FLL82, Ber82, OK82, Pra80, Pra82]. Penalty-hybrid [FLL82]. Penrose [Loe89]. pentangle [Loe89]. perceptual [LN89]. perennial [Sla86]. Perfect [CKB83]. perfection [Har86a]. perfectly [Sch86b]. performance [LdT85, McL86, Ran87]. period [Ano89n, Cho84, Rob87a, Rob87b, dPLBA87]. period-application [Ano89n]. Periodic [Cus86, KL83, LP85, BM85a, BP85, BC88, Cus83, DHT86, EO82, GS86, Jen86, MS88b, Ran88, Sey87, VS88, Yac82]. peripheral [SvDHH89]. Persistence [Li88]. person [LT87, SS89]. personal [Con87a]. perturbation [Bro87, Li89]. perturbations [BCG87, RZ86]. perturbed [Lot87]. Petersen [Cox86]. petroleum [CW85]. phase [Che89, KA89a, SC88, TT89]. phases [Kar88]. phenomena [SBCC86]. phenomenon [Sla86]. philosophy [Den86a]. phyllotaxis [Dix89]. Physical [STM88]. physics [Ado88c, BR86, Gav89, Kop88b, Man86, OK88, PPR^{+}89]. physiological [WLR87]. picture [Asm87, Doc86]. pictures [Ha89, Wal85]. Piecewise [PR86b, CW86, FBS88, Ger82, HL86]. piecewise-steady [HL86]. Pierre [Shu88b]. pilot [SvDHH89]. pins [XS87]. pipeline [CW85]. pipelines [Tur88]. pivot [Pan85b]. pivotal [Zla82]. pivoting [Tho86a]. placement [AB81a, Ran87]. plan [Alb88]. planar [SGG89]. plane [BS80a, Blo86, Blo88a, EC83, HF86, Pac87, Ung82, Yav87b, YK87, Zla82]. plane-wave [Blo86]. planning [BKL88, BM89a, SI89, SB85]. plans [Sid83]. plants [dlSGA87]. plate [Fra82, JM84, UW80b]. plates [MDR85]. players [Yav87b]. Plenum [Wal87]. Point [Gin80, AAM86, Bar86, Cas83b, Cas84, Cas85a, Cas86, CL89a, CK88b, Eva88, KKRS84, STM87b, Tro80, Ubh84, Usm85]. points [EC83, GR88, Wer84]. Poiseuille [SM80]. Poisson [Mar86c, McC85]. Pol [Spa80]. policy [JM89]. pollutant [Ame88]. pollution [LL88a, ZB88]. polygon [Orl85, OP87]. polygons [ATB82, Cha89, MT85]. polyhedra [Dav86, Mol88]. polyhedral [Fle89a, Fle89b, MW89]. Polyhedron [STY87]. polymer [Chû86]. polymers [Col85]. Polynomial [BT89, PS84, BP86, Pan82a, Pan85d, Pan87a, Pan87b, Pan89, PMS86, Pet87, PP89, Sal84a]. polynomial-exponential [Sal84a]. polynomials [AW89, EW84, Ger82, HZ86, PR86b, TSS88]. polytopes [HL89]. polytropes [STM87c]. Polytypism [Zvy88]. Population [Eld83, Bla83, Bra83, Cus83, Cus85, Cus86, Eld87, Kra87, KLS89, KSW85, Li88, Mac83, Mar83, SvDHH89, Swa88, Web83, Web86a, Web89, WSR87]. populations [AK89b, CMN88, HD83, Mar85a, Prü83, Ven87, Web87]. porous [CW89, Kha86, MS85, Ros80, Soc88, Wan87, Wan88]. portfolio [Sen89b]. position [Har88c]. Positive [Fra89, Lam88, Lay86, LM86a, Zen89]. possible [SBCC86]. postage [Ste82]. potential [PB81a, PB81b]. potentials [Kob89a, Kob89b, Men86]. pottery [Cam89]. power [Che87b, CL89b, CS89, MM84, Pan83, TSS88]. powerful [SL88]. powers [BB83, Zen89]. pp [Eld87, Lev82, Lov82, Nic87, O’C87a, Ste81a, Ste81d, Ste82, Zad92]. pp. [Col82, Con87a, Gra87, Gre87a, Gro87, Kam87, Ste87, Wal87]. Practical [He87]. Practice [Web86b, Esp88]. preceding [Ver86]. Precise [Mor86]. preconditioning [Eva81a, Wac84a]. predation [Lev83]. predicting [SB88]. Prediction [BM85c, CARG89, CF88, FBG88]. predictor [Ben85, DD83, Kha89]. Preface [Ano85j, Ano87-28, Ano88-32, Ano89s, Ano89t, AK89a, Cer83, HP88, Har86b, HV88, LM87, Mit89b, Ode80, PY89]. prehistoric [Cam89]. prejudices [Los89b]. Prentice [Col82]. prepaid [Ste82]. presence [CF84b, Krz86]. presentation [HF86, Mol88]. presented [BL81]. preserve [HP89b]. preserving [STP88]. Press [Kam87, Ste81b, Ste81c, Wal87, O’C87b]. pressure [CN89]. preventative [Kra87]. Price [Lev82, Ste82]. prices [HC89]. pricing [Cer89b]. primal [TM87]. primal-dual [TM87]. prime [AHL83, Hag88, MM84]. Principal [TSY88, Los89b, Rab86]. principal-value [Rab86]. principle [AFM88, Ben89, Bra86, Hua89, Roc88, Yan87, Yan89]. principles [Tho80]. priority [Asz80]. probabilistic [Hag88, TO87]. Probability [Ara83, BR86, LL88b, Vla86, YJ81, YdV89]. problem [AB86, AF87, AB81a, Avi81, BE85, BS80a, BKL88, Blo86, Blo88a, CL89b, Cho85, CK81, Dav86, Dem82, Eva81b, FBG88, Ger80, Ira87, Kha86, Kra82, MPL86, Med88a, Med88b, MM88, MN83, MR88d, Mur88, Mur89, Neu80, OK82, Pan81a, Pan83, Ras85, Roh83, Tou83, UM85, Was84, Zem85, Zen89, Ric81]. problems [AC86, Aga88, AW89, AAM86, Ano87-27, Ant89, Ata85, BGT85, Bir85, Blo85, BN86, BC88, Cas83a, Cas83b, Cas84, Cas85a, Cas86, Che84a, Che85, CB88, DeS85, DD89, DOS88, ES87, Eng88a, Eva88, ES88b, Far87, Fat82, Fat86, Fit83, FG80, FGP83, FS80, GP89, GNP82, HK85, KKRS84, KL83, Kar87, Kob89a, Kob89b, KS85, LP81, LJP84, Lin86, LE81, Lip84, LO86, Man87, Mar88, McC85, Mil86, Mor81a, MU82, NP81, NS88, Pan85d, Pas82, PR87, Ran88, Rud88, Sch85, STM87a, Sny89, SL88, Tak88, TT86, TD88, UW80a, UW80b, Usm85, Ver86, VS88, Was84, ZN88, vN87, vN88]. procedure [Gei81, MS85, VMO89]. procedures [But85, NR86, Sza88]. Proceedings [Kam87, Ste82]. process [Jur83, Sza89, TSJ89, TK89a, TK89b]. processes [Ben85, Den86b, McC85, Oda88, Sar88, Sho80]. processing [BS80b, EW88, GT88, Kit83, ML83, Pou89, WW89a]. processors [Gro88, Lip84, Sch85]. products [Pan82a]. Professional [Ste82]. Professor [Ort86]. programmable [SFK87, Lev82]. Programming [GB85, Ric81, Bel83, Boz81, CM88, CB88, CF84a, EW86, EW88, EL88, Lov82, NS88, OP87, PR86a, Pra82, Sni88, TD88, Ste81a]. Programs [Ede80]. progression [SMH87]. Project [ML88]. projection [Mar85b, Wai81, Wai82]. projections [Mar88]. projective [Pan83]. projectors [Den86b]. Proof [SB88, And86, Kar81]. proofs [BHM88]. Propagation [Gin80, Vic85, Ano87-27, Ara83, BGT85, EG85, KKK88, KTPB85, LP81, LJP84, SJ85]. properties [Ara83, Bha82, CG86, FBS88, Men86, Web87, YWLB87]. Proportional [YV89]. proportionally [Ser80]. proportionally-damped [Ser80]. proportions [Cho84]. protocols [AWYU82]. PROTRAN [MR88a]. proving [LO82]. pseudorandom [Møl86]. pseudosymmetry [Fin89]. public [HHL89, KL84, Web88]. public-key [HHL89, KL84]. Publisher [Ano84d]. Publishers [Lev82, Lov82]. Publishing [Mea86a]. Pueblo [Cam89]. pulse [MU82]. Pure [Mea86b, Web86b, Eld87, WW88]. pursuer [YV89]. Pursuit [Rod89a, Yav87a, AR87c, CN87, GS87a, GZ89, GEG87, Jär87, Mer87, MP89, MPW87, Pac87, PR87, Rod87d, SGG89, SS87, Yav87b, dVWY87].

Qin [Van87]. qua [Par88]. quadratic [AB87b, CN87, Ger82, Wer86, Zam89]. quadrature [GR88, JJS83, Squ81]. quadtrees [Abe84, Gar82, Wal85]. Qualitative [Ira87, Prü83, SS89]. quality [Ant89, Pou89]. quantifiers [Zad83]. quasi [Ubh84, Ubh87]. quasi-convex [Ubh84, Ubh87]. quasicrystals [Pau89]. quasilinear [Gip87]. Quasilinearization [Rot86, And86]. Quasimolecular [Gre87b, Gre88b]. quaternion [WD86]. query [Had85]. questions [Ste81c]. queue [Gei81]. queueing [Pou87]. queues [Asz80, LL89]. queuing [Chu88b, Pou88]. Quintic [UW80a, WA89]. quotient [EO82]. quotient-difference [EO82].

R [Col82, Mea86b, Ste81b, Wac85]. Rabbit [BCP87]. radar [Ali87]. radial [Che84a]. Ramsey [Exo88]. Random [BdSM88, BS88, Cah88, CK89, Coh81, Dev80, Dev81c, Dev82, KSW85, MS88a, Rei87, STB88, Sod85]. Range [Ano87-29, HC89, Jär89, MPW87]. Ranger [Ung82]. ranging [BS85]. rank [CF87b, KTW83]. rapid [Dun80]. rat [SMH87]. rate [AWYU82, CMN88, CNM88, Chu88b, HF87, Han87, Rig87]. rates [DHT86, Kno86, O’N86]. Rational [Ber88, DS85, vN88, Bas89, Chr89, Dal85, Die89, KLT80, Pic89a, Rao84, SPF86, Wac85, vNvN89]. ray [Hau88, SDB89]. reaction [BM89b, BM90, FS83, MLDR87, Sod85, Sod87]. reaction-diffusion [Sod85, Sod87]. reactions [Bal86a, Pea86]. reactor [PAC85]. real [Mar89, Pan89]. realistic [Jär89, SB85]. realization [AF87]. reasoning [RC83]. recall [McC83]. received [Ano81a, Ano82a]. reciprocal [Dem82]. Recognition [Hav83, Asm87, Big88, Sid83]. reconnaissance [SvR88]. Reconstruction [Ger89, Mur88]. records [AB81a]. recovery [Riv86]. rectangles [Hay80, Sha86a]. recurrences [Wim83]. recurring [Cum86]. Recursive [BM84a, AS89, SLPC87, Sza88, YNA89, dM84]. reduce [Dev81a]. Reducing [MM84]. reduction [NS88]. Reference [Ano87-30, Pie85]. refinement [Gro88, Kan88]. reflection [Vic85]. Reflectors [WLL88]. refracting [EG85]. regimes [SI89]. regional [SM88b]. regions [Abe84, CW86, Gar82]. regression [KT89]. regular [Cha89, CK88b, PG87]. Regularity [Har86c]. regularization [MD86]. regulation [EGZK87, GE87]. Reidel [Sha82]. Reimannian [Rud88]. related [Ant89, BCG87, MWBM89, Omo82, Pan85c, Rig87]. Relation [CK88a]. relations [BPXY88, SCH87]. relationship [GHN89, Len89]. relative [RT81]. relativistic [Kup87]. relaxation [CSD88, IE88, MB89]. relevance [SBCC86]. Reliability [YWLB87, Ano88-44, RL88]. Remarks [Str89, Tho86b, Wit87b]. Remes [KLT80]. Remes-differential [KLT80]. renaissance [Cum86]. renewable [AR89]. renewal [Sin83]. repertoire [Lum87]. repetition [MB80, Was86]. Reply [Los92]. report [Ano89v, Ano89w]. Reports [Rod85a, Ano86c, Ano87f, Ano87g, Ano87h, Ano87i, Ano88h, Ano88i, Ano88j, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89x, Ano89y, Ano89z, Rod84a, Rod84b, Rod84c, Rod85b]. representation [Asm87, BM88, Eng88b, Kup86, Tuz88, Wim83]. representational [RC83]. representations [CK88a, Fel85a, Hav83, Kop88a, SC88]. Representative [Ano87-31]. represented [Abe84, Gar82]. representing [HTY87]. requirements [Sla86]. Resampling [CNM88, CMN88]. Resampling-type [CNM88]. research [Ano86l]. reserve [CW85]. residual [SM88a]. residue [Bos87, LLM84]. resolution [Let87, LO82]. resolution-based [LO82]. resonance [PB81a, PB81b]. resource [Koh89, RR88a, RR88b]. resources [AR89, EL88]. response [TH80]. restricted [JáJ85, Kit83]. restrictions [Med88b, SU88]. results [Dem82, EGZK87, Exo88, Gyö89, Han86, KLT80, Sen88a, Sza87a, TB87]. retransmitted [Lau80]. retrials [Pou87]. Review [Col82, Con87a, Eld87, Gra87, Gre87a, Gro87, Kam87, Lev82, Lov82, Mea86a, Mea86b, Nic87, O’C87b, O’C87a, Sha82, Ste87, Ste81a, Ste81b, Ste81c, Ste81d, Ste82, Wal87, Web86b, Mea86a]. revised [Pan85b]. revisited [CK81, Sen89a]. ribosomal [Rig87]. Rich [Ste81c]. Richard [Ste81a, Ano86l, KK86, Pan86]. Riemann [Lin86, MPL86, Ras85]. rigor [Lan89]. rigorous [Hua89]. ring [Ric81]. rival [Rus89]. river [LL88a]. Robert [Lov82, Ste81d, Rod88a]. robot [SLHC89, Gra87]. robust [Kno86, Sen88a, dlS87]. rocket [LO89]. Roger [Lev82]. role [Haa89, Hau88, Wil86]. Root [Fra89, Gar80, PS84, TSY88, Bur82, CH89]. Roots [Gar80, Bur82, EW84, Pan85d]. Rosenbrock [GS87c, Ver82, Zed87]. Rosenbrock-type [GS87c, Zed87]. Rotating [KTW89, AG80, OB88, STM87c]. rotation [BKL88]. rotations [Zla82]. rotelet [Ung82]. rough [BBPW85]. round [Oza81]. round-off [Oza81]. Rounding [CdC85]. row [AF87]. Ruin [Cam89]. rule [Gup86, Lam86, McL86, UA85]. rules [Big88, O’N86, Pea86, Squ81, SBCC86]. Runge [Bru87, Cas84, Cas85a, Cas85b, DP86, DP87, DLMP89, Gip85, JT89, Kar87, ODS86, SE88b, Sha84, Sha87, SPF86, Tho87, Ver82].

S [Gro87, Mea86b, Van87]. saddle [CL89a]. saddle-point [CL89a]. Saint [Hua89]. Saint-Venant [Hua89]. same [Rus89]. samples [CA86]. Sampling [Alb88]. Sarafyan [Tho87]. satellite [Mer87]. saturated [CW89, Wan87]. scalar [HHHK88]. scale [He87, Mar88, SB88, ZB88]. scaling [Tel89]. scan [CK87]. scans [LV86]. scattered [EC83, Fol87, Fra82]. Scattering [Men86, BBPW85, Den86b, WCB85]. scenarios [MPW87]. schedules [AB87a, Rig87]. schema [Hav83]. schema-based [Hav83]. scheme [Ano87c, CSL85b, CS88a, Gip87, Hag89, HL86, Kha89, KL84, Kov80, PO85, PR85, St.85, WW89b]. schemes [Ano87u, CSL85a, Far87, GS87c, Los89b, Mil86, MP86, ND88, Soc88, Yee86]. Science [Mea86a, Chû86, Ros89, Mea86b]. sciences [Tuz88, Tar88]. Scientific [Gin80, Ral87]. screening [SDB89]. sea [BKL88]. Search [CKB83, Omo82]. seasonal [Kit89, YNA89]. Second [Ano88-33, AE87, Aga88, Bal86b, Bal88, Bru87, CO83, DS81, EC83, Fat84, Fat85, GS87c, Kar86, Luk81, SDS81, Ser80]. second-order [AE87, Bal86b, Bru87, DS81, Fat85, GS87c, Kar86]. Section [Ano89-30, Ano85k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87-32, Ano87-33, Ano88w, Ano88x, Ano88y, Ano88z, Ano88-27, Ano88-28, Ano88-38, Ano88-39, Ano88-40, Ano88-41, Ano88-42, Ano89l, Ano89m, Ano89-27, Ano89-28, Ano89-29, Ano89-31, Ano89-32, Rod89b, Sta89]. Sector [TSY88]. Seen [Doc86]. segment [DV87]. segmented [Pan85c]. seismic [MU82]. selection [Gei81]. Selective [CK87]. self [Eva81a, JA87]. self-checking [JA87]. Semantic [Kit83, ML83]. semi [Lip84]. semi-direct [Lip84]. semiconductor [PPR^{+}89]. semigroup [Web83]. semigroups [BL81, FN86]. Semilinear [Ado87b, AFM88]. sensorineural [Møl86]. separability [BZ89]. September [Ano88e]. sequence [Lam86, OP89]. sequences [Smi83]. sequencing [Che87b, CL89b]. Sequential [MR89a, Pan87b, Cho84, Cho85, Kra82, Pan85d, TD88]. series [Aok89, CG89, CW85, HC89, KW81, Mit89a, Ng81]. Serra [O’C87b]. server [Gei81]. servers [Pou87]. set [BPXY88, EC83, Kou89, Kou90, LT87, Ver89]. set-valued [BPXY88]. sets [Lou83]. Seventh [Ano88-34, Bér89a]. Several [Hua89, DGL88, HTY87]. severe [Fle88]. Shadow [Had85, Tan89]. shallow [ND88]. shallow-water [ND88]. Shaul [Ste81d]. shear [Ber82]. shell [Kro89]. shells [SP89, Tar89]. shift [CS89, Ver86]. Ship [MP89]. Shock [Kan88, Blo88b]. shore [BKL88]. short [Jär89, TH80]. short-range [Jär89]. shortest [Pan81a]. showing [Pau89]. shrunken [Han87]. Shubnikov [SU88, Tar88, Vai88]. si [Nic87]. side [Ali87]. side-looking [Ali87]. sided [PM84]. sight [SD87]. sign [LLM84]. signature [SHPT87]. Signorini [Dem82, OK82]. Similarity [HKW89, Her88, Shu88c, WW88]. Simple [GEG87, Mac83, Pac87, AR89, EW84, Fer82, Fin89, FHS84, FBS88, Orl85, Sch86c, SBCC86]. Simple-motion [Pac87]. simplest [Mar89]. simplex [Pan85b]. Simplification [Ng81]. Simplified [BM85b, SSS87]. simulating [DV87]. Simulation [Ran87, ZA88, Gre88a, Gre89, Koh89, SMH87, SHPT87, TT86]. simulations [Spa80, WLR87]. simultaneous [PMS86]. simultaneously [SB88]. single [Bal86b, Bal88, Gei81, Kar86, MR88c, MHM87, Pic89b]. single-step [Bal86b, Kar86]. singular [BCG87, DS85, FMWW86, Fat86, Ger82, Ger86, GS87b, JJS83, Mao86, PR86b, PB86]. Singularities [McC86, CF84b, YK87]. singularity [Blo86]. Sippl [Lev82]. sister [Web89]. situations [SA87]. sixth [ODS86]. sixth-order [ODS86]. size [AWYU82, Ano87-29, Dev88, DHT86, Wal85, WSR87]. sizes [Han86]. skeleton [OP87]. slabs [BD85]. slicks [CARG89]. SLOR [YE87a]. slot [HKW89]. Slotine [Gra87]. small [CA86, Pie84]. Smedsaas [Mea86a]. Smooth [Fra82, UW80b, SZ89]. smoothing [EC83, EC86, KA89b]. snowflakes [CBC89]. so-called [Pau89]. soap [Fom86]. Sobolev [AG87, Aga88]. social [Nic87]. sociale [Nic87]. Society [Kam87]. Soft [Ste81c]. Softstrip [Ano88-37, Ano88-35, Ano88-36, Ano89u]. Software [Ano85k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87-32, Ano87-33, Ano88w, Ano88x, Ano88y, Ano88z, Ano88-27, Ano88-28, Ano88-38, Ano88-39, Ano88-40, Ano88-41, Ano88-42, Ano89l, Ano89m, Ano89v, Ano89x, Ano89y, Ano89z, Ano89w, Ano89-27, Ano89-28, Ano89-29, Ano89-30, Ano89-31, Ano89-32, Rod89b, Mea86a, Ros87, Ste81c]. solid [CSD88, MB89, OK88]. solids [Bar86, CO83]. solitary [IJ80]. solute [MLDR87]. Solution [Ado86, BN86, Che84a, Che85, Luk81, LTW83, PR89, Ado87a, Ado88a, AW86, AAM86, Ano87c, AG80, BGT85, BS80a, BZ89, BD85, Cas83b, CLF86, Chu88b, CF87a, CF87c, Cod88, Cuy82, CvdC83, DD89, DOS88, Dey88, Die89, DS85, ERH89, ES88b, FC83, Ger82, HS87, HS88, IJ80, KKS86, Kha86, Lam88, LCG86, LE81, LO86, MR89a, MK85, MS88a, MLDR87, MR88c, MR88d, Mur89, OS84, OP86, PR86b, Rok85, SE88a, STB88, Sar84, Sar88, SJ85, Sha86b, SZ89, Tan89, Wer86, Wu88, Yav84, YE86, ZL80, ZWS85]. Solutions [BdSM88, Ame85, Blo87, Cah88, Cus83, Cus85, Gyö88a, HKW89, Ira87, KL83, KESC89, Kos88, LP85, LM86a, Lin86, McD85, PB81a, PB81b, Sey87, Sin83, UW80a, UW80b, Wan87, WW88, WFW81, WO87, Yac82, Zem85]. solve [TD88]. solvents [TSS88]. solver [Dey85, Lot87, Tho86b]. solvers [Tho86a]. Solving [Ado88c, HWL89, JJS83, Lip84, ZN88, AE87, BM81, Eva88, FS80, HGP86, Kov80, LP81, Mar85b, Mar88, Muk81, MU82, PO85, Pur83a, SL88, Wai81, Wai82, YE87b, Zed87]. Soma [BCC^{+}86]. Some [Ata85, BHM88, Bas89, Bir85, CL89b, CL89c, DeS85, Exo88, Mar83, PMS86, Pet87, Qin86, RC83, ST88, TB87, Usm85, Was84, WCB85, Zen83, Cas86, Cod88, Dem82, Dev81b, Dev88, DS81, Gyö88a, Pan85d, Tho87, Ver82, ZL80]. Sommerfeld [Ger80, VP85]. SOR [GHN89]. sorting [Gre89, Sha81b]. sound [CO83]. source [Asz80, STM87b, Sla86]. sources [CSD88, MB89]. South [Ger89]. Southampton [Ano88e]. Soviet [Tar88]. Space [Zhe88, Zou89, Aok89, Bok88, Cer89b, CK88b, CB89, DGL88, Fic86, Ker89, KA89b, Kop88a, Low86, Mal89, Mit89a, OvD89, SC88, Str89, TB87, TW80, eZ86]. space-time [Mal89, eZ86]. spaces [CN89, GP85, WLL88]. span [HL88]. spanning [HWYL88]. sparse [Tho86a, Zla82, ZN88]. sparsity [ZWS85]. Spatial [ByMS86]. Spatio [PPR^{+}89]. Spatio-temporal [PPR^{+}89]. speaker [Sid83]. Special [Ste82, Ano88r, Duf83, Fra80, Mei86a, PM84, Tak88]. specific [Prü83]. specification [AR85, Kar88]. specimen [MR89b]. spectral [Fel85a, Ste88, TSS88]. spectroscopy [Roc88]. speed [MP89, YdV89, YV89]. speed-loss [MP89]. speeds [PR87]. sphere [KC80]. spherical [STP88, Tar89]. Spiral [Dix89]. spirals [LRBK89]. spline [AAM86, BT89, Che84b, Die84, HS87, HS88, MR88c, UW80a, UW80b, WA89]. spline-based [AAM86]. spline-like [MR88c]. splines [Fra82]. Split [Wac84b, SB88]. Split-level [Wac84b]. spontaneously [Man86]. sports [Ste81d]. spread [DV87]. Springer [Gro87]. spurious [Vic85]. Square [Fra89, Gar80, LRBK89, Bur82, Sza89]. Square-Root [Gar80]. squares [AHL83, BM88, CF84b, Die84, KT89, Los89b, MC84, O’N89, PR86a, Ubh87, ZN88]. Stability [CW86, FS83, Gri83, PS84, Sha84, Tho87, BN81, BC86, DHT86, GS87b, He87, JT89, JM82, Mar86a, MB80, Sey87, Sin83, Swa88, TB87]. stabilization [JM89, ZA88]. Stabilizing [AV82]. Stable [Lay83, TSY88, CSL85b, Fel85b, Lam86, Rab86, RZ86, Sha88, SLPC87, UA85, Zed87, dlS87, dlSGA87]. stage [Cho84]. Stagnation [GS80]. star [BHO82]. Starkweather [Cam89]. starlike [HL88]. State [Cer89b, OvD89, Aok89, BC88, CW89, Chu88b, CB89, Eld83, KA89b, Med88b, Mit89a, OK88, Str89]. State-space [OvD89]. station [Pou89]. stationary [AG80, BJ87, BD85, KTW89, Man87]. statistical [But85, GG87, O’C87a, Yak89]. statistics [BPXY88, SS88]. Steady [Chu88b, BC88, CW89, CF88, Eld83, HL86, STM88]. stellation [Fle89a, Fle89b, MW89]. step [AM81, Ano87-29, Aok89, Bal86b, Bal88, BSPSM81, Cas86, Kar86, MR88c, Pan85b, Sha86c, VS88, vN87, vN88]. steplengths [Lam86]. stepping [Ewi80]. steps [SvR88]. stepsize [SE88b]. stepwise [CA86]. stereochemistry [Sok86]. stereographic [Eng88b]. stereotactic [LV86]. Sternberg [Mea86b]. stiff [AR87b, Cas83a, Far87, FS80, Lam86, Tho86b, Zed87]. stiffness [BM84a].

Stiintifica [Nic87]. stimulation [EGZK87]. stimuli [Sla86]. Stochastic [Mer87, RR88b, SL88, Sza87b, AB87b, BR86, BdSM88, BCP87, Cah88, CN87, Gai86, OvD89, Pan86, SSB88, Sza87a, YV81, Yav84, Yav87b, dVWY87]. stochasticity [Wit83c]. stock [Ber89b, MS81]. stock-outs [MS81]. Stokes [Ado86, BJ87, CN89, Yee86]. stopped [LT87]. stopping [FN86]. storm [CLF86]. Strange [SBCC86, Sha81b]. strapdown [Mao86]. strategic [CW85]. strategies [BK87, HRA89, Ran87, Ste81d, Zla82, dVWY87]. strategy [AE87]. stratified [SJ85]. Stream [OB88, CN89]. street [Gre88b]. stress [YK87]. Stresses [MR89b]. strictly [Ras85]. strip [Ano88-37, Ano88-35, Ano88-36, Ano89u]. strong [GS80, Sha81a]. strong-connectivity [Sha81a]. structered [Cus86]. Structural [Mal89, Ste82, Ano88-43, Che89, Her88]. Structure [STM87b, BM85a, BC86, MK86, MRV85, RTF^{+}88, STM87c, Sla86, TK89a]. structured [Bla83, Cus85, Li88, Mur83, Web86a, Web87]. structures [Bal86a, CG86, Fic88, GS86, Mei86b, PPR^{+}89, SB85, Tar89, Zvy88]. stuck [XS87]. stuck-at [XS87]. studies [Ano89n, Rob87a, Rob87b, RI89]. Study [BGG^{+}87, MR88b, MS88e, BN81, CA86, IM89b, Jär89, Kap86b, MS88f, SvDHH89, STM88, SRP87, TA89, TK89b]. studying [Che89]. Sturm [Kob89a, Kob89b]. subcommunities [Smi89]. Subcritical [SM80]. subcubes [GH88]. Subject [Ano85l, Ano87-34]. sublanguages [Kit83]. subscribers [Ano84d]. subsequence [OP89]. Successive [MS88a, GHN89, MP88, Mar88]. Sufficient [Yan89, Qin86]. Sukhanov [KKRS84]. Summation [Met80]. sums [Ake87]. supercomputer [Gre88a, Gre88b, Gre89]. supercomputers [Ano88f, Ano88e, Ano88d]. superlinearly [Rus89]. supraconvergent [WW89b]. sure [CK89]. surface [CARG89, HKW89, IR89, MLDR87, PM87, RI89, Ste82]. surface-tension [HKW89]. surges [CLF86]. surveillance [LO89]. Survey [Ano89-30, Ano85k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano87s, Ano87t, Ano87-32, Ano87-33, Ano88w, Ano88x, Ano88y, Ano88z, Ano88-27, Ano88-28, Ano88-38, Ano88-39, Ano88-40, Ano88-41, Ano88-42, Ano89l, Ano89m, Ano89-27, Ano89-28, Ano89-29, Ano89-31, Ano89-32, Cas86, Ell86, HHW88, Lir89, Rod89b]. Survival [SDB89, TO87]. survivor [Ano89n, Rob87a, Rob87b]. suspension [LdT85]. sustained [Ano89n, Rob87a, Rob87b]. switching [BM84b, Mór81b, SI89]. Symbiosis [Gli85]. symbol [Joh86]. symmetric [HL89, Kob89a, Kob89b, Lay86, Low86, McC85, Mei86b, Orl85, Sch86b, Tou83, dF89]. symmetrical [Tel89]. Symmetries [Apa89, Fom86, Gav89, Har88b, Doc86, Ger89, Mak86a, Sch86a, eZ86]. Symmetrology [Mak86a, Mak89]. Symmetry [Bal86a, Bar86, Bér89a, BHW86, Bro89, CSUZ89, Chû86, Don86, Fab89, Fle89a, Fle89b, Gra86, GGS86, Ha89, Har86c, Har89b, Har88c, HP86, Ker89, Loe86b, Loe86a, Mak86b, Man86, Mar89, MW89, MM86a, Osb86, OK88, PPR^{+}89, Pea86, RTF^{+}88, Ros89, Róz86, RR89, SC88, Sen89a, Sen89c, Shu88c, Tót86, Urm86, Vai86, Wil86, Yu89, Ano86k, Ano89r, Asl88, Ben89, Bok88, Cer89a, CK88a, Cho89, CG86, Dun86, Fic86, Fom89, GS86, Haa89, Har89a, Hau88, Her86, Hil89, HP89b, Hos86, Jen86, Kop88b, Kou89, Kou90, LN89, Loe89, Mac86a, Mac86b, Mal89, MWBM89, Pac89, Pau89, PT86, San86, Sen86, Sen88b, Sha86a, SS88, Shu88b, Sla86, Was86, Zhe88, Zou89]. Symmetry-breaking [PPR^{+}89]. Symmetry-making [MM86a]. symmetry-oriented [MWBM89].

Symposium

[Ano88-43, Kam87, Ste82, Ano88-33].

synovial [TA89]. System

[HC89, LV86, All83, Asz80, BM85b, BM85c, Ber89b, Fit87, KA89a, Kop88a, Lam88, Los89a, MS81, Mar85a, MS88a, Møl86, Pan85a, Per87, Pou88, Qin84, Ras85, Rus89, TH80, TK89a, TK89b, Urm86, Wit83c, WFW81, YE86, Gro87]. systematic [JV88, XS87]. Systems [Blo85, Møl86, PR89, dM84, AA85, Ado87a, Ado88a, AB87b, AR87b, BCG87, BP85, BR86, BC86, BZ89, BMC85, CK87, CK88b, CF87a, CF87c, Cod88, Den86b, DGL88, EW86, EW88, Eso88, EL88, GGR^{+}87, Gli85, GS87c, Har86d, He87, HWL89, JJS83, Kra82, KW89, Lam86, Lay83, Lay85, Lay86, LL88a, LLM84, LRG89, Lir89, Mar85b, MC85, MR88a, Muk81, Neu80, OvD89, Pan86, Pou87, PB81a, PB81b, Pur83a, Sar84, Sen88a, Ser80, STY83, SGN87, Sod85, Sod87, Tuz88, Wai81, Wai82, WW89b, Wer86, YV81, YNA89, ZA88, Mea86a, Sha82]. systolic [BB87]. sytem [CvdC83].

T

[Los92, Mea86a, O’C87a, SvDHH89, SA87]. T-cell [SvDHH89]. t-characterization [SA87]. T7 [RTF^{+}88]. table [Ano88-37, Jen86, Sha86a]. tableau [TC89]. tables [Ort86]. tactical [LRG89]. T’ai [Har89b]. Tamil [Ger89]. Tandem [Pou88]. tanks [FS83]. target [SD87, YJ81, Yav87b]. targets [SS87]. tasks [McC83]. tau [LTW83, OS84, OP86, LO86]. tau-elements [OP86]. Tau-lines [LO86]. teaching [Apa89, Iwa87]. Technical [Nic87]. technique [AR84, CK87, LM80, SJ85, Sch86c, VO85, XS87]. technique-1 [VO85]. Techniques [JA87, MC85, Ano84a, CF87c, Gyö88b, Omo82, Pan85d, Wai81, Wai82, WCB85]. technological [Nic87]. technology [Bér89a, JA87]. tehnice [Nic87]. tehnologice [Nic87]. telecommunications [BM89a]. temporal [PPR^{+}89]. tension [HKW89, Sch86c]. Tenth [Bér89a]. terminal [Rot86]. terms [MH89]. terrain [LL88c]. Tesselation [Pic89b]. Test [AB87a, AR87a, Ano87-31, CW85, Cho85, PP89]. testable [SFK87]. Testing [CP87, AR85, Ake87, SSS87]. tests [Cho84, KTW83, MHM87]. text [ML83]. textbooks [Eld87]. texts [Kit83]. textures [Shu88a]. th [CH89, PM84, TSY88]. their [And86, Ber82, CG86, Gyö89, Len89, PB82, Sar84, Vai86, Ver89, WCB85]. them [Lev82]. theodicy [Joh86]. theorem [Akr81, LO82, O’N89, SA87, Zen83]. theorem-proving [LO82]. Theorems [DD89, Cah88, Wim83]. theoretic [Ami89b, DT88, HC89, Kou89, Kou90, Mei86a, SDD88]. Theoretical [Mak89, Che89, Per87, ST88]. Theory [Bok88, CM88, Esp88, Gro87, VP85, Web86b, Asm87, Bla86, Blo85, BEW86a, BEW86b, Chu88a, Col82, Cuy82, Exo88, HHW88, KO88, Mal89, Mar86a, RI89, SC88, Sha82, SI89, Sta89, UW80b, Wil82, Eld87]. therapy [Rig87]. thermal [CSD88, MB89]. thermo [CSD88, MB89]. thermodynamics [Len89]. thin [Fra82]. third [St.85]. third-order [St.85]. thoughts [Bas89]. Three [FS85, Wai81, AR87c, BBPW85, DV87, Gip85, HF86, KSTK84, Los89a, Net80, SL85b, St.85, Yav84, Yav87b]. Three-dimensional [FS85, AR87c, BBPW85, DV87, HF86, SL85b]. three-variable [Los89a]. thrombopoiesis [EGZK87, GE87]. tibia [MS88c]. tidal [CLF86, Duf83, LBC82, LCG86]. Tidally [SM88a]. Tight [Tro80]. Tikhonov [MD86]. Tilings [Cha89]. Time [BR86, CW85, DHT86, KT89, MK85, SvDHH89, Aok89, Ber89b, BN86, CG89, Che87a, Chu88a, Chu88b, Cus83, Ewi80, Gei81, HC89, Kar87, Ker89, Lay85, LM80, Mal89, Mar86a, MT85, Men86, Mit89a, NP81, SLPC87, SGN87, Ubh83, Zhe88, eZ86]. time-dependent [Chu88b, Kar87, Men86]. Time-domain [MK85]. time-harmonic [NP81]. Time-history [SvDHH89]. Time-periodic [DHT86]. time-stepping [Ewi80]. Time-varying [KT89, SGN87]. times [ML88, Pou89, Ver86]. tip [IR89]. tissue [DV87]. tool [Iwa87]. Topological [CBC89, RL88, Bai89, WLL88]. topology [RPL87]. total [Che87b]. total-work-content-power [Che87b]. totally [JA87]. towed [LK85]. Tower [WBZB88, Cho89]. traffic [SB88, dPLBA87]. trajectory [Eng88b]. transcription [Rig87]. transfer [CF86, MLDR87, ND88]. transform [DD89, Jul88, Mor86, Mor89, MU82, UM85, Fer82, MS88f]. transformals [Kar81]. transformation [SZ88, TT89]. Transformations [SGN87, FBS88, HGP86, STY83, Smi83]. transformers [Ver89]. transforms [DT88, PB82, PB86]. transient [FS80]. transitions [Che89, SC88]. transmission [AWYU82, Blo87, Blo88b, Fit87]. transonic [HL86, Neu80]. transport [Ame88, CO83, TA89]. trapezoidal [ES88b, UA85]. treatment [BM85a, Fat86, LK85, Per87, SSB88, ZB88]. trees [Chu84, GM88, HL88, HWYL88, MHM87, Ric88]. trends [Ano88-43, Web86b]. triangular [Che84b, LB86]. Tricomi [AB86]. tridiagonal [EO82, YE86, YE87b]. trigonometric [Pic89a]. Trilinear [Pan82c]. trip [SB88]. Tripartite [Eso88]. triples [DP86, DP87, DLMP89]. Trisecting [Cox89]. tsunami [KKK88]. Tucson [Ste82]. tumor [AK89b, DV87]. turbulence [SE81, TG88]. turn [MP89]. tutorial [Li89]. TVD [Yee86]. twins [Pau89]. TWK [CL89b]. TWK-power [CL89b]. Two [BCP87, Cho84, Gyö88b, Jär89, PM84, Raj89, SS89, VS88, AF87, AAM86, Aok89, AB81b, BJ87, ByMS86, Cas83b, Cas84, Cas85a, Cas86, CLF86, CP87, Eva88, GS86, Jay92, KKRS84, KW81, KSTK84, LCG86, LM86a, Lin86, Los92, McC86, MT85, ND88, PB81a, PB81b, Pur83b, SD87, SS87, Sza89, Usm85, Yav87b, YdV89, YV89, Zla82]. two-dimensional [BJ87, ByMS86, CLF86, CP87, LCG86, LM86a, Lin86, ND88]. Two-on-one [Raj89]. Two-person [SS89]. two-point [AAM86, Cas84, Cas85a, Cas86, Eva88, KKRS84, Usm85]. two-row [AF87]. Two-stage [Cho84]. Two-step [VS88, Aok89]. two-target [SD87, Yav87b]. type [AG87, Aga88, CNM88, Chu88a, DS85, Ger82, Gip85, GS87c, HK85, KO88, LK85, Neu80, SDS81, Zed87]. type-independent [Neu80]. type-insensitive [SDS81].

U.K [Ste81b]. U.S [Ste82]. U.S.A [Ste81c]. ultimate [LL88a]. uncertain [CLS87, CL89c, LL88a]. uncertainties [GS87a]. Uncertainty [Ben89, Kra82, Roc88]. unconstrained [SvR88]. Understanding [DW83, Ano89r, LK83]. underwater [Ame85, BGT85, LP81, LJP84]. unidirectional [JV88]. unified [SA87]. Uniform [Far87, Gro88, LL88a, Mil86, YK87]. Unifying [Ano89r]. unique [MP88]. Uniqueness [Zem85, TB87]. univariate [Hag89]. universal [SFK87]. University [Ano88e, Ste82]. unknown [ÅH86]. unmodelled [dlSGA87]. unsaturated [Kha86]. unseen [Doc86]. unstable [Lay83]. updating [AR84]. upon [Bok88, Wer86]. upper [Kra87]. Urban [Cer89a]. use [Ake87, FBG88, Ger82, Kar88, Lev82, LTW83, McC83, VO85, WCB85]. used [LdT85]. using [AR87a, AE87, AMR89, BN86, Cas83a, Cas86, CLF86, Che85, DT88, Hag88, HHL89, HGP86, JJS83, KKRS84, LCG86, Mao86, MR88a, MRV85, Møl86, Mor86, Mor89, Pan82a, PB86, VMO89, WLR87, WBZB88, ZA88, vNvN89]. utilizing [RR88b].

V [Ste81a, Vai88, Web86b, GH88]. value [AC86, Aga88, AW89, AAM86, Cas83a, Cas83b, Cas84, Cas85a, Cas86, DD89, DOS88, ES87, Eva88, ES88b, Far87, Fat82, Fat86, Fit83, FBG88, Ger86, Ger80, HK85, JT89, KKRS84, KS85, LE81, Lip84, LN89, Mao86, McC85, Mil86, NP81, Rab86, STM87a, UW80a, UW80b, Usm85, VS88, Zem85, vN87, vN88]. valued [BPXY88, LT87, MSH80]. values [FMWW86]. Variable [Pra80, AR87b, BB84, BG88, Cas83b, Los89a, OS84, ODS86, Pra82, PR87, Sha87, YdV89, YV89]. variables [Coh81, Dev80, Dev81c, DGL88, McC86, Ran88, Rei87, Vla86, WBZB88]. variance [O’N86, Oza81, Tel89]. variational [BM81, MM88, Tho80]. variations [CS89, Kri85]. various [MSH80, ND88]. varying [KT89, SGN87]. vault [MR88b]. Vector [CL89a, Ben85, BM85c, BS80b, CP88, LT87, MR89a]. velocity [SRP87]. Venant [Hua89]. ventricle [MS88e]. Version [Rod87d, Rod89a]. vertebra [MS88d]. vertex [MT85]. vertical [JM84, Wan87]. vertices [KSTK84, Orl85]. vessels [MS88b]. vi [Eld87]. via [EL88, Gre88a, KT89, Lev83, Luk81, STP88]. vibration [MS88c]. vibrations [MR88b]. view [Bai89, Low86, Sen89b]. viewpoint [CL81]. VII [CNM88]. VIII [Pic89b]. Vincent [Akr81]. violated [Mar85b]. virus [Kra87]. viscoelastic [MB89]. viscoelasticity [Had81]. viscous [GNP82, Man87, SF88, STM88]. vision [Gli85]. Visual [Fom89, Den86a, Lan89, MM86a]. visualization [Don86]. Viterbi [WW89a]. VLSI [AB87a, CK87, XS87]. Vol [Gro87, Mea86b, Ste87, Web86b, Zad92, Eld87, Ste81c]. Volterra [Bru87, GP85, JM82, MB80, MS88a].

Volume [Ano88-31, Ano86i, Ano86j, Ano87y, Ano88-30, Pas82, Rok85]. Voronoi [Dav86]. vortex [Gre88b]. Vries [BDK86]. VSEPR [HC86].

W [O’C87a]. Wachspres [Wac85]. Wachspress [Dal85]. wall [SE81, Wan87]. Waring [Zen89]. water [Ant89, CW89, EL88, Kra82, ND88, RR88a, RR88b, WW88]. wave [Ado87b, Ame85, Ano87-27, AFM88, BGT85, Blo86, Blo88a, CSL85a, CSL85b, IJ80, KKK88, KTPB85, LP81, LJP84, McD85, Men86, Mor81a, PO85, SE88a, SJ85]. wave- [Blo88a]. waveguide [EG85]. wavenumber [Pie85]. wayward [Sen89c]. weak [GS80]. weather [FBG88, SM88b, Ste88]. Webb [Eld87]. Weber [CK81, MR88d]. wedge [YK87]. weevil [FC83]. weevil-cotton [FC83]. weights [ODS86]. well [Ber88, Sla86]. well-conditioned [Ber88]. which [AB81b, Big88, HL88, Kri85]. white [Sch86b]. whole [TB87]. Wide [Ano87-35]. Wide-angle [Ano87-35]. width [GM88]. Wiley [Con87a, Gra87, Gre87a, O’C87a, Ste81a]. Williams [KL84]. wind [CLF86, CARG89]. wind-driven [CARG89]. wind-induced [CLF86]. Winthrop [Lov82]. within [(Co85]. without [Pan82b]. Woods [Cro86]. work [Che87b, KK86]. worker [Ano89n, Rob87a, Rob87b]. works [Shu88b]. Workshop [Ano88-44]. world [SS88, Sok86]. Worst [Avi81, BK87].

x [Wai82, Hau88, SDB89]. x-projection [Wai82]. X-ray [Hau88, SDB89].

York [Con87a, Eld87, Gra87, Gre87a, Gro87, Mea86a, Mea86b, O’C87b, O’C87a, Ste81a, Wal87, Web86b]. young [Ha89]. Yudell [Ort86].

Zellner [Los92]. zero [HWL89, MD86, Pet87, PP89, ZA88]. zero-order [ZA88]. Zeros [PS84, FL85, Pan87a, Pan87b, Pan89, PMS86]. Zolotarev [HZ86].

References

Adomian:1985:CSA

[AA85] G. Adomian and G. E. Adomian. Cellular systems and aging models. Computers and Mathematics with Applications, 11(1–3):283– 291, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                                   tronic).  URL http://www.
                                                   sciencedirect.com/science/
                                                   article/pii/0898122185901531.

Angeles:1986:SBM

[AAM86] J. Angeles, M. A. Alvarado, and A.-R. Mansouri. A spline-based method of solution of nonlinear two-point boundary-value problems. Computers and Mathematics with Applications, 12 (Part A)(9):969–985, September 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                                   tronic).  URL http://www.
                                                   sciencedirect.com/science/
                                                   article/pii/0898122186900222.

Arato:1981:DPR

[AB81a] M. Arató and A. Benczúr. Dynamic placement of records and the classical occupancy problem. Computers and Mathematics with Applications, 7(2): 173–185, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                                   tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122181901176.

Asmuth:1981:EAC

[AB81b] C. A. Asmuth and G. R. Blakley.

An efficient algorithm for constructing a cryptosystem which is harder to break than two other cryptosystems. Computers and Mathematics with Applications, 7(6):447–450, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900298.

Adomian:1986:TP

[AB86] G. Adomian and N. Bellomo. On the Tricomi problem. Computers and Mathematics with Applications, 12(4–5): 557–563, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901811.

Abadir:1987:TSV

[AB87a] M. S. Abadir and M. A. Breuer.

Test schedules for VLSI circuits having built-in test hardware. Computers and Mathematics with Applications, 13(5– 6):519–536, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900800.

Aihara:1987:LQS

[AB87b] S.-I. Aihara and A. Bagchi. Linear–quadratic stochastic differential games for distributed parameter systems. Computers and Mathematics with Applications, 13(1–3):247–259, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901088.

Abel:1984:CDC

[Abe84] D. J. Abel. Comments on “Detection of connectivity for regions represented by linear quadtrees”. Computers and Mathematics with Applications, 10(2):167–170, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900464.

See [Gar82].

Agarwal:1986:FDM

[AC86] R. P. Agarwal and Y. M. Chow. Finite-difference methods for boundary-value problems of differential equations with deviating arguments. Computers and Mathematics with Applications, 12 (Part A)(11):1143– 1153, November 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900180.

Acs:1989:ATE

[Ács89] M. Ács. Asymmetry through the eyes of a(nother) chemist. Computers and Mathematics with Applications, 17(1–3):425–441, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901727.

Adomian:1983:PDE

[Ado83] G. Adomian. Partial differential equations with integral boundary conditions. Computers and Mathematics with Applications, 9(3):443–445, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900585.

Adomian:1986:SNS

[Ado86] G. Adomian. Solution of the Navier–Stokes equation — I. Computers and Mathematics with Applications, 12 (Part A) (11):1119–1124, November 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900167.

Adomian:1987:GAS

[Ado87a] G. Adomian. A general approach to solution of partial differential equation systems. Computers and Mathematics with Applications, 13(9–11):741–747, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901593.

See corrigenda and comment [Ado88a].

Adomian:1987:SWE

[Ado87b] G. Adomian. Semilinear wave equations. Computers and Mathematics with Applications, 14 (6):497–499, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900617.

Adomian:1988:CCG

[Ado88a] G. Adomian. Corrigenda and comment on “A general approach to solution of partial differential equation systems”. Computers and Mathematics with Applications, 15(6–8):711, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902908.

See [Ado87a].

Adomian:1988:EED

[Ado88b] G. Adomian. Elliptic equations and decomposition. Computers and Mathematics with Applications, 15(1):65–67, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900727.

Adomian:1988:SNE

[Ado88c] G. Adomian. Solving the nonlinear equations of physics. Computers and Mathematics with Applications, 16(10–11): 903–914, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902003.

Abdullah:1987:NSS

[AE87] A. R. B. Abdullah and D. J.

Evans. A new strategy for solving second-order hyperbolic equations using asymmetric formulae. Computers and Mathematics with Applications, 13(9– 11):831–838, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901660.

Al-Fedaghi:1987:TRC

[AF87] S. S. Al-Fedaghi. The tworow constraints realization problem. Computers and Mathematics with Applications, 14(3): 173–183, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901507.

Astaburuaga:1988:HPS

[AFM88] M. A. Astaburuaga, C. Fernandez, and G. Perla Menzala. Huygens’ principle and semilinear wave equations. Computers and Mathematics with Applications, 15(6–8):603–609, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902830.

Anwar:1980:NSM

[AG80] M. Anwar and G. S. Guram. Numerical solution of a micropolar flow between a rotating and a stationary disc. Computers and Mathematics with Applications, 6(2):235–245, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900310.

Agarwal:1987:LMD

[AG87] R. P. Agarwal and R. C. Gupta. Linear methods for differential equations of Sobolev type. Computers and Mathematics with Applications, 14(7):519–525, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187900472.

Agarwal:1988:BVP

[Aga88] R. P. Agarwal. Boundary value problems for second order differential equations of Sobolev type. Computers and Mathematics with Applications, 15 (2):107–118, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900806.

Aastrom:1986:DCI

[ÅH86] K. J. Åström and A. Helmersson. Dual control of an integrator with unknown gain. Computers and Mathematics with Applications, 12 (Part A)(6): 653–662, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900520.

Atkin:1983:ECP

[AHL83] A. O. L. Atkin, L. Hay, and R. G.

Larson. Enumeration and construction of pandiagonal Latin squares of prime order. Computers and Mathematics with Applications, 9(2):267–292, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218390130X.

Andrews:1981:EFM

[AJMT81] M. Andrews, D. Jaeger, S. F.

McCormick, and G. D. Taylor. Evaluation of functions on microcomputers: exp(x). Computers and Mathematics with Applications, 7(6):503–508, 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900341.

Arino:1989:P

[AK89a] O. Arino and M. Kimmel. Preface. Computers and Mathematics with Applications, 18(10– 11):xi–xii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900035.

Axelrod:1989:CHP

[AK89b] D. E. Axelrod and T. Kuczek. Clonal heterogeneity in populations of normal cells and tumor cells. Computers and Mathematics with Applications, 18(10– 11):871–881, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900059.

Akers:1987:ULS

[Ake87] S. B. Akers. The use of linear sums in exhaustive testing. Computers and Mathematics with Applications, 13(5–6):475–483, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187900770.

Akritas:1981:VFT

[Akr81] Alkiviadis G. Akritas. Vincent’s forgotten theorem, its extension and application. Computers and Mathematics with Applications, 7(4):309–317, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900602.

Albert:1986:MGL

[AL86] A. Albert and E. Lesaffre. Multiple group logistic discrimination. Computers and Mathematics with Applications, 12 (Part A) (2):209–224, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690074X.

Alborzi:1988:SP

[Alb88] M. Alborzi. Sampling plan. Computers and Mathematics with Applications, 15(2):153–157, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900843.

Ali:1987:MMS

[Ali87] A. E. Ali. Mathematical modelling of side-looking radar imaging geometry and errors. Computers and Mathematics with Applications, 14(2):91–117, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901404.

Allen:1983:ASO

[All83] James F. Allen. Argot: a system overview. Computers and Mathematics with Applications, 9(1):97–109, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218390010X.

Allen:1981:FSD

[AM81] K. Allen and S. Mckee. Fixed step discretisation methods for delay differential equations. Computers and Mathematics with Applications, 7(5): 413–423, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901280.

Ames:1985:ISU

[Ame85] W. F. Ames. Invariant solutions of the underwater acoustic wave equation. Computers and Mathematics with Applications, 11(7– 8):681–685, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901646.

Ames:1988:AMM

[Ame88]  W. F. Ames.  Analysis of

mathematical models for pollutant transport and dissipation. Computers and Mathematics with Applications, 16(10– 11):939–965, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122188902040.

Amir:1989:GOP

[Ami89a] R. Amir. Globally optimal paths in the nonclassical growth model. Computers and Mathematics with Applications, 18(6– 7):663–674, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901168.

Amir:1989:LTA

[Ami89b] R. Amir. A lattice-theoretic approach to a class of dynamic games. Computers and Mathematics with Applications, 17(8– 9):1345–1349, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901004.

Angeles:1989:AID

[AMR89] J. Angeles, O. Ma, and A. Rojas. An algorithm for the inverse dynamics of n-axis general manipulators using Kane’s equations. Computers and Mathematics with Applications, 17(12): 1545–1561, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900540.

Andrus:1986:DIT

[And86] J. F. Andrus. Differential inequalities and their application to a convergence proof of quasilinearization. Computers and Mathematics with Applications, 12 (Part A)(8):909–919, August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186900271.

Anonymous:1980:A

[Ano80a] Anonymous. Announcement. Computers and Mathematics with Applications, 6(4):449, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122180900528.

Anonymous:1980:EB

[Ano80b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 6(1):i, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812218090053X.

Anonymous:1981:BR

[Ano81a] Anonymous. Books received. Computers and Mathematics with Applications, 7(4):353, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122181900675.

Anonymous:1981:EB

[Ano81b] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 7(1):iii, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122181900018.

Anonymous:1981:O

[Ano81c] Anonymous. Obituary. Computers and Mathematics with Applications, 7(4):281–283, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900560.

Anonymous:1982:BR

[Ano82a] Anonymous. Books received.

Computers and Mathematics with Applications, 8(3):243, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122182900499.

Anonymous:1982:EB

[Ano82b] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 8(1):i, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122182900347.

Anonymous:1983:ACP

[Ano83a] Anonymous. Announcement and call for papers. Computers and Mathematics with Applications, 9(5):701, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218390127X.

Anonymous:1983:EB

[Ano83b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 9(1):iii, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122183900019.

Anonymous:1984:CTA

[Ano84a] Anonymous. Computational techniques and applications conference. Computers and Mathematics with Applications, 10 (6):iii, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900695.

Anonymous:1984:EB

[Ano84b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 10(1):ii–iii, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/0898122184900798.

Anonymous:1984:HPD

[Ano84c] Anonymous. Hyperbolic partial differential equations. Computers and Mathematics with Applications, 10(2):95, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900403.

Anonymous:1984:PNA

[Ano84d] Anonymous. Publisher’s notice to authors and subscribers. Computers and Mathematics with Applications, 10(6):477, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900786.

Anonymous:1985:AIa

[Ano85a] Anonymous. About this issue. Computers and Mathematics with Applications, 11(1–3): vii, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901312.

Anonymous:1985:AIb

[Ano85b] Anonymous. Author index.

Computers and Mathematics with Applications, 11(7–8):iii– iv, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901853.

Anonymous:1985:AIc

[Ano85c] Anonymous. Author index. Computers and Mathematics with Applications, 11(12):xvi, December 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0898122185901130.

Anonymous:1985:AG

[Ano85d] Anonymous. Author’s guidelines. Computers and Mathematics with Applications, 11(1–3): ix, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901324.

Anonymous:1985:EBa

[Ano85e] Anonymous. Editorial Board. Computers and Mathematics with Applications, 11(1–3):iii, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901300.

Anonymous:1985:EBb

[Ano85f] Anonymous. Editorial board. Computers and Mathematics with Applications, 11(5):ii, May 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812218590046X.

Anonymous:1985:I

[Ano85g] Anonymous. Index. Computers and Mathematics with Applications, 11(1–3):i, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122185901579.

Anonymous:1985:LC

[Ano85h] Anonymous. List of contents. Computers and Mathematics with Applications, 11(12): ix–xv, December 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901129.

Anonymous:1985:NPC

[Ano85i]  Anonymous.  NBS parallel

computer benchmark collection. Computers and Mathematics with Applications, 11(12): 1243, December 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901105.

Anonymous:1985:P

[Ano85j] Anonymous. Preface. Computers and Mathematics with Applications, 11(5):v, May 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900471.

Anonymous:1985:SSS

[Ano85k] Anonymous. Software survey section. Computers and Mathematics with Applications, 11 (12):i–iii, December 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901117.

Anonymous:1985:SI

[Ano85l]  Anonymous.  Subject index.

Computers and Mathematics with Applications, 11(7–8):i– ii, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901841.

Anonymous:1986:AI

[Ano86a] Anonymous. About this issue. Computers and Mathematics with Applications, 12(4– 5):vii, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901689.

Anonymous:1986:AG

[Ano86b] Anonymous. Author’s guidelines. Computers and Mathematics with Applications, 12(4– 5):ix, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901690.

Anonymous:1986:BR

[Ano86c] Anonymous. Book reports. Computers and Mathematics with Applications, 12 (Part A)(8): 945–949, August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900295.

Anonymous:1986:EB

[Ano86d] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 12 (Part B) (1–2):i, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901318.

Anonymous:1986:E

[Ano86e] Anonymous. Erratum. Computers and Mathematics with Applications, 12(3):375, March 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901963.

Anonymous:1986:Fa

[Ano86f] Anonymous. Foreword. Computers and Mathematics with Applications, 12 (Part A)(2): vii, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900696.

Anonymous:1986:Fb

[Ano86g] Anonymous. Foreword. Computers and Mathematics with Applications, 12 (Part A)(2): vii, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900696.

Anonymous:1986:I

[Ano86h] Anonymous. Index. Computers and Mathematics with Applications, 12(4–5):i, April/May 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901896.

Anonymous:1986:LCAa

[Ano86i] Anonymous. List of contents and author index volume 12a, 1986. Computers and Mathematics with Applications, 12 (Part A)(12):i–x, December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900076.

Anonymous:1986:LCAb

[Ano86j] Anonymous. List of contents and author index volume 12b, 1986. Computers and Mathematics with Applications, 12 (Part 2)(5– 6):i–v, vii, September/December 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186902555.

Anonymous:1986:PAS

[Ano86k] Anonymous. Papers to appear in symmetry (part 2). Computers and Mathematics with Applications, 12 (Part B)(1–2): ix, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690132X.

Anonymous:1986:RCR

[Ano86l] Anonymous. The research contributions of Richard Bellman. Computers and Mathematics with Applications, 12 (Part A) (6):633–651, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900519.

Anonymous:1987:AIa

[Ano87a] Anonymous. About this issue. Computers and Mathematics with Applications, 13 (9–11):vii, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218790157X.

Anonymous:1987:AIb

[Ano87b] Anonymous. About this issue. Computers and Mathematics with Applications, 14 (9–12):vii, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902185.

Anonymous:1987:ASM

[Ano87c] Anonymous. Applicable solution methods other than the implicit finite difference scheme. Computers and Mathematics with Applications, 14(5):361–368, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902136.

Anonymous:1987:AGa

[Ano87d] Anonymous. Author’s guidelines. Computers and Mathematics with Applications, 13 (9–11):ix, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901581.

Anonymous:1987:AGb

[Ano87e] Anonymous. Author’s guidelines. Computers and Mathematics with Applications, 14 (9–12):ix, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902197.

Anonymous:1987:BRa

[Ano87f] Anonymous. Book reports. Computers and Mathematics with Applications, 13(4):407–412, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187900095.

Anonymous:1987:BRb

[Ano87g] Anonymous. Book reports. Computers and Mathematics with Applications, 13(7):657–659, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901283.

Anonymous:1987:BRc

[Ano87h] Anonymous. Book reports. Computers and Mathematics with Applications, 14(3):229–232, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901556.

Anonymous:1987:BRd

[Ano87i] Anonymous. Book reports. Computers and Mathematics with Applications, 14(6):501–503, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187900629.

Anonymous:1987:EBa

[Ano87j] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 13(1–3):iii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/0898122187900873.

Anonymous:1987:EBb

[Ano87k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 14(1):i, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218790174X.

Anonymous:1987:ESSa

[Ano87l] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 13(4):i–iii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900101.

Anonymous:1987:ESSb

[Ano87m] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 13(7):i–iii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901295.

Anonymous:1987:ESSc

[Ano87n] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 13(8):i–iii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900447.

Anonymous:1987:ESSd

[Ano87o] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 13(12):i–iii, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900691.

Anonymous:1987:ESSe

[Ano87p] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 14(2):i–iii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901453.

Anonymous:1987:ESSf

[Ano87q] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 14(4):i–iii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901386.

Anonymous:1987:ESSg

[Ano87r] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 14(6):i–iii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900630.

Anonymous:1987:ESSh

[Ano87s] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 14(7):i–iii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900538.

Anonymous:1987:ESSi

[Ano87t] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 14(8):i–iii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218790191X.

Anonymous:1987:FDS

[Ano87u] Anonymous. Finite difference schemes. Computers and Mathematics with Applications, 14 (5):317–325, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187902094.

Anonymous:1987:F

[Ano87v] Anonymous. Foreword. Computers and Mathematics with Applications, 13(1–3):vii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900885.

Anonymous:1987:IBC

[Ano87w] Anonymous. Initial and boundary conditions. Computers and Mathematics with Applications, 14(5):327–340, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902100.

Anonymous:1987:I

[Ano87x] Anonymous. Introduction. Computers and Mathematics with Applications, 14(5):305–307, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902070.

Anonymous:1987:LCA

[Ano87y] Anonymous. List of contents and author index volume 13, 1987. Computers and Mathematics with Applications, 13 (12):i–ix, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900708.

Anonymous:1987:LCC

[Ano87z] Anonymous. Listing of computer codes. Computers and Mathematics with Applications, 14(5):385–419, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218790215X.

Anonymous:1987:OAW

[Ano87-27] Anonymous. Ocean acoustic wave propagation problems. Computers and Mathematics with Applications, 14 (5):309–316, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902082.

Anonymous:1987:P

[Ano87-28] Anonymous. Preface. Computers and Mathematics with Applications, 13(1–3):ix–x, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900897.

Anonymous:1987:RSS

[Ano87-29] Anonymous. Range step size analysis. Computers and Mathematics with Applications, 14 (5):341–349, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902112.

Anonymous:1987:RCI

[Ano87-30] Anonymous. Reference citation index. Computers and Mathematics with Applications, 14(5):421, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902161.

Anonymous:1987:RTE

[Ano87-31] Anonymous. Representative test examples. Computers and Mathematics with Applications, 14 (5):369–383, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902148.

Anonymous:1987:SSSa

[Ano87-32] Anonymous. Software survey section. Computers and Mathematics with Applications, 14 (1):i–v, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901830.

Anonymous:1987:SSSb

[Ano87-33] Anonymous. Software survey section. Computers and Mathematics with Applications, 14 (3):i–iii, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901568.

Anonymous:1987:SI

[Ano87-34] Anonymous. Subject index. Computers and Mathematics with Applications, 14 (5):423, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187902173.

Anonymous:1987:WAC

[Ano87-35] Anonymous. Wide-angle capability. Computers and Mathematics with Applications, 14 (5):351–360, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187902124.

Anonymous:1988:AI

[Ano88a] Anonymous. About this issue. Computers and Mathematics with Applications, 15 (6–8):vii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902672.

Anonymous:1988:Aa

[Ano88b] Anonymous. Announcement. Computers and Mathematics with Applications, 15(1):iv, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188900764.

Anonymous:1988:Ab

[Ano88c] Anonymous. Announcements.

Computers and Mathematics with Applications, 16(12):v– vii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902659.

Anonymous:1988:ASEb

[Ano88d] Anonymous. Applications of supercomputers in engineering. Computers and Mathematics with Applications, 15(3):iv, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901782.

Anonymous:1988:ASEa

[Ano88e] Anonymous. Applications of supercomputers in engineering: Southampton University, England 5–7 September 1989. Computers and Mathematics with Applications, 15(2):iv, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900880.

Anonymous:1988:AAS

[Ano88f] Anonymous. ASE 89: Applications of supercomputers in engineering. Computers and Mathematics with Applications, 15(5):v, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901654.

Anonymous:1988:AG

[Ano88g] Anonymous. Author’s guidelines. Computers and Mathematics with Applications, 15 (6–8):ix, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902684.

Anonymous:1988:BRa

[Ano88h] Anonymous. Book reports. Computers and Mathematics with Applications, 15(2):159–160, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188900855.

Anonymous:1988:BRb

[Ano88i] Anonymous. Book reports. Computers and Mathematics with Applications, 15(3):235–240, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188901757.

Anonymous:1988:BRc

[Ano88j] Anonymous. Book reports. Computers and Mathematics with Applications, 15(5):425–426, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901629.

Anonymous:1988:BRd

[Ano88k] Anonymous. Book reports. Computers and Mathematics with Applications, 15(9):795–796, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901307.

Anonymous:1988:BRe

[Ano88l] Anonymous. Book reports. Computers and Mathematics with Applications, 15(11):975–976, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900430.

Anonymous:1988:BRf

[Ano88m] Anonymous. Book reports.

Computers and Mathematics with Applications, 15(12):1051– 1052, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890140X.

Anonymous:1988:BRg

[Ano88n] Anonymous. Book reports. Computers and Mathematics with Applications, 16(3):267–268, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218890185X.

Anonymous:1988:BRh

[Ano88o] Anonymous. Book reports. Computers and Mathematics with Applications, 16(9):767–768, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900120.

Anonymous:1988:CPb

[Ano88p] Anonymous. Call for papers. Computers and Mathematics with Applications, 15(3):iv, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901770.

Anonymous:1988:CPc

[Ano88q] Anonymous. Call for papers. Computers and Mathematics with Applications, 15(5):v, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901642.

Anonymous:1988:CPA

[Ano88r] Anonymous. Call for papers: Advances in partial differential equations (an annual special issue of Computers & Mathematical with Applications). Computers and Mathematics with Applications, 15(2):iv, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900879.

Anonymous:1988:CG

[Ano88s] Anonymous. Chaos & graphics. Computers and Mathematics with Applications, 16 (9):v–vi, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890017X.

Anonymous:1988:C

[Ano88t] Anonymous. Corrigendum.

Computers and Mathematics with Applications, 16(9):771, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/0898122188900144.

Anonymous:1988:EBa

[Ano88u] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 15(1):i, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900661.

Anonymous:1988:EBb

[Ano88v] Anonymous. Editorial Board.

Computers and Mathematics with Applications, 16(1–2):i, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900181.

Anonymous:1988:ESSa

[Ano88w] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 15(2):i–iii, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900867.

Anonymous:1988:ESSb

[Ano88x] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 15(5):i–iii, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901630.

Anonymous:1988:ESSc

[Ano88y] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 15(9):i–iii, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901319.

Anonymous:1988:ESSd

[Ano88z] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 15(11):i–iii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900442.

Anonymous:1988:ESSe

[Ano88-27] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 15(12):i–iii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901411.

Anonymous:1988:ESSf

[Ano88-28] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 16(4):i–iii, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901514.

Anonymous:1988:F

[Ano88-29] Anonymous. Foreword. Computers and Mathematics with Applications, 15(6–8):xi, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902696.

Anonymous:1988:LCAa

[Ano88-30] Anonymous. List of contents and author index volume 15, 1988. Computers and Mathematics with Applications, 15 (12):i–xi, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901423.

Anonymous:1988:LCAb

[Ano88-31] Anonymous. List of contents and author index volume 16, 1988. Computers and Mathematics with Applications, 16 (12):i–x, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902660.

Anonymous:1988:P

[Ano88-32] Anonymous. Preface. Computers and Mathematics with Applications, 15(10):vii, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901162.

Anonymous:1988:SII

[Ano88-33] Anonymous. Second IMACS international symposium on computational acoustics. Computers and Mathematics with Applications, 15(5):vi, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901666.

Anonymous:1988:SIC

[Ano88-34] Anonymous. Seventh international conference on mathematical and computer modelling.

Computers and Mathematics with Applications, 16(4):vi, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901538.

Anonymous:1988:SDSb

[Ano88-35] Anonymous. Softstrip data strip.

Computers and Mathematics with Applications, 16(1–2):i, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900314.

Anonymous:1988:SDSc

[Ano88-36] Anonymous. Softstrip data strip.

Computers and Mathematics with Applications, 16(9):ii, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900016.

Anonymous:1988:SDSa

[Ano88-37] Anonymous. Softstrip data strip containing the table of contents of this issue. Computers and Mathematics with Applications, 15(11):iv, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900326.

Anonymous:1988:SSSa

[Ano88-38] Anonymous. Software survey section. Computers and Mathematics with Applications, 15 (1):i–iii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900752.

Anonymous:1988:SSSb

[Ano88-39] Anonymous. Software survey section. Computers and Mathematics with Applications, 15 (3):i–iii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901769.

Anonymous:1988:SSSc

[Ano88-40] Anonymous. Software survey section. Computers and Mathematics with Applications, 16 (3):i–iii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901861.

Anonymous:1988:SSSd

[Ano88-41] Anonymous. Software survey section. Computers and Mathematics with Applications, 16 (9):i–iv, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900156.

Anonymous:1988:SSSe

[Ano88-42] Anonymous. Software survey section. Computers and Mathematics with Applications, 16 (12):i–iii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902647.

Anonymous:1988:SAT

[Ano88-43] Anonymous. Symposium on advances and trends in computational structural mechanics and fluid dynamics. Computers and Mathematics with Applications, 16(4):v, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901526.

Anonymous:1988:WRC

[Ano88-44] Anonymous. Workshop on reliability in computational mechanics. Computers and Mathematics with Applications, 16(9):v, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900168.

Anonymous:1989:AIb

[Ano89a] Anonymous. About this issue. Computers and Mathematics with Applications, 18 (10–11):vii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900011.

Anonymous:1989:AIa

[Ano89b] Anonymous. Author index. Computers and Mathematics with Applications, 17(12):x, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189900503.

Anonymous:1989:AG

[Ano89c] Anonymous. Author’s guidelines. Computers and Mathematics with Applications, 18 (10–11):ix, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900023.

Anonymous:1989:BRa

[Ano89d] Anonymous. Book reports. Computers and Mathematics with Applications, 17(7):1161, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900461.

Anonymous:1989:BRb

[Ano89e] Anonymous. Book reports. Computers and Mathematics with Applications, 17(10):1429– 1430, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122189900254.

Anonymous:1989:BRc

[Ano89f] Anonymous. Book reports.

Computers and Mathematics with Applications, 17(12):1587– 1588, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900588.

Anonymous:1989:BRd

[Ano89g] Anonymous. Book reports. Computers and Mathematics with Applications, 18(4):399–400, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189902241.

Anonymous:1989:BRe

[Ano89h] Anonymous. Book reports. Computers and Mathematics with Applications, 18(5):479–480, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189902447.

Anonymous:1989:BRf

[Ano89i] Anonymous. Book reports. Computers and Mathematics with Applications, 18(9):847–848, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218990182X.

Anonymous:1989:EBa

[Ano89j] Anonymous. Editorial Board. Computers and Mathematics with Applications, 17(1–3):i–iii, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

                                        www.sciencedirect.com/science/
                                        article/pii/0898122189901405.

Anonymous:1989:EBb

[Ano89k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 18(1–3):i, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189901181.

Anonymous:1989:ESSa

[Ano89l] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 17(7):i–iii, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900485.

Anonymous:1989:ESSb

[Ano89m] Anonymous. Editorial: Software survey section. Computers and Mathematics with Applications, 17(10):i–iii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900266.

Anonymous:1989:EAN

[Ano89n] Anonymous. Errata to “Addendum to ‘A new approach to causal inference in mortality studies with a sustained exposure period-application to control of the healthy worker survivor effect’ ”. Computers and Mathematics with Applications, 18(5):477, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902435.

See [Rob87a].

Anonymous:1989:EFC

[Ano89o] Anonymous. Errata to “The foundations of confounding in epidemiology”. Computers and Mathematics with Applications, 18(5):475, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902423.

See [Ano89o].

Anonymous:1989:LC

[Ano89p] Anonymous. List of contents. Computers and Mathematics with Applications, 17 (12):iii–ix, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900497.

Anonymous:1989:LCA

[Ano89q] Anonymous. List of contents and author index. Computers and Mathematics with Applications, 18(12):iii–x, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900345.

Anonymous:1989:PAS

[Ano89r] Anonymous. Papers to appear in symmetry 2: Unifying human understanding (part 2). Computers and Mathematics with Applications, 17(1–3):ix, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901417.

Anonymous:1989:Pa

[Ano89s] Anonymous. Preface. Computers and Mathematics with Applications, 17(1–3):xi, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901429.

Anonymous:1989:Pb

[Ano89t] Anonymous. Preface. Computers and Mathematics with Applications, 17(8–9):vii–viii, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900862.

Anonymous:1989:SDS

[Ano89u] Anonymous. Softstrip data strip. Computers and Mathematics with Applications, 18 (10–11):i, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900175.

Anonymous:1989:SRa

[Ano89v] Anonymous. Software report. Computers and Mathematics with Applications, 17 (7):1163, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900473.

Anonymous:1989:SRe

[Ano89w] Anonymous. Software report. Computers and Mathematics with Applications, 18 (9):849, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901831.

Anonymous:1989:SRb

[Ano89x] Anonymous. Software reports. Computers and Mathematics with Applications, 17(11): 1507–1509, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900849.

Anonymous:1989:SRc

[Ano89y] Anonymous. Software reports. Computers and Mathematics with Applications, 17(12): 1589–1590, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218990059X.

Anonymous:1989:SRd

[Ano89z] Anonymous. Software reports. Computers and Mathematics with Applications, 18 (8):769–770, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902332.

Anonymous:1989:SSSa

[Ano89-27] Anonymous. Software survey section. Computers and Mathematics with Applications, 17 (12):i–ii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900606.

Anonymous:1989:SSSb

[Ano89-28] Anonymous. Software survey section. Computers and Mathematics with Applications, 18 (4):i–ii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902253.

Anonymous:1989:SSSc

[Ano89-29] Anonymous. Software survey section. Computers and Mathematics with Applications, 18 (5):i–ii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902459.

Anonymous:1989:SSSd

[Ano89-30] Anonymous. Software survey section. Computers and Mathematics with Applications, 18 (8):i–ii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902344.

Anonymous:1989:SSSe

[Ano89-31] Anonymous. Software survey section. Computers and Mathematics with Applications, 18 (9):i–ii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901843.

Anonymous:1989:SSSf

[Ano89-32] Anonymous. Software survey section. Computers and Mathematics with Applications, 18 (12):i–ii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900333.

Antonios:1989:OPR

[Ant89] M. N. Antonios. Optimization problems related to water quality control in aquatic ecosystems. Computers and Mathematics with Applications, 18(10– 11):851–870, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900047.

Aoki:1989:TSS

[Aok89] M. Aoki. A two-step state space time series modeling method. Computers and Mathematics with Applications, 17(8– 9):1165–1176, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900874.

Apagyi:1989:SMT

[Apa89] M. Apagyi. Symmetries in music teaching. Computers and Mathematics with Applications, 17(4– 6):671–695, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902551.

Al-Rabeh:1984:NEU

[AR84]  A. H. Al-Rabeh.  A note

on the Enright updating technique. Computers and Mathematics with Applications, 10 (6):405–411, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900701.

Abadir:1985:FST

[AR85] M. S. Abadir and H. K. Reghbati. Functional specification and testing of logic circuits. Computers and Mathematics with Applications, 11(12):1143– 1153, December 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901026.

Abadir:1987:FTG

[AR87a] M. S. Abadir and H. K. Reghbati. Functional test generation using binary decision diagrams. Computers and Mathematics with Applications, 13(5– 6):413–430, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900721.

Al-Rabeh:1987:VPE

[AR87b] A. H. Al-Rabeh. A variable parameter embedded Dirk algorithm for the numerical integration of stiff systems of ODEs. Computers and Mathematics with Applications, 13 (4):373–379, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187900058.

Ardema:1987:ATD

[AR87c] M. D. Ardema and N. Rajan. An approach to three-dimensional aircraft pursuit–evasion. Computers and Mathematics with Applications, 13(1–3):97–110, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187900964.

Adelani:1989:SNM

[AR89] L. A. Adelani and E. Y. Rodin. A simple nonlinear model for the exploitation of renewable economic resources. Computers and Mathematics with Applications, 18(5):411–420, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902368.

Arato:1983:PBA

[Ara83]  Matyas Arato.  Probability

bounds and asymptotic properties of error propagation. Computers and Mathematics with Applications, 9(2):307–310, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122183901323.

Achuthan:1988:NAE

[AS88] P. Achuthan and S. Sundar. A new application of the extended Euclidean algorithm for matrix padé approximants. Computers and Mathematics with Applications, 16(4):287–296, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901459.

Achuthan:1989:RAM

[AS89] P. Achuthan and S. Sundar. A recursive algorithm for matrix Padé approximants-the divideand-conquer approach. Computers and Mathematics with Applications, 17(10):1359–1367, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189900199.

Aslanov:1988:CSA

[Asl88] L. A. Aslanov. Crystal symmetry and atomic interactions. Computers and Mathematics with Applications, 16(5–8):443–451, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902349.

Asmuth:1987:AGR

[Asm87] C. Asmuth. An application of group representation theory to picture recognition. Computers and Mathematics with Applications, 13(4):363–365, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187900034.

Asztalos:1980:OCF

[Asz80] D. Asztalos. Optimal control of finite source priority queues with computer system applications. Computers and Mathematics with Applications, 6(4): 425–431, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900498.

Atallah:1985:SDC

[Ata85] Mikhail J. Atallah. Some dynamic computational geometry problems. Computers and Mathematics with Applications, 11 (12):1171–1181, December 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901051.

Avis:1982:MDC

[ATB82] David Avis, Godfried T. Toussaint, and Binay K. Bhattacharya. On the multimodality of distances in convex polygons. Computers and Mathematics with Applications, 8(2): 153–156, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122182900542.

Alt:1982:SBM

[AV82] R. Alt and J. Vignes. Stabilizing Bairstow’s method. Computers and Mathematics with Applications, 8(5):379–387, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900311.

Avis:1981:WCB

[Avi81] David Avis. Worst case bounds for the Euclidean matching problem. Computers and Mathematics with Applications, 7(3): 251–257, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900845.

Ahmed:1986:EEC

[AW86] A. H. Ahmed and K. Wright. Error estimation for collocation solution of linear ordinary differential equations. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1053–1059, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902300.

Agarwal:1989:LPB

[AW89] R. P. Agarwal and P. J. Y.

Wong. Lidstone polynomials and boundary value problems. Computers and Mathematics with Applications, 17(10):

1397–1421, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900230.

Aho:1982:BST

[AWYU82] A. V. Aho, A. D. Wyner, M. Yannakakis, and J. D. Ullman. Bounds on the size and transmission rate of communications protocols. Computers and Mathematics with Applications, 8(3): 205–214, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900438.

Baik:1989:TTV

[Bai89] K. H. Baik. Towards a topological view of databases. Computers and Mathematics with Applications, 18(5):401–409, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902356.

Balaban:1986:SCS

[Bal86a] Alexandru T. Balaban. Symmetry in chemical structures and reactions. Computers and Mathematics with Applications, 12(3– 4):999–1020, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904372.

Bales:1986:HOS

[Bal86b] Laurence A. Bales. Higherorder single-step fully discrete approximations for nonlinear second-order hyperbolic equations. Computers and Mathematics with Applications, 12(4–5): 581–604, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901835.

Bales:1988:EES

[Bal88] L. A. Bales. Error estimates for single step fully discrete approximations for nonlinear second order hyperbolic equations. Computers and Mathematics with Applications, 15(6–8):535–553, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902799.

Balaban:1989:CN

[Bal89] A. T. Balaban. Carbon and its nets. Computers and Mathematics with Applications, 17(1– 3):397–416, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901703.

Bartram:1986:SPI

[Bar86] Ralph H. Bartram. Symmetry of point imperfections in solids. Computers and Mathematics with Applications, 12 (Part B)(1– 2):185–196, January/April 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901501.

Basar:1989:STR

[Bas89] T. Basar. Some thoughts on rational expectations models, and alternate formulations. Computers and Mathematics with Applications, 18(6–7):591–604, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901107.

Blakley:1983:MAI

[BB83] G. R. Blakley and I. Borosh. Modular arithmetic of iterated powers. Computers and Mathematics with Applications, 9(4): 567–581, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901141.

Balasundaram:1984:MFE

[BB84] S. Balasundaram and P. K. Bhattacharyya. A mixed finite element method for fourth order elliptic equations with variable coefficients. Computers and Mathematics with Applications, 10 (3):245–256, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/089812218490052X.

Bojanczyk:1987:SAE

[BB87] Adam W. Bojanczyk and Richard P.

Brent. A systolic algorithm for extended GCD computation. Computers and Mathematics with Applications, 14 (4):233–238, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901301.

Baer:1985:TDM

[BBPW85] R. N. Baer, D. H. Berman, J. S. Perkins, and E. B. Wright. A three-dimensional model for acoustic scattering from rough ocean bathymetry. Computers and Mathematics with Applications, 11(7–8):863– 871, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901816.

Beretta:1986:GSE

[BC86] E. Beretta and V. Capasso. On the general structure of epidemic systems. global asymptotic stability. Computers and Mathematics with Applications, 12 (Part A)(6):677–694, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900544.

Bucy:1988:DSS

[BC88] R. S. Bucy and L. A. Campbell. Determination of steady state behavior for periodic discrete filtering problems. Computers and Mathematics with Applications, 15(2):131–140, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218890082X.

Brunvoll:1986:CSC

[BCC^{+}86] Jon Brunvoll, Bjørg Cyvin, Einar Cyvin, Sven Cyvin, Aage Paus, Martin Stølevik, and Reidar Stølevik. The computerized Soma cube. Computers and Mathematics with Applications, 12 (Part B)(1–2):113– 121, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901458.

Bardos:1987:EER

[BCG87] C. Bardos, V. Cea, and P. Grisvard. Error estimates related to singular perturbations of first-order equations and systems. Computers and Mathematics with Applications, 13(9– 11):801–829, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901659.

Bernhard:1987:RHG

[BCP87] P. Bernhard, A.-L. Colomb, and G. P. Papavassilopoulos. Rabbit and hunter game: Two discrete stochastic formulations. Computers and Mathematics with Applications, 13(1–3):205–225, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901052.

Bouzidi:1985:NSH

[BD85] M. Bouzidi and P. Duhamel.

Non-stationary heat conduction in composite slabs with coupling. application to enclosures: numerical computation of the analytical solution. Computers and Mathematics with Applications, 11(10):1043–1055, October 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900239.

Bona:1986:FDG

[BDK86] Jerry L. Bona, Vassilios A.

Dougalis, and Ohannes A. Karakashian. Fully discrete Galerkin methods for the Korteweg– de Vries equation. Computers and Mathematics with Applications, 12 (Part A)(7): 859–884, July 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900313.

Bellomo:1988:RHE

[BdSM88] N. Bellomo, L. M. de Socio, and R. Monaco. Random heat equation: Solutions by the stochastic adaptive interpolation method. Computers and Mathematics with Applications, 16 (9):759–766, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900119.

Barrett:1985:FMF

[BE85] John W. Barrett and Charles M. Elliott. Fixed mesh finite element approximations to a free boundary problem for an elliptic equation with an oblique derivative boundary condition. Computers and Mathematics with Applications, 11(4):335–345, April 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900586.

Barefoot:1988:COC

[BE88] C. A. Barefoot and R. Entringer. Computer observations of cycles in cubic graphs. Computers and Mathematics with Applications, 15(4):241–246, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902076.

Bellman:1983:NHP

[Bel83] Richard Bellman. Nonlinear hyperbolic partial differential equations and dynamic programming. Computers and Mathematics with Applications, 9(3): 547–548, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900676.

Bencsik:1985:APF

[Ben85] István Bencsik. Approximate predictor and filter for partially observed vector ARMA processes. Computers and Mathematics with Applications, 11(9): 943–947, September 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900987.

Bencze:1989:UPS

[Ben89] L. Bencze. Uncertainty principle and symmetry in metaphors. Computers and Mathematics with Applications, 17(4–6):697– 708, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902563.

Bercovier:1982:BES

[Ber82] Michel Bercovier. On c^{0} beam elements with shear and their corresponding penalty function formulation. Computers and Mathematics with Applications, 8(4):245–256, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900074.

Berrut:1988:RFG

[Ber88] J.-P. Berrut. Rational functions for guaranteed and experimentally well-conditioned global interpolation. Computers and Mathematics with Applications, 15(1):1–16, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900673.

Berczi:1989:STO

[Bér89a]  Sz. Bérczi.  Symmetry and

technology in ornamental art of old Hungarians and Avar– Onogurians from the archaeological finds of the Carpathian Basin, Seventh to Tenth Century A.D. Computers and Mathematics with Applications, 17(4– 6):715–730, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902587.

Bergstrom:1989:OFD

[Ber89b]  A. R. Bergstrom.  Optimal

forecasting of discrete stock and flow data generated by a higher order continuous time system. Computers and Mathematics with Applications, 17(8–9): 1203–1214, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900904.

Butzer:1986:EDCa

[BEW86a] P. L. Butzer, W. Engels, and U. Wipperfürth. An extension of the dyadic calculus with fractional order derivatives. further theory and applications. Computers and Mathematics with Applications, 12 (Part A)(8): 921–943, August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900283.

Butzer:1986:EDCb

[BEW86b] P. L. Butzer, W. Engels, and U. Wipperfürth. An extension of the dyadic calculus with fractional order derivatives: General theory. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1073–1090, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902324.

Bhattacharyya:1988:EFE

[BG88] P. K. Bhattacharyya and S. Gopalsamy. An equilibrium finite element method for fourth order elliptic equations with variable coefficients. Computers and Mathematics with Applications, 15 (5):405–424, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901617.

Bertuzzi:1987:SCK

[BGG^{+}87] A. Bertuzzi, A. Gandolfi, A. Germani, R. Vitelli, G. Badaracco, and G. Starace. Study of cell kinetics by computer-analyzed flow cytometric histograms. Computers and Mathematics with Applications, 14(9–12):763–770, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187902252.

Bayliss:1985:NSH

[BGT85] A. Bayliss, C. I. Goldstein, and E. Turkel. The numerical solution of the Helmholtz equation for wave propagation problems in underwater acoustics. Computers and Mathematics with Applications, 11(7–8): 655–665, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901622.

Bhat:1982:PAH

[Bha82] K. V. S. Bhat. On the properties of arbitrary hypercubes. Computers and Mathematics with Applications, 8(5):339–342, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122182900281.

Baringhaus:1988:SEP

[BHM88] L. Baringhaus, N. Henze, and D. Morgenstern. Some elementary proofs of the normality of XY (X^{2} + Y^{2})^{1}2 when X and Y are normal. Computers and Mathematics with Applications, 15(11):943–944, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900387.

Bokhari:1982:FOD

[BHO82] Shahid H. Bokhari, M. Yousuff Hussaini, and Steven A. Orszag. Fast orthogonal derivatives on the star. Computers and Mathematics with Applications, 8(5): 367–377, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218290030X.

Brandmuller:1986:SA

[BHW86] J. Brandmüller, B. Hrouda, and A. V. Wickede. Symmetry in archaeology. Computers and Mathematics with Applications, 12(3– 4):783–787, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186904244.

Bigham:1988:IIP

[Big88] J. Bigham. The inductive inference of pattern recognition rules which are capable of a linguistic interpretation. Computers and Mathematics with Applications, 15(10):839–862, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901216.

Birkhoff:1985:SMP

[Bir85] Garrett Birkhoff. Some mathematical problems of numerical ocean acoustics. Computers and Mathematics with Applications, 11(7–8):643–654, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122185901610.

Baker:1987:NFE

[BJ87] G. A. Baker and W. N. Jureidini. A nonstandard finite element method for the stationary two-dimensional Navier–Stokes equations. Computers and Mathematics with Applications, 13 (4):387–400, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900071.

Basar:1987:WCD

[BK87] T. Basar and P. R. Kumar.

On worst case design strategies. Computers and Mathematics with Applications, 13(1– 3):239–245, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901076.

Blanco:1988:ADM

[BKL88] T. A. Blanco, I. A. Krass, and C. Louie. An application of dynamic modeling to the sea shore rotation planning problem in the Navy. Computers and Mathematics with Applications, 16(12): 1065–1077, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902635.

Ballantyne:1981:NDA

[BL81] A. M. Ballantyne and D. S. Lankford. New decision algorithms for finitely presented commutative semigroups. Computers and Mathematics with Applications, 7(2):159–165, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901152.

Blasio:1983:ABA

[Bla83] Gabriella Di Blasio. Asymptotic behavior of an age-structured fish population. Computers and Mathematics with Applications, 9(3):377–381, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900524.

Blackmore:1986:MTC

[Bla86] Denis Blackmore. The mathematical theory of chaos. Computers and Mathematics with Applications, 12(3–4):1039–1045, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0898122186904396.

Bloom:1985:SNH

[Blo85] Frederick Bloom. Systems of nonlinear hyperbolic equations associated with problems of classical electromagnatic theory. Computers and Mathematics with Applications, 11(1–3): 261–279, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901518.

Bloom:1986:INC

[Blo86]  Frederick Bloom.  The in-

fluence of nonlinear conduction on singularity formation in the intense plane-wave, nonlinear dielectric interaction problem. Computers and Mathematics with Applications, 12(4–5): 477–489, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901756.

Bloom:1987:CIG

[Blo87] F. Bloom. A condition implying the global existence of C^{1} solutions for a class of nonlinear transmission lines. Computers and Mathematics with Applications, 13(9–11):861–879, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901696.

Bloom:1988:AGE

[Blo88a] F. Bloom. Almost global existence in the plane wave- nonlinear dielectric interaction problem. Computers and Mathematics with Applications, 15(6– 8):491–510, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902763.

Bloom:1988:FSD

[Blo88b] F. Bloom. Formation of shock discontinuities for a class of nonlinear transmission lines. Computers and Mathematics with Applications, 15(6–8):459–472, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902738.

Bermudez:1981:DMS

[BM81] A. Bermúdez and C. Moreno. Duality methods for solving variational inequalities. Computers and Mathematics with Applications, 7(1):43–58, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900067.

Baart:1984:RCC

[BM84a] M. L. Baart and R. J. Y. McLeod. Recursive calculation of curved finite elements stiffness matrices. Computers and Mathematics with Applications, 10(4– 5):387–397, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122184900671.

Berman:1984:EFS

[BM84b] Joel Berman and Masao Mukaidono. Enumerating fuzzy switching functions and free Kleene algebras. Computers and Mathematics with Applications, 10 (1):25–35, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218490083X.

Babuska:1985:CPS

[BM85a] I. Babuška and R. C. Morgan. Composites with a periodic structure: mathematical analysis and numerical treatment. Computers and Mathematics with Applications, 11(10): 995–1005, October 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900203.

Bellman:1985:SAH

[BM85b] Richard Bellman and Semyon M.

Meerkov. Simplified analysis of a hyperbolic system. Computers and Mathematics with Applications, 11(1–3):281–282, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218590152X.

Bencsik:1985:PFP

[BM85c] István Bencsik and György Michaletzky. Prediction and filtration of a partially observed vector ARMA system of first order. Computers and Mathematics with Applications, 11 (4):321–328, April 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900562.

Bencsik:1988:GLS

[BM88] I. Bencsik and Gy. Michaletzky.

Generalized least squares innovation representation. Computers and Mathematics with Applications, 15(5):359–365, ???? 1988.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901575.

Bogdanowicz:1989:NOP

[BM89a] Z. R. Bogdanowicz and T. G. Moore. A new optimal packing algorithm for telecommunications networks planning. Computers and Mathematics with Applications, 18(8):739–744, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189902307.

Busenberg:1989:CRD

[BM89b] S. N. Busenberg and J. M. Mahaffy. A compartmental reaction–diffusion cell cycle model. Computers and Mathematics with Applications, 18(10– 11):883–892, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900060.

See [BM90].

Busenberg:1990:ECR

[BM90] S. N. Busenberg and J. M. Mahaffy. Errata to “A compartmental reaction–diffusion cell cycle model”: Computers Math. Applic. 18 (10/11), 883–892 (1989). Computers and Mathematics with Applications, 20 (7):73–74, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812219090350S.

See [BM89b].

Bylander:1985:CLE

[BMC85] Tom Bylander, Sanjay Mittal, and B. Chandrasekaran. CSRL: a language for expert systems for diagnosis. Computers and Mathematics with Applications, 11(5):449–456, May 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900501.

Beaudoin:1981:CCC

[BN81] Y. Beaudoin and M. A. H. Nerenberg. Computer codes for computational assistance in the study of asymptotic stability. Computers and Mathematics with Applications, 7(1):7–16, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900031.

Brebbia:1986:SPH

[BN86] C. A. Brebbia and D. Nardini. Solution of parabolic and hyperbolic time dependent problems using boundary elements. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1061–1072, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186902312.

Bokic:1988:TCS

[Bok88] G. B. Bokiç. Theory of crystal symmetry upon the derivation of the 230 space groups. Computers and Mathematics with Applications, 16(5–8):479–485, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188902386.

Bose:1987:DAR

[Bos87] B. Bose. 2-dimensional arithmetic residue check codes. Computers and Mathematics with Applications, 13(5–6):547–554, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187900824.

Boyarsky:1986:CO

[Boy86] A. Boyarsky. Computer orbits. Computers and Mathematics with Applications, 12 (Part A)(10):1057–1064, October 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900118.

Bozoki:1981:APF

[Boz81] Eva Bozoki. An algorithm for programming function generators. Computers and Mathematics with Applications, 7 (1):1–5, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218190002X.

Begue:1985:HSP

[BP85] C. Begue and O. Pironneau. Hyperbolic systems with periodic boundary conditions: an example with Euler equations. Computers and Mathematics with Applications, 11(1–3):113–128, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901415.

Bini:1986:FPA

[BP86] D. Bini and V. Pan. Fast parallel algorithms for polynomial division over an arbitrary field of constants. Computers and Mathematics with Applications, 12 (Part A)(11):1105– 1118, November 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900155.

Baowen:1988:FBR

[BPXY88] L. Baowen, W. Peizhuang, L. Xihui, and S. Yong. Fuzzy bags and relations with set-valued statistics. Computers and Mathematics with Applications, 15 (10):811–818, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901186.

Bellomo:1986:TEF

[BR86] N. Bellomo and R. Riganti. Time evolution and fluctuations of the probability density and entropy function for a class of nonlinear stochastic systems in mathematical physics. Computers and Mathematics with Applications, 12 (Part A)(6):663–675, June 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186900532.

Brauer:1983:NAD

[Bra83] Fred Brauer. Nonlinear agedependent population growth under harvesting. Computers and Mathematics with Applications, 9(3):345–352, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900226.

Brandmuller:1986:ENM

[Bra86] J. Brandmüller. An extension of the Neumann–Minnigerode– Curie principle. Computers and Mathematics with Applications, 12 (Part B)(1–2):97– 100, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122186901434.

Brossier:1987:ABP

[Bro87] F. Brossier. Asymptotic behaviour of the perturbation of a given mean flow. Computers and Mathematics with Applications, 13(8):687–709, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900423.

Brown:1989:SGH

[Bro89] R. Brown. Symmetry, groupoids and higher-dimensional analogues. Computers and Mathematics with Applications, 17 (1–3):49–57, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901478.

Brunner:1987:IRK

[Bru87] H. Brunner. Implicit Runge– Kutta–Nyström methods for general second-order Volterra integro-differential equations. Computers and Mathematics with Applications, 14(7):549– 559, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900502.

Bhattacharyya:1980:NFS

[BS80a] P. K. Bhattacharyya and G. T. Symm. A novel formulation and solution of the plane elastostatic displacement problem. Computers and Mathematics with Applications, 6(4):443–448, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900516.

Bucy:1980:NFA

[BS80b] Richard S. Bucy and Kenneth D. Senne. Nonlinear filtering algorithms for vector processing machines. Computers and Mathematics with Applications, 6(3): 317–338, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900401.

Bernstein:1985:GAA

[BS85] Ira B. Bernstein and Timothy M. Smith. Geometric acoustics as applied to ocean ranging and improvements. Computers and Mathematics with Applications, 11(7–8):787–793, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122185901737.

Bakry:1988:PFG

[BS88] S. H. Bakry and M. Shatila. Pascal functions for the generation of random numbers. Computers and Mathematics with Applications, 15(11):969–973, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900429.

Blanchet-Sadri:1989:GED

[BS89] F. Blanchet-Sadri. Games, equations and the dot-depth hierarchy. Computers and Mathematics with Applications, 18 (9):809–822, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990179X.

Brezillon:1981:NEF

[BSPSM81] Patrick Brezillon, Jean-François Staub, Anne-Marie Perault-Staub, and Gérard Milhaud. Numerical estimation of the first order derivative: approximate evaluation of an optimal step. Computers and Mathematics with Applications, 7(4):333–347, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122181900626.

Brunner:1989:PSC

[BT89] H. Brunner and T. Tang. Polynomial spline collocation methods for the nonlinear Basset equation. Computers and Mathematics with Applications, 18 (5):449–457, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902393.

Burr:1982:CCR

[Bur82] S. A. Burr. Computing cube roots when a fast square root is available. Computers and Mathematics with Applications, 8(3):181–183, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122182900414.

Butler:1985:SBN

[But85] John Scott Butler. The statistical bias of numerically integrated statistical procedures. Computers and Mathematics with Applications, 11(6):587–593, June 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122185900410.

Ben-yu:1986:SET

[ByMS86] G. Ben-yu, A. R. Mitchell, and B. D. Sleeman. Spatial effects in a two-dimensional model of the budworm-balsam fir ecosystem. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1117–1132, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122186902361.

Bjorck:1989:ESS

[BZ89] Å. Björck and Z. Zlatev. Exploiting the separability in the solution of systems of linear ordinary differential equations. Computers and Mathematics with Applications, 18(5):421–438, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990237X.

Costanza:1986:MCS

[CA86] M. C. Costanza and T. Ashikaga. Monte Carlo study of forward stepwise discrimination based on small samples. Computers and Mathematics with Applications, 12 (Part A)(2): 245–252, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900775.

Cahlon:1988:ETR

[Cah88] B. Cahlon. Existence theorems of random solutions to stochastic functional integral equations. Computers and Mathematics with Applications, 15(6– 8):473–481, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890274X.

Campbell:1989:GDP

[Cam89] P. J. Campbell. The geometry of decoration on prehistoric Pueblo pottery from Starkweather Ruin. Computers and Mathematics with Applications, 17(4–6):731–749, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902599.

Cekirge:1989:DWD

[CARG89] H. M. Cekirge, A. H. Al-Rabeh, and N. Gunay. Determining the wind-driven surface currents for prediction of movements of oil slicks in the Arabian Gulf. Computers and Mathematics with Applications, 17(11): 1449–1453, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900771.

Cash:1983:ISI

[Cas83a] J. R. Cash. The integration of stiff initial value problems in ODEs using modified extended backward differentiation formulae. Computers and Mathematics with Applications, 9(5): 645–657, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183901220.

Cash:1983:VOD

[Cas83b] J. R. Cash. A variable order deferred correction algorithm for the numerical solution of nonlinear two point boundary value problems. Computers and Mathematics with Applications, 9(2): 257–265, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901293.

Cash:1984:DIR

[Cas84] J. R. Cash. Diagonally implicit Runge–Kutta formulae for the numerical integration of nonlinear two-point boundary value problems. Computers and Mathematics with Applications, 10 (2):123–137, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900439.

Cash:1985:ARK

[Cas85a] J. R. Cash. Adaptive Runge– Kutta methods for nonlinear two-point boundary value problems with mild boundary layers. Computers and Mathematics with Applications, 11 (6):605–619, June 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900434.

Cash:1985:BEE

[Cas85b] J. R. Cash. Block embedded explicit Runge–Kutta methods. Computers and Mathematics with Applications, 11 (4):395–409, April 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900616.

Cash:1986:NIN

[Cas86] J. R. Cash. Numerical integration of non-linear two-point boundary-value problems using iterated deferred corrections — I: a survey and comparison of some one-step formulae. Computers and Mathematics with Applications, 12 (Part A)(10): 1029–1048, October 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218690009X.

Chowdhury:1988:OAC

[CB88] S. Chowdhury and M. A. Breuer. Optimization algorithms for a class of nonlinear programming problems. Computers and Mathematics with Applications, 15 (3):175–184, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901691.

Craine:1989:SSM

[CB89] R. Craine and D. Bowman. A state space model of the economic fundamentals. Computers and Mathematics with Applications, 18(6–7):555–564, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901077.

Cyvin:1989:TAB

[CBC89] S. J. Cyvin, J. Brunvoll, and B. N. Cyvin. Topological aspects of benzenoids and coronoids, including “snowflakes” and “laceflowers”. Computers and Mathematics with Applications, 17(1– 3):355–374, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901685.

Cuyt:1985:REA

[CdC85] Annie Cuyt and Paul Van der Cruyssen. Rounding error analysis for forward continued fraction algorithms. Computers and Mathematics with Applications, 11(6):541–564, June 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900379.

Cercone:1983:P

[Cer83] Nick Cercone. Preface. Computers and Mathematics with Applications, 9(1):v, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900020.

Cerbo:1989:US

[Cer89a] M. Di Cerbo. Urban symmetry. Computers and Mathematics with Applications, 17(4– 6):907–911, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902678.

Cerchi:1989:SSM

[Cer89b] M. Cerchi. State space methods in asset pricing. Computers and Mathematics with Applications, 18(6–7):581–590, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901090.

Cooper:1984:DPA

[CF84a] Mary W. Cooper and Keyvan Farhangian. A dynamic programming algorithm for multiple-choice constraints. Computers and Mathematics with Applications, 10(3):279– 282, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900555.

Cox:1984:ALS

[CF84b] C. L. Cox and G. J. Fix. On the accuracy of least squares methods in the presence of corner singularities. Computers and Mathematics with Applications, 10(6):463–475, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900774.

Chen:1986:NFP

[CF86] C. F. Chen and G. Freeman.

A new formula for partial fraction expansion of a transfer matrix. Computers and Mathematics with Applications, 12 (Part A)(9):963–968, September 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900210.

Codenotti:1987:IMP

[CF87a] B. Codenotti and P. Favati. Iterative methods for the parallel solution of linear systems. Computers and Mathematics with Applications, 13(7):631–633, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901258.

Codenotti:1987:LRM

[CF87b] B. Codenotti and P. Favati. Low rank modifications of Jacobi and JOR iterative methods. Computers and Mathematics with Applications, 13(7):617–621, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901234.

Codenotti:1987:NTS

[CF87c] B. Codenotti and P. Favati. New techniques for the solution of linear systems by iterative methods. Computers and Mathematics with Applications, 14(6): 425–428, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218790054X.

Chiang:1988:LAP

[CF88] H.-W. D. Chiang and S. Fleeter. Locally analytical prediction of the steady inviscid incompressible flow through an airfoil cascade. Computers and Mathematics with Applications, 15 (3):221–233, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901745.

Cyvin:1986:KST

[CG86] S. J. Cyvin and I. Gutman. Kekulé structures and their symmetry properties. Computers and Mathematics with Applications, 12(3–4):859–876, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218690430X.

Chen:1989:NTS

[CG89] H.-F. Chen and L. Guo. Nonstationary time series identification. Computers and Mathematics with Applications, 17(8– 9):1247–1258, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900928.

Chen:1989:FCTa

[CH89] S.-G. Chen and P. Y. Hsieh.

Fast computation of the N-th root. Computers and Mathematics with Applications, 17(10): 1423–1427, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900242.

Chavey:1989:TRP

[Cha89] D. Chavey. Tilings by regular polygons — II: a catalog of tilings. Computers and Mathematics with Applications, 17(1– 3):147–165, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901569.

Chen:1984:SLP

[Che84a] Reuven Chen. Solution of location problems with radial cost functions. Computers and Mathematics with Applications, 10 (1):87–94, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900890.

Cherenack:1984:CCS

[Che84b] P. Cherenack. Conditions for cubic spline interpolation on triangular elements. Computers and Mathematics with Applications, 10(3):235–244, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900518.

Chen:1985:SMP

[Che85]  Reuven Chen.  Solution of

minimax problems using equivalent differentiable functions. Computers and Mathematics with Applications, 11(12):1165– 1169, December 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218590104X.

Cheng:1987:MMD

[Che87a] T. C. E. Cheng. Minimizing the maximum deviation of job completion time about a common due-date. Computers and Mathematics with Applications, 14(4):279–283, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901362.

Cheng:1987:OTW

[Che87b] T. C. E. Cheng. Optimal total-work-content-power duedate determination and sequencing. Computers and Mathematics with Applications, 14(8): 579–582, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901842.

Chechin:1989:CGT

[Che89] G. M. Chechin. Computers and group-theoretical methods for studying structural phase transitions. Computers and Mathematics with Applications, 17(1– 3):255–278, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901624.

Choi:1984:TSS

[Cho84] S. C. Choi. Two-stage sequential tests for proportions and means with follow-up period. Computers and Mathematics with Applications, 10(4–5):343–349, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122184900622.

Choi:1985:AST

[Cho85] Sung C. Choi. Approximate sequential test for the Behrens– Fisher problem. Computers and Mathematics with Applications, 11(9):897–902, September 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900926.

Choi:1986:DCO

[Cho86] S. C. Choi. Discrimination and classification: Overview. Computers and Mathematics with Applications, 12 (Part A)(2): 173–177, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900702.

Chorbachi:1989:TBB

[Cho89] W. K. Chorbachi. In the Tower of Babel: Beyond symmetry in Islamic design. Computers and Mathematics with Applications, 17(4–6):751–789, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902605.

Christodoulakis:1989:ELL

[Chr89] N. M. Christodoulakis. Extensions of linearisation to large econometric models with rational expectations. Computers and Mathematics with Applications, 18(6–7):629–642, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901132.

Chung:1984:OLA

[Chu84] F. R. K. Chung. On optimal linear arrangements of trees. Computers and Mathematics with Applications, 10(1):43–60, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122184900853.

Chujo:1986:SPS

[Chû86] Riichirô Chûjô. Symmetry in polymer science. Computers and Mathematics with Applications, 12(3–4):849–857, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186904293.

Chukwu:1988:TCT

[Chu88a] E. N. Chukwu. The time control theory of linear differential equations of neutral type. Computers and Mathematics with Applications, 16(10–11):851–866, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901952.

Chung:1988:SSS

[Chu88b] K. P. Chung. Steady state solution and convergence rate of time-dependent Markov chains of queuing networks. Computers and Mathematics with Applications, 15(5):351–358, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901563.

Cooper:1981:WPR

[CK81] Leon Cooper and I. Norman Katz. The Weber problem revisited. Computers and Mathematics with Applications, 7(3): 225–234, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900821.

Chau:1987:SSD

[CK87] K. K. Chau and C. R. Kime. Selective I/O scan: a diagnosable design technique for VLSI systems. Computers and Mathematics with Applications, 13(5– 6):485–502, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900782.

Chechin:1988:RBM

[CK88a] G. M. Chechin and V. A. Koptsik. Relation between multidimensional representations of the fedorov groups and the groups of color symmetry. Computers and Mathematics with Applications, 16(5–8):521–536, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902428.

Chuprunov:1988:DSG

[CK88b] E. V. Chuprunov and T. S.

Kuntsevich. n-dimensional space groups and regular point systems. Computers and Mathematics with Applications, 16(5– 8):537–543, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890243X.

Choi:1989:ASC

[CK89] U. Jin Choi and Do Y. Kwak.

Almost sure convergence of Galerkin approximations for a heat equation with a random initial condition. Computers and Mathematics with Applications, 18(12):1057–1063, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900321.

Cercone:1983:MAM

[CKB83] Nick Cercone, Max Krause, and John Boates. Minimal and almost minimal perfect hash function search with application to natural language lexicon design. Computers and Mathematics with Applications, 9(1): 215–231, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900160.

Chen:1981:NLP

[CL81] C. F. Chen and K. K. Leung. A new look at partial fraction expansion from a high-level language viewpoint. Computers and Mathematics with Applications, 7(5):361–367, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181901218.

Chan:1989:VSP

[CL89a] W. L. Chan and W. T. Lau. Vector saddle-point and distributed parameter differential games. Computers and Mathematics with Applications, 18(1– 3):195–207, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901375.

Cheng:1989:SOE

[CL89b] T. C. E. Cheng and S. Li. Some observations and extensions of the optimal TWK-power duedate determination and sequencing problem. Computers and Mathematics with Applications, 17(7):1103–1107, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900400.

Corless:1989:ACA

[CL89c] M. Corless and G. Leitmann.

Adaptive controllers for avoidance or evasion in an uncertain environment: Some examples. Computers and Mathematics with Applications, 18(1– 3):161–170, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901338.

Cekirge:1986:AST

[CLF86] H. M. Cekirge, R. W. Lardner, and R. J. Fraga. Adaptation of the solution of the twodimensional tidal equations using the method of characteristics to wind-induced currents and storm surges. Computers and Mathematics with Applications, 12 (Part A)(10):1081–1090, October 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186900131.

Chiccoli:1987:NMG

[CLM87] C. Chiccoli, S. Lorenzutta, and G. Maino. A numerical method for generalized exponential integrals. Computers and Mathematics with Applications, 14 (4):261–268, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901349.

Clenshaw:1986:GEL

[CLOT86] C. W. Clenshaw, Daniel W. Lozier, Frank W. J. Olver, and P. R. Turner. Generalized exponential and logarithmic functions. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1091–1101, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186902336.

Corless:1987:ACA

[CLS87] M. Corless, G. Leitmann, and J. M. Skowronski. Adaptive control for avoidance or evasion in an uncertain environment. Computers and Mathematics with Applications, 13(1–3):1–11, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900903.

Carraway:1988:TAG

[CM88] R. L. Carraway and T. L. Morin. Theory and applications of generalized dynamic programming: an overview. Computers and Mathematics with Applications, 16(10–11):779–788, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901885.

Chernick:1988:EER

[CMN88] M. R. Chernick, V. K. Murthy, and C. D. Nealy. Estimation of error rate for linear discriminant functions by resampling: Non-Gaussian populations. Computers and Mathematics with Applications, 15(1):29–37, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900697.

Chan:1987:POL

[CN87] W. L. Chan and S. K. Ng.

Partially observable linear– quadratic stochastic pursuit– evasion games. Computers and Mathematics with Applications, 13(1–3):181–189, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901039.

Cayco:1989:ANS

[CN89] M. E. Cayco and R. A. Nicolaides. Analysis of nonconforming stream function and pressure finite element spaces for the Navier–Stokes equations. Computers and Mathematics with Applications, 18(8):745–760, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189902319.

Chernick:1988:RTE

[CNM88] M. R. Chernick, C. D. Nealy, and V. K. Murthy. Resamplingtype error rate estimation for linear discriminant functions: Pearson VII distributions. Computers and Mathematics with Applications, 15(11):897–902, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900338.

Coleman:1983:NBS

[CO83] Bernard D. Coleman and David R. Owen. On the nonequilibrium behavior of solids that transport heat by second sound. Computers and Mathematics with Applications, 9(3):527–546, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900664.

Jones:1985:AAI

[(Co85] Marlene Jones (Colbourn). Applications of artificial intelligence within education. Computers and Mathematics with Applications, 11(5):517–526, May 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900549.

Codenotti:1988:BCS

[Cod88] B. Codenotti. The bit-cost of some algorithms for the solution of linear systems. Computers and Mathematics with Applications, 15(11):933–937, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900363.

Cohen:1981:NNF

[Coh81] Edgar A. Cohen, Jr. A note on normal functions of normal random variables. Computers and Mathematics with Applications, 7(5):395–400, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901243.

Colbourn:1982:BRE

[Col82] Charles J. Colbourn. Book review: Elements of the theory of computation: By H. R. Lewis and C. H. Papadimitriou. Prentice– Hall, Englewood Cliffs, New Jersey 1981. 466 pp. Computers and Mathematics with Applications, 8(3):241, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900475.

Coleman:1985:CDP

[Col85]  Bernard D. Coleman.  On

the cold drawing of polymers.

Computers and Mathematics with Applications, 11(1–3):35– 65, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901373.

Conway:1987:BRA

[Con87a] G. Conway. Book review: Advanced graphics with the IBM personal computer: By I. O. Angell. Wiley, New York (1985). 273 pp. Computers and Mathematics with Applications, 14(9– 12):951–952, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902471.

Conway:1987:F

[Con87b] G. Conway. Fastback. Computers and Mathematics with Applications, 14(9–12):956–957, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902501.

Conway:1987:L

[Con87c] G. Conway. Lightning. Computers and Mathematics with Applications, 14(9–12):957–958, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902513.

Coxeter:1986:GPG

[Cox86] H. S. M. Coxeter. The generalized Petersen graph g(24, 5). Computers and Mathematics with Applications, 12(3–4): 579–583, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904128.

Coxeter:1989:TO

[Cox89] H. S. M. Coxeter. Trisecting an orthoscheme. Computers and Mathematics with Applications, 17(1–3):59–71, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990148X.

Cheng:1987:TTD

[CP87] W.-T. Cheng and J. H. Patel.

Testing in two-dimensional iterative logic arrays. Computers and Mathematics with Applications, 13(5–6):443–454, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900745.

Codenotti:1988:MVM

[CP88] B. Codenotti and C. Puglisi.

Matrix–vector multiplication: Parallel algorithms and architectures. Computers and Mathematics with Applications, 16(12): 1057–1063, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902623.

Crowe:1986:MPH

[Cro86] D. W. Crowe. The mosaic patterns of H. J. Woods. Computers and Mathematics with Applications, 12 (Part B)(1–2):407– 411, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901616.

Craig:1988:ADI

[CS88a] I. J. D. Craig and A. D. Sneyd. An alternating-direction implicit scheme for parabolic equations with mixed derivatives. Computers and Mathematics with Applications, 16(4):341–350, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901502.

Cullen:1988:NMB

[CS88b] M. J. P. Cullen and G. J. Shutts. Numerical modelling of balanced atmospheric flow. Computers and Mathematics with Applications, 16(1–2):69–91, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900259.

Craig:1989:AMP

[CS89] I. J. D. Craig and A. D. Sneyd.

The acceleration of matrix power methods by cyclic variations of the shift parameter. Computers and Mathematics with Applications, 17(7):1149–1159, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218990045X.

Chandrasekharaiah:1988:MTE

[CSD88] D. S. Chandrasekharaiah, K. S.

Srinath, and L. Debnath. Magneto-thermo-elastic disturbances with thermal relaxation in a solid due to heat sources. Computers and Mathematics with Applications, 15(6–8):483–490, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902751.

Chan:1985:DSP

[CSL85a] Tony F. Chan, Longjun Shen, and Ding Lee. Difference schemes for the parabolic wave equation in ocean acoustics. Computers and Mathematics with Applications, 11(7–8):747– 754, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901695.

Chan:1985:SES

[CSL85b] Tony F. Chan, Longjun Shen, and Ding Lee. A stable explicit scheme for the ocean acoustic wave equation. Computers and Mathematics with Applications, 11(9):929–936, September 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900963.

Chernikov:1989:SC

[CSUZ89] A. A. Chernikov, R. Z. Sagdeev, D. A. Usikov, and G. M. Zaslavsky. Symmetry and chaos. Computers and Mathematics with Applications, 17 (1–3):17–32, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901454.

Cummings:1986:RGP

[Cum86] L. A. Cummings. A recurring geometrical pattern in the early renaissance imagination. Computers and Mathematics with Applications, 12(3–4):981– 997, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186904360.

Cushing:1983:BTP

[Cus83] J. M. Cushing. Bifurcation of time periodic solutions of the McKendrick equations with applications to population dynamics. Computers and Mathematics with Applications, 9(3): 459–478, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900603.

Cushing:1985:GBE

[Cus85] J. M. Cushing. Global branches of equilibrium solutions of the McKendrick equations for agestructured population growth. Computers and Mathematics with Applications, 11(1–3):175– 188, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901452.

Cushing:1986:PME

[Cus86]  J. M. Cushing.  Periodic

McKendrick equations for agestructered population growth. Computers and Mathematics with Applications, 12(4–5):513– 526, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690177X.

Cuyt:1982:PAO

[Cuy82] Annie A. M. Cuyt. Padéapproximants in operator theory for the solution of nonlinear differential and integral equations. Computers and Mathematics with Applications, 8(6): 445–466, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122182900190.

Cuyt:1983:APA

[CvdC83] Annie Cuyt and Paul van der Cruyssen. Abstract Padéapproximants for the solution of a sytem of nonlinear equations. Computers and Mathematics with Applications, 9(4): 617–624, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183901190.

Chernick:1985:TSA

[CW85] Michael R. Chernick and David M. Wincelberg. Time series analysis of the strategic petroleum reserve’s brine pipeline test data. Computers and Mathematics with Applications, 11(9): 967–982, September 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901002.

Cooke:1986:SRL

[CW86] Kenneth L. Cooke and Joseph Wiener. Stability regions for linear equations with piecewise continuous delay. Computers and Mathematics with Applications, 12 (Part A)(6):695–701, June 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186900556.

Chen:1989:MSS

[CW89] Y.-Y. Chen and C.-A. Wang.

Multiple steady state horizontal buoyancy induced flows in porous media saturated with cold water. Computers and Mathematics with Applications, 18 (5):439–448, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902381.

Dalton:1985:AIC

[Dal85] G. R. Dalton. Automatic indexed calculation of Wachspress’ rational finite element functions. Computers and Mathematics with Applications, 11 (6):625–641, June 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900458.

David:1986:VPD

[Dav86] Carl W. David. Voronoi polyhedra and the docking problem. Computers and Mathematics with Applications, 12(3–4): 763–766, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186904220.

Davenport:1988:PEF

[DBH88] J. W. Davenport, J. C. Bezdek, and R. J. Hathaway. Parameter estimation for finite mixture distributions. Computers and Mathematics with Applications, 15(10):819–828, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901198.

Dey:1983:EFD

[DD83] Charlie Dey and Suhrit K. Dey. Explicit finite difference predictor and convex corrector with applications to hyperbolic partial differential equations. Computers and Mathematics with Applications, 9(3):549–557, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900688.

Debnath:1989:TML

[DD89] Joyati Debnath and R. S. Dahiya. Theorems on multidimensional Laplace transform for solution of boundary value problems. Computers and Mathematics with Applications, 18(12): 1033–1056, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218990031X.

Demkowicz:1982:SRC

[Dem82] Leszek Demkowicz. On some results concerning the reciprocal formulation for the Signorini’s problem. Computers and Mathematics with Applications, 8(1):57–74, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900396.

Denes:1986:NVP

[Den86a] Agnes Denes. Notes on a visual philosophy. Computers and Mathematics with Applications, 12(3–4):835–848, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186904281.

Denman:1986:IIS

[Den86b] E. D. Denman. Invariant imbedding, scattering processes and projectors in systems. Computers and Mathematics with Applications, 12 (Part A)(6): 703–717, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900568.

DeSanto:1985:SCP

[DeS85] John A. DeSanto. Some computational problems in ocean acoustics. Computers and Mathematics with Applications, 11(7–8):

755–763, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901701.

Devroye:1980:GMI

[Dev80] Luc Devroye. Generating the maximum of independent identically distributed random variables. Computers and Mathematics with Applications, 6(3): 305–315, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900395.

Devroye:1981:HRA

[Dev81a] Luc Devroye. How to reduce the average complexity of convex hull finding algorithms. Computers and Mathematics with Applications, 7(4):299–308, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900596.

Devroye:1981:ACS

[Dev81b] Luc Devroye. On the average complexity of some bucketing algorithms. Computers and Mathematics with Applications, 7(5):407–412, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901279.

Devroye:1981:CGR

[Dev81c] Luc Devroye. On the computer generation of random variables with a given characteristic function. Computers and Mathematics with Applications, 7(6):547–552, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900389.

Devroye:1982:CGR

[Dev82] Luc Devroye. On the computer generation of random convex hulls. Computers and Mathematics with Applications, 8 (1):1–13, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900359.

Devroye:1988:ESS

[Dev88] L. Devroye. The expected size of some graphs in computational geometry. Computers and Mathematics with Applications, 15 (1):53–64, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900715.

Dey:1985:INS

[Dey85] S. K. Dey. An implicit numerical solver for nonlinear hyperbolic partial differential equations. Computers and Mathematics with Applications, 11(1– 3):79–91, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901397.

Dey:1988:EBC

[Dey88] S. K. Dey. Effects of boundary conditions on numerical solution of implicit hyperbolic difference equations. Computers and Mathematics with Applications, 15(6– 8):619–622, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902854.

deFreitas:1989:SND

[dF89] L. de Freitas. 515 — a symmetric number in Dante. Computers and Mathematics with Applications, 17(4–6):887–897, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902654.

Du:1988:LDH

[DGL88] Q. Du, M. Gunzburger, and W. Layton. A low dispersion, high accuracy finite element method for first order hyperbolic systems in several space variables. Computers and Mathematics with Applications, 15(6– 8):447–457, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902726.

Diekmann:1986:SCS

[DHT86] O. Diekmann, H. J. A. M. Heijmans, and H. R. Thieme. On the stability of the cell-size distribution II: Time-periodic developmental rates. Computers and Mathematics with Applications, 12(4–5):491–512, April/May 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186901768.

Dierckx:1984:CLS

[Die84] Paul Dierckx. Computation of least-squares spline approximations to data over incomplete grids. Computers and Mathematics with Applications, 10 (3):283–289, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900567.

Diebold:1989:SDL

[Die89] F. X. Diebold. The solution of dynamic linear rational expectations models. Computers and Mathematics with Applications, 18(6–7):605–616, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901119.

Dixon:1989:SP

[Dix89] R. Dixon. Spiral phyllotaxis. Computers and Mathematics with Applications, 17(4–6):535– 538, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902496.

Dormand:1989:GEE

[DLMP89] J. R. Dormand, M. A. Lockyer, N. E. McGorrigan, and P. J. Prince. Global error estimation with Runge–Kutta triples. Computers and Mathematics with Applications, 18(9):835–846, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189901818.

delaSen:1985:MDA

[dlS85] M. de la Sen. Multirate digital adaptive control. Computers and Mathematics with Applications, 11(12):1193–1210, December 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122185901075.

delaSen:1987:DSR

[dlS87] M. de la Sen. A discrete stable robust MRAC design with an extra adaptation parameter. Computers and Mathematics with Applications, 14(6):429–435, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187900551.

delaSen:1987:SAC

[dlSGA87] M. de la Sen and C. Gil-Aguirrebeitia. A stable adaptive control design for discrete plants with unmodelled dynamics. Computers and Mathematics with Applications, 13(4): 341–350, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900010.

dAlessandro:1984:SMC

[dM84] P. d’Alessandro and M. Dalla Mora. Systems, memory, causality, evolution and recursive equations. Computers and Mathematics with Applications, 10 (1):61–69, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900865.

Doczi:1986:SUS

[Doc86] György Doczi. Seen and unseen symmetries: a picture essay. Computers and Mathematics with Applications, 12 (Part B)(1– 2):39–62, January/April 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901380.

Donnini:1986:VMS

[Don86] Roberto Donnini. The visualization of music: Symmetry and asymmetry. Computers and Mathematics with Applications, 12 (Part B)(1–2):435– 463, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901641.

Deer:1988:HOC

[DOS88] L. Deer, C. Outlaw, and D. Sarafyan. Higher order continuous solution of y′′ = f(x, y, y′) and y′′ = f(x, y) initial value problems. Computers and Mathematics with Applications, 16 (10–11):895–902, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218890199X.

Dormand:1986:RKT

[DP86] J. R. Dormand and P. J. Prince. Runge–Kutta triples. Computers and Mathematics with Applications, 12 (Part A)(9):1007– 1017, September 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900258.

Dormand:1987:RKN

[DP87] J. R. Dormand and P. J. Prince. Runge–Kutta–Nystrom triples. Computers and Mathematics with Applications, 13 (12):937–949, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900666.

Dubois:1988:ICI

[DP88] D. Dubois and H. Prade. Incomplete conjunctive information. Computers and Mathematics with Applications, 15 (10):797–810, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901174.

dePalma:1987:DMP

[dPLBA87] A. de Palma, C. Lefèvre, and M. Ben-Akiva. A dynamic model of peak period traffic flows and delays in a corridor. Computers and Mathematics with Applications, 14(3):201–223, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901532.

Dougalis:1981:ESF

[DS81] Vassilios A. Dougalis and Steven M.

Serbin. On the efficiency of some fully discrete Galerkin methods for second-order hyperbolic equations. Computers and Mathematics with Applications, 7(3): 261–279, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900869.

Driscoll:1985:RFA

[DS85] Michael A. Driscoll and R. P. Srivastav. Rational function approximations in the numerical solution of Cauchy-type singular integral equations. Computers and Mathematics with Applications, 11(9):949–965, September 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122185900999.

Drauschke:1988:EDU

[DT88] G. Drauschke and M. Tasche. Exact deconvolution using numbertheoretic transforms. Computers and Mathematics with Applications, 15(9):757–768, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901289.

Duff:1983:SAM

[Duf83] G. F. D. Duff. A special ADI model for the Laplace tidal equations. Computers and Mathematics with Applications, 9(3): 507–517, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900640.

Dunlop:1980:RCM

[Dun80] G. R. Dunlop. A rapid computational method for improvements to nearest neighbour interpolation. Computers and Mathematics with Applications, 6(3): 349–353, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900425.

Dunham:1986:HS

[Dun86]  Douglas Dunham.  Hyper-

bolic symmetry. Computers and Mathematics with Applications, 12 (Part B)(1–2):139– 153, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901471.

Duchting:1987:AMS

[DV87] W. Düchting and Th. Vogelsaenger. An approach of modelling and simulating the spread of heterogeneous tumor cells in a three-dimensional tissue segment. Computers and Mathematics with Applications, 14(9– 12):783–792, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902276.

deVilliers:1987:SPE

[dVWY87] R. de Villiers, C. J. Wright, and Y. Yavin. A stochastic pursuit– evasion differential game in 3-D: Optimal evasion strategies. Computers and Mathematics with Applications, 13(7):623–630, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901246.

DeJong:1983:UNL

[DW83] Gerald F. DeJong and David L. Waltz. Understanding novel language. Computers and Mathematics with Applications, 9(1): 131–147, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900123.

deWolff:1988:DNL

[dW88] P. M. de Wolff. Definition of Niggli’s lattice characters. Computers and Mathematics with Applications, 16(5–8):487–492, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902398.

Evans:1985:NEM

[EA85] D. J. Evans and A. R. Abdullah. A new explicit method for the diffusion-convention equation. Computers and Mathematics with Applications, 11(1–3): 145–154, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122185901439.

Ech-Charif:1983:OSF

[EC83] I. B. Ech-Charif. Overall smoothing of first and second derivatives of known functions in a set of points scattered through a domain in the plane (R^{2}). Computers and Mathematics with Applications, 9(4):595–608, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901177.

Enas:1986:CSP

[EC86] Gregory G. Enas and Sung C.

Choi. Choice of the smoothing parameter and efficiency of k-nearest neighbor classification. Computers and Mathematics with Applications, 12 (Part A) (2):235–244, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900763.

Edelen:1980:PCI

[Ede80] Dominic G. B. Edelen. Programs for computer implementation of the exterior calculus with applications to isovector calculations. Computers and Mathematics with Applications, 6(4):415–424, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122180900486.

Evans:1985:APR

[EG85] R. B. Evans and K. E. Gilbert. Acoustic propagation in a refracting ocean waveguide with an irregular interface. Computers and Mathematics with Applications, 11(7–8): 795–805, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901749.

Eller:1987:MTR

[EGZK87] J. Eller, I. Györi, M. Zöllei, and F. Krizsa. Modelling thrombopoiesis regulation — I: Model description and stimulation results. Computers and Mathematics with Applications, 14(9– 12):841–848, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187902331.

Esogbue:1988:OAL

[EL88] A. O. Esogbue and C. Y. Lee. Optimal analysis of large complex water resources conveyance systems via nonserial dynamic programming. Computers and Mathematics with Applications, 16(10–11):967–981, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122188902052.

Elderkin:1983:PMG

[Eld83] Richard H. Elderkin. Population models with globally agedependent dynamics: on computing the steady state. Computers and Mathematics with Applications, 9(3):371–376, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900512.

Elderkin:1987:BRT

[Eld87] R. Elderkin. Book review: Theory of nonlinear age-dependent population dynamics (Monographs and textbooks in pure and applied mathematics, vol. 89): By G. F. Webb. Dekker, New York (1985). vi + 294 pp. $65.00. Computers and Mathematics with Applications, 13(9– 11):913–915, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901738.

Ellacott:1986:SFM

[Ell86] S. W. Ellacott. A survey of Faber methods in numerical approximation. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1103–1107, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186902348.

Engel:1988:MPM

[Eng88a] P. Engel. Mathematical problems in modern crystallography. Computers and Mathematics with Applications, 16(5–8):425–436, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188902325.

Engersbach:1988:GCS

[Eng88b] N. Engersbach. Great circle stereographic trajectory representation. Computers and Mathematics with Applications, 15 (12):1019–1028, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218890137X.

Evans:1982:QDA

[EO82] D. J. Evans and S. O. Okolie. A quotient-difference algorithm for the determination of eigenvalues of periodic tridiagonal matrices. Computers and Mathematics with Applications, 8(2): 157–164, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122182900554.

Ehtamo:1989:CNB

[ERH89] H. Ehtamo, J. Ruusunen, and R. P. Hämäläinen. The computation of the Nash bargaining solution. Computers and Mathematics with Applications, 18(1– 3):189–193, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901363.

Eckhardt:1987:FMI

[ES87] U. Eckhardt and M. G. El Sheikh. A Fourier method for initial value problems with mixed boundary conditions. Computers and Mathematics with Applications, 14(3):189–199, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901520.

Evans:1988:GEM

[ES88a] D. J. Evans and M. S. Sahimi.

Group explicit methods for hyperbolic equations. Computers and Mathematics with Applications, 15(6–8):659–697, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218890288X.

Evans:1988:NTF

[ES88b] D. J. Evans and B. B. Sanugi.

A nonlinear trapezoidal formula for the solution of initial value problems. Computers and Mathematics with Applications, 15 (1):77–79, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900740.

Esogbue:1988:MCS

[Eso88] A. O. Esogbue. Mathematics, computers and systems: Tripartite interconnections. Computers and Mathematics with Applications, 16(10–11):773–777, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188901873.

Esposito:1988:GGT

[Esp88] C. Esposito. Graph graphics: Theory and practice. Computers and Mathematics with Applications, 15(4):247–253, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902088.

Evans:1981:ADI

[Eva81a] D. J. Evans. Alternating direction implicit preconditioning methods for self adjoint elliptic differential equations. Computers and Mathematics with Applications, 7(2):151–158, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181901140.

Evans:1981:MAP

[Eva81b] James R. Evans. The multicommodity assignment problem: a network aggregation heuristic. Computers and Mathematics with Applications, 7(2): 187–194, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901188.

Evans:1988:AGM

[Eva88]  D. J. Evans.  The age

Galerkin method for solving two-point boundary value problems. Computers and Mathematics with Applications, 16(12): 1045–1055, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902611.

Ellis:1984:PAS

[EW84] George H. Ellis and Layne T.

Watson. A parallel algorithm for simple roots for polynomials. Computers and Mathematics with Applications, 10(2): 107–121, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900427.

Esogbue:1986:HLC

[EW86] Augustine O. Esogbue and Nazir A. Warsi. A high-level computing algorithm for diverging and converging branch nonserial dynamic programming systems. Computers and Mathematics with Applications, 12 (Part A)(6):719–732, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690057X.

Esogbue:1988:HLD

[EW88] A. O. Esogbue and N. A. Warsi. A high level dynamic programming algorithm for processing nonserial looped systems. Computers and Mathematics with Applications, 16(10– 11):801–813, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901903.

Ewing:1980:ETS

[Ewi80] Richard E. Ewing. On efficient time-stepping methods for nonlinear partial differential equations. Computers and Mathematics with Applications, 6 (1):1–13, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900553.

Exoo:1988:SNC

[Exo88]  G. Exoo.  Some new com-

puter generated results in Ramsey graph theory. Computers and Mathematics with Applications, 15(4):255–257, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890209X.

ekabsZvilna:1986:CSS

[eZ86] ekabs Zvilna. Colored symmetries in space-time. Computers and Mathematics with Applications, 12(3–4):895–910, IN1, 911, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904335.

Fabricand:1989:SFM

[Fab89] B. P. Fabricand. Symmetry in free markets. Computers and Mathematics with Applications, 17(4–6):653–669, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990254X.

Farrell:1987:UOS

[Far87] P. A. Farrell. Uniform and optimal schemes for stiff initialvalue problems. Computers and Mathematics with Applications, 13(12):925–936, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900654.

Fatunla:1982:NMM

[Fat82] Simeon Ola. Fatunla. Nonlinear multistep methods for initial value problems. Computers and Mathematics with Applications, 8(3):231–239, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122182900463.

Fatunla:1984:OLM

[Fat84] Simeon O. Fatunla. One leg multistep method for second order differential equation. Computers and Mathematics with Applications, 10(1):1–4, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122184900804.

Fatunla:1985:OLH

[Fat85] Simeon Ola Fatunla. One leg hybrid formulas for second-order differential equations. Computers and Mathematics with Applications, 11(4):329–333, April 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900574.

Fatunla:1986:NTS

[Fat86]  S. O. Fatunla.  Numerical

treatment of singular initial value problems. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1109–1115, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690235X.

Franke:1988:UOD

[FBG88] R. Franke, E. Barker, and J. Goerss. The use of observed data for the initial-value problem in numerical weather prediction. Computers and Mathematics with Applications, 16(1– 2):169–184, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900302.

Friedman:1988:EPC

[FBS88] N. Friedman, A. Boyarsky, and M. Scarowsky. Ergodic properties of computer orbits for simple piecewise monotonic transformations. Computers and Mathematics with Applications, 15 (12):997–1006, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901356.

Feldman:1983:PFN

[FC83] Richard M. Feldman and Guy L.

Curry. A PDE formulation and numerical solution for a boll weevil-cotton crop model. Computers and Mathematics with Applications, 9(3):393–402, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122183900548.

Felsen:1985:NES

[Fel85a]  L. B. Felsen.  Numerically

efficient spectral representations for guided ocean acoustics. Computers and Mathematics with Applications, 11(7–8): 687–698, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901658.

Felsen:1985:GUD

[Fel85b] Leopold B. Felsen. Give us this day a stable code. Computers and Mathematics with Applications, 11(7–8):xi, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122185901609.

Ferguson:1982:SDG

[Fer82] Warren E. Ferguson, Jr. A simple derivation of Glassman’s general N Fast Fourier Transform. Computers and Mathematics with Applications, 8(6):401–411, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122182900165.

Fix:1980:NMA

[FG80] George J. Fix and Max D. Gunzburger. On numerical methods for acoustic problems. Computers and Mathematics with Applications, 6(2):265–278, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900346.

Fix:1983:FEA

[FGP83] George J. Fix, Max D. Gunzburger, and Janet S. Peterson. On finite element approximations of problems having inhomogeneous essential boundary conditions. Computers and Mathematics with Applications, 9(5):687–700, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901268.

Ford:1984:SAC

[FHS84] H. Ford, L. R. Hunt, and Renjeng Su. A simple algorithm for computing canonical forms. Computers and Mathematics with Applications, 10(4–5):315–326, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122184900592.

Fichtner:1986:NSG

[Fic86] Konrad Fichtner. Non-space-group symmetry in crystallography. Computers and Mathematics with Applications, 12(3– 4):751–762, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904219.

Fichtner:1988:ODS

[Fic88]  K. Fichtner.  Order-disorder

structures. Computers and Mathematics with Applications, 16(5– 8):469–477, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902374.

Finney:1989:LPS

[Fin89] J. L. Finney. Local pseudosymmetry in simple liquids. Computers and Mathematics with Applications, 17(1–3):341–354, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901673.

Fitzgibbon:1983:IBV

[Fit83] W. E. Fitzgibbon. Initial boundary value problems for the Carleman equation. Computers and Mathematics with Applications, 9(3):519–525, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900652.

Fitzgibbon:1987:HSN

[Fit87] W. E. Fitzgibbon. A hyperbolic system from a neural transmission model. Computers and Mathematics with Applications, 13(9–11):767–769, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901623.

Farmer:1985:LMZ

[FL85] M. R. Farmer and G. Loizou.

Locating multiple zeros interactively. Computers and Mathematics with Applications, 11 (6):595–603, June 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900422.

Fletcher:1988:CMS

[Fle88] C. A. J. Fletcher. Computational modelling of severe gradients with mass operator constructions. Computers and Mathematics with Applications, 16 (1–2):31–39, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900223.

Fleurent:1989:SPSa

[Fle89a] G. M. Fleurent. Symmetry and polyhedral stellation — Ia. Computers and Mathematics with Applications, 17(1–3):167–175, ????

1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901570.

Fleurent:1989:SPSb

[Fle89b] G. M. Fleurent. Symmetry and polyhedral stellation — Ib. Computers and Mathematics with Applications, 17(1–3):177–193, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901582.

Fix:1982:PHF

[FLL82] G. J. Fix, G. Liang, and D. N. Lee. Penalty-hybrid finite element method. Computers and Mathematics with Applications, 8(5):393–399, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900335.

Faber:1986:ABS

[FMWW86] V. Faber, Thomas A. Manteuffel, Andrew B. White, Jr., and G. Milton Wing. Asymptotic behavior of singular values and singular functions of certain convolution operators. Computers and Mathematics with Applications, 12 (Part A)(6):733–747, June 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186900581.

Fujita:1986:NSA

[FN86] Yasuhiro Fujita and Makiko Nisio. Nonlinear semigroups associated with optimal stopping of controlled diffusions under partial observation. Computers and Mathematics with Applications, 12 (Part A)(6):749–760, June 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186900593.

Foley:1987:IAD

[Fol87] T. A. Foley. Interpolation and approximation of 3-D and 4-D scattered data. Computers and Mathematics with Applications, 13(8):711–740, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900435.

Fomenko:1986:SSF

[Fom86] A. T. Fomenko. Symmetries of soap films. Computers and Mathematics with Applications, 12(3– 4):825–834, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690427X.

Fomenko:1989:VHS

[Fom89] A. T. Fomenko. Visual and hidden symmetry in geometry. Computers and Mathematics with Applications, 17(1–3):301–320, ????

1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901648.

Francis:1980:CGI

[Fra80] N. D. Francis. Computation of generalised inverse of a special class of matrices. Computers and Mathematics with Applications, 6(4):411–413, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900474.

Franke:1982:SIS

[Fra82] Richard Franke. Smooth interpolation of scattered data by local thin plate splines. Computers and Mathematics with Applications, 8(4):273–281, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900098.

Franca:1989:ACS

[Fra89] L. P. Franca. An algorithm to compute the square root of a 3×3 positive definite matrix. Computers and Mathematics with Applications, 18(5):459–466, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218990240X.

Franke:1980:EMS

[FS80] Richard Franke and David Salinas. An efficient method for solving stiff transient field problems arising from FEM formulations. Computers and Mathematics with Applications, 6 (1):15–21, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900565.

Frauenthal:1983:SBR

[FS83] James C. Frauenthal and Kenneth E. Swick. Stability of biochemical reaction tanks. Computers and Mathematics with Applications, 9(3):499–506, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122183900639.

Fix:1985:TDM

[FS85] George J. Fix and Manil Suri. Three-dimensional massconserving elements for compressible flows. Computers and Mathematics with Applications, 11(7–8):765–776, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122185901713.

Fusaro:1987:MM

[Fus87] B. A. Fusaro. Mathematics and the microcomputer. Computers and Mathematics with Applications, 14(1):1–2, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901763.

Franke:1987:MIH

[FWPT87] H. Franke, H. E. Wichmann, C. S. Potten, and L. Todd. Modelling of the influence of ^{3} HTdR on cell kinetics in mouse epidermis. Computers and Mathematics with Applications, 14(9– 12):741–749, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902239.

Gaidov:1986:NES

[Gai86] Svetoslav D. Gaidov. Nashequilibrium in stochastic differential games. Computers and Mathematics with Applications, 12 (Part A)(6):761–768, June 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218690060X.

Gargantini:1980:PSR

[Gar80]  Irene Gargantini.  Parallel

square-root iterations for multiple roots. Computers and Mathematics with Applications, 6(3): 279–288, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122180900358.

Gargantini:1982:DCR

[Gar82] Irene Gargantini. Detection of connectivity for regions represented by linear quadtrees. Computers and Mathematics with Applications, 8(4):319–327, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122182900141.

See comments [Abe84].

Gavroglu:1989:SNE

[Gav89] K. Gavroğlu. Symmetries and the notion of elementarity in high energy physics. Computers and Mathematics with Applications, 17(1–3):457–465, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901740.

Glasgow:1985:PLA

[GB85] J. Glasgow and R. Browse. Programming languages for artificial intelligence. Computers and Mathematics with Applications, 11(5):431–448, May 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900495.

Gora:1988:WCL

[GB88] P. Góra and A. Boyarsky. Why computers like Lebesgue measure. Computers and Mathematics with Applications, 16 (4):321–329, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901484.

Gyori:1987:MTR

[GE87] I. Györi and J. Eller. Modelling thrombopoiesis regulation — II: Mathematical investigation of the model. Computers and Mathematics with Applications, 14(9–12):849–859, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902343.

Gutman:1987:SLP

[GEG87] S. Gutman, M. Esh, and M. Gefen. Simple linear pursuit– evasion games. Computers and Mathematics with Applications, 13(1–3):83–95, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900952.

Geiszler:1981:NPS

[Gei81] Carl Geiszler. A numerical procedure for the selection of the constant interarrival time to a single server queue. Computers and Mathematics with Applications, 7(6):537–546, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900377.

Gersting:1980:NME

[Ger80] John M. Gersting, Jr. Numerical methods for eigensystems: the Orr–Sommerfeld problem as an initial value problem. Computers and Mathematics with Applications, 6(2):167–174, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900267.

Gerasoulis:1982:UPQ

[Ger82] Apostolos Gerasoulis. The use of piecewise quadratic polynomials for the solution of singular integral equations of Cauchy type. Computers and Mathematics with Applications, 8 (1):15–22, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900360.

Gerasoulis:1986:SVD

[Ger86]  Apostolos Gerasoulis.  The

singular value decomposition of the Gauss–Chebyshev and Lobatto–Chebyshev methods for Cauchy Singular Integral Equations. Computers and Mathematics with Applications, 12 (Part A) (8):895–907, August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812218690026X.

Gerdes:1989:REL

[Ger89] P. Gerdes. Reconstruction and extension of lost symmetries: Examples from the Tamil of south India. Computers and Mathematics with Applications, 17(4– 6):791–813, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902617.

Gordon:1987:MGS

[GG87] F. S. Gordon and S. P. Gordon. Microcomputer graphics for statistical education. Computers and Mathematics with Applications, 14(1):71–80, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901829.

Gaglio:1987:ESC

[GGR^{+}87] S. Gaglio, M. Genovesi, C. Ruggiero, G. Spinelli, C. Nicolini, G. Bonadonna, and P. Valagussa. Expert systems for cancer chemotherapy. Computers and Mathematics with Applications, 14(9–12):793–802, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187902288.

Grunbaum:1986:SMO

[GGS86] Branko Grünbaum, Zdenka Grünbaum, and G. C. Shepard. Symmetry in moorish and other ornaments. Computers and Mathematics with Applications, 12(3–4):641–653, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186904165.

Graham:1988:CPG

[GH88] Niall Graham and Frank Harary.

Covering and packing in graphs — V: Mispacking subcubes in hypercubes. Computers and Mathematics with Applications, 15 (4):267–270, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902118.

Galanis:1989:RBJ

[GHN89] S. Galanis, A. Hadjidimos, and D. Noutsos. The relationship between the Jacobi and the successive overrelaxation (SOR) matrices of a k-cyclic matrix. Computers and Mathematics with Applications, 17(10):1351–1357, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189900187.

Giger:1986:M

[Gig86] Hans Giger. Moirés. Computers and Mathematics with Applications, 12 (Part B)(1–2):329– 361, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901586.

Ginsberg:1980:MFP

[Gin80] Myron Ginsberg. Monitoring floating-point error propagation in scientific computation. Computers and Mathematics with Applications, 6(1):23–43, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900577.

Gipson:1985:BET

[Gip85] G. Steven Gipson. On bounding the errors in three Runge–Kutta type formulations. Computers and Mathematics with Applications, 11(11):1115–1125, November 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/0898122185901890.

Gipson:1987:EDS

[Gip87] G. S. Gipson. An effective difference scheme for quasilinear diffuion. Computers and Mathematics with Applications, 13 (7):635–646, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218790126X.

Grechushnikov:1988:COA

[GK88] B. N. Grechushnikov and A. F.

Konstantinova. Crystal optics of absorbing and gyrotropic media. Computers and Mathematics with Applications, 16(5– 8):637–655, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902520.

Glicksman:1985:SCV

[Gli85]  Jay Glicksman.  Symbiosis

in computational vision systems. Computers and Mathematics with Applications, 11 (5):505–515, May 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900537.

Goldberg:1988:PAB

[GM88] M. Goldberg and Z. Miller. A parallel algorithm for bisection width in trees. Computers and Mathematics with Applications, 15(4):259–266, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902106.

Gunzburger:1982:CMF

[GNP82] Max D. Gunzburger, R. A. Nicolaides, and Janet S. Peterson. On conforming mixed finite element methods for incompressible viscous flow problems. Computers and Mathematics with Applications, 8(3):167–179, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900402.

Grimmer:1985:LVE

[GP85] R. Grimmer and J. Prüss. On linear Volterra equations in Banach spaces. Computers and Mathematics with Applications, 11 (1–3):189–205, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122185901464.

Ghose:1989:MDG

[GP89] D. Ghose and U. R. Prasad. Multicriterion differential games with applications to combat problems. Computers and Mathematics with Applications, 18(1– 3):117–126, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901296.

Geppini:1988:AQC

[GR88] M. Geppini and F. Romani. Automatic quadrature on Chebyshev points. Computers and Mathematics with Applications, 16(3):185–193, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122188901794.

Gray:1986:SNF

[Gra86] Norman H. Gray. Symmetry in a natural fracture pattern: the origin of columnar joint networks. Computers and Mathematics with Applications, 12(3– 4):531–545, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904098.

Graham:1987:BRR

[Gra87]  J. H. Graham.  Book re-

view: Robot analysis and control: By Haruhiko Asada and Jean-Jacque E. Slotine. Wiley– Interscience, New York (1986). 266 pp. Computers and Mathematics with Applications, 14(9– 12):948–949, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902422.

Greenberg:1987:BRD

[Gre87a] R. A. Greenberg. Book review:

The design and analysis of clinical experiments: By Joseph L. Fleiss. Wiley, New York (1986). 432 pp. Computers and Mathematics with Applications, 14(9– 12):950–951, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187902458.

Greenspan:1987:QMC

[Gre87b] D. Greenspan. Quasimolecular modeling of cavity flow. Computers and Mathematics with Applications, 14(4):239–248, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187901313.

Greenspan:1988:MCS

[Gre88a] D. Greenspan. Mechanisms of capillarity via supercomputer simulation. Computers and Mathematics with Applications, 16(4):331–340, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901496.

Greenspan:1988:QCV

[Gre88b] D. Greenspan. Quasimolecular channel and vortex street modeling on a supercomputer. Computers and Mathematics with Applications, 15(2):141–151, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188900831.

Greenspan:1989:PSB

[Gre89] D. Greenspan. Particle simulation of biological sorting on a supercomputer. Computers and Mathematics with Applications, 18(9):823–834, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901806.

Gripenberg:1983:SAD

[Gri83] Gustaf Gripenberg. Stability analysis of a distributed parameter model for the growth of micro-organisms. Computers and Mathematics with Applications, 9(3):431–442, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900573.

Grossman:1987:BRS

[Gro87] Z. Grossman. Book review:

System Theory in Immunology (Lecture Notes in Biomathematics): Vol. 32. Edited by S. Levin Springer, New York (1979). 273 pp. Computers and Mathematics with Applications, 14(9–12):948, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902410.

Gropp:1988:LUM

[Gro88]  W. D. Gropp.  Local uni-

form mesh refinement on looselycoupled parallel processors. Computers and Mathematics with Applications, 15(5):375– 387, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901599.

Guram:1980:SFM

[GS80] G. S. Guram and A. C. Smith. Stagnation flows of micropolar fluids with strong and weak interactions. Computers and Mathematics with Applications, 6(2): 213–233, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900309.

Gatto:1981:NEM

[GS81] M. A. Gatto and J. B. Seery. Numerical evaluation of the modified Bessel functions I and K. Computers and Mathematics with Applications, 7(3): 203–209, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900808.

Gavezzotti:1986:SPS

[GS86] A. Gavezzotti and M. Simonetta. On the symmetry of periodic structures in two dimensions. Computers and Mathematics with Applications, 12 (Part B)(1–2):465–476, January/April 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186901653.

Galperin:1987:PED

[GS87a] E. A. Galperin and J. M.

Skowronski. pursuit–evasion differential games with uncertainties in dynamics. Computers and Mathematics with Applications, 13(1–3):13–35, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900915.

Gerasoulis:1987:SGC

[GS87b] A. Gerasoulis and R. P. Srivastav. The stability of the Gauss– Chebyshev method for Cauchy singular integral equations. Computers and Mathematics with Applications, 14(2):81–90, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901398.

Goyal:1987:CRT

[GS87c] S. Goyal and S. M. Serbin. A class of Rosenbrock-type schemes for second-order nonlinear systems of ordinary differential equations. Computers and Mathematics with Applications, 13 (4):351–362, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900022.

Grimm:1989:OLC

[GS89] W. Grimm and J. M. Schaeffer.

Optimal launch conditions on the “no-escape envelope”. Computers and Mathematics with Applications, 18(1–3):45–59, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901235.

Green:1988:IPC

[GT88] H. S. Green and T. Triffet. Information processing by the cortex. Computers and Mathematics with Applications, 15(9): 743–756, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901277.

Gupta:1986:CRM

[Gup86] A. K. Gupta. On a classification rule for multiple measurements. Computers and Mathematics with Applications, 12 (Part A)(2):301–308, February 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186900829.

Gyori:1988:MLS

[Gyö88a] I. Györi. The method of lines for the solutions of some nonlinear partial differential equations. Computers and Mathematics with Applications, 15(6– 8):635–658, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122188902878.

Gyori:1988:TAT

[Gyö88b] I. Györi. Two approximation techniques for functional differential equations. Computers and Mathematics with Applications, 16(3):195–214, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901800.

Gyori:1989:OCR

[Gyö89] I. Györi. Oscillation and comparison results in neutral differential equations and their applications to the delay logistic equation. Computers and Mathematics with Applications, 18(10– 11):893–906, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900072.

Galperin:1989:IGO

[GZ89] E. A. Galperin and Q. Zheng.

Integral global optimization method for differential games with application to pursuit– evasion games. Computers and Mathematics with Applications, 18(1–3):209–243, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901387.

Ha:1989:SPY

[Ha89] Y.-M. Ha. Symmetry in pictures by young Chinese children. Computers and Mathematics with Applications, 17(4–6):921–930, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189902691.

Haake:1989:RSJ

[Haa89] A. Haake. The role of symmetry in Javanese batik patterns. Computers and Mathematics with Applications, 17(4–6):815–826, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189902629.

Haddad:1981:NAN

[Had81] Y. M. Haddad. Numerical analysis in nonlinear viscoelasticity. Computers and Mathematics with Applications, 7(2): 139–145, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181901127.

Hadley:1985:SNL

[Had85] Robert F. Hadley. Shadow: a natural language query analyzer. Computers and Mathematics with Applications, 11 (5):481–504, May 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900525.

Hager:1987:BD

[Hag87] W. W. Hager. Bidiagonalization and diagonalization. Computers and Mathematics with Applications, 14(7):561–572, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900514.

Haghighi:1988:CPN

[Hag88] M. Haghighi. Computation of prime numbers by using a probabilistic algorithm. Computers and Mathematics with Applications, 15(11):939–942, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900375.

Hager:1989:DBB

[Hag89] W. W. Hager. A derivative-based bracketing scheme for univariate minimization and the conjugate gradient method. Computers and Mathematics with Applications, 18(9):779–795, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901776.

Hand:1986:ECS

[Han86]  D. J. Hand.  Estimating

class sizes by adjusting fallible classifier results. Computers and Mathematics with Applications, 12 (Part A)(2): 289–299, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900817.

Hand:1987:SLO

[Han87]  D. J. Hand.  A shrunken

leaving-one-out estimator of error rate. Computers and Mathematics with Applications, 14 (3):161–167, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901489.

Hannaby:1988:MMM

[Han88]  S. A. Hannaby.  A map-

ping method for mesh generation. Computers and Mathematics with Applications, 16 (9):727–735, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900089.

Hargittai:1986:LP

[Har86a] István Hargittai. Limits of perfection. Computers and Mathematics with Applications, 12 (Part B)(1–2):1– 17, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122186901355.

Hargittai:1986:P

[Har86b] István Hargittai. Preface. Computers and Mathematics with Applications, 12 (Part B)(1–2):xiii– xv, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901343.

Harries:1986:SNR

[Har86c]  John G. Harries.  Symme-

try and notation: Regularity and symmetry in notated computer graphics. Computers and Mathematics with Applications, 12 (Part B)(1–2):303– 314, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901562.

Hartley:1986:CMM

[Har86d] Michael J. Hartley. Calibration of macroeconomic models with incomplete data — a systems approach. Computers and Mathematics with Applications, 12 (Part A)(6):769–775, June 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900611.

Harary:1988:CGC

[Har88a] Frank Harary. Cubical graphs and cubical dimensions. Computers and Mathematics with Applications, 15(4):271–275, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218890212X.

Hargittai:1988:SML

[Har88b] I. Hargittai. Symmetries in Moscow and leningrad. Computers and Mathematics with Applications, 16(5–8):663–669, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902544.

Hartmann:1988:SEP

[Har88c] E. Hartmann. Symmetry in an equilibrium position. Computers and Mathematics with Applications, 16(5–8):465–468, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902362.

Hargittai:1989:JS

[Har89a] I. Hargittai. The joy of symmetry. Computers and Mathematics with Applications, 17(4– 6):1067–1072, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902769.

Harries:1989:SMA

[Har89b] J. G. Harries. Symmetry in the movements of t’ai chi chuan. Computers and Mathematics with Applications, 17(4– 6):827–835, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902630.

Hauptman:1988:RCS

[Hau88] H. Hauptman. The role of crystallographic symmetry in the direct methods of X-ray crystallography. Computers and Mathematics with Applications, 16(5– 8):385–396, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902295.

Havens:1983:RMS

[Hav83] William S. Havens. Recognition mechanisms for schema-based knowledge representations. Computers and Mathematics with Applications, 9(1):185–199, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122183900147.

Hayes:1980:ADC

[Hay80] Linda J. Hayes. An alternatingdirection collocation method for finite element approximations on rectangles. Computers and Mathematics with Applications, 6(1):45–50, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900589.

Hargittai:1986:VMM

[HC86] I. Hargittai and B. Chamberland. The VSEPR model of molecular geometry. Computers and Mathematics with Applications, 12(3– 4):1021–1038, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904384.

Havenner:1989:STT

[HC89] A. M. Havenner and K. R. Criddle. System theoretic time series: an application to inventories and prices of California range cattle. Computers and Mathematics with Applications, 17(8– 9):1177–1187, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900886.

Hadeler:1983:NHP

[HD83] K. P. Hadeler and K. Dietz. Nonlinear hyperbolic partial differential equations for the dynamics of parasite populations. Computers and Mathematics with Applications, 9(3):415–430, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900561.

He:1987:PSD

[He87] J. He. Practical stability of discontinuous large-scale systems. Computers and Mathematics with Applications, 14 (2):119–125, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901416.

Herrmann:1986:ASC

[Her86]  Heinz Herrmann.  Asymme-

try and symmetry in cellular organization. Computers and Mathematics with Applications, 12 (Part B)(1–2):155– 167, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901483.

Herndon:1988:GCD

[Her88] W. C. Herndon. Graph codes and a definition of structural similarity. Computers and Mathematics with Applications, 15 (4):303–309, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902167.

Hendler:1986:IPH

[HF86] Lawrence Hendler and Stephen S. Friedland. Image plane holography for holographic presentation of a three-dimensional data base. Computers and Mathematics with Applications, 12 (Part A)(6):777–784, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900623.

Hand:1987:ERE

[HF87] D. J. Hand and G. M. Fitzmaurice. Error rate estimation by mixture decomposition. Computers and Mathematics with Applications, 14(7):573–578, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187900526.

Head-Gordon:1986:PAS

[HGP86] M. Head-Gordon and P. Piela. Parallel algorithms for solving linear equations using Givens transformations. Computers and Mathematics with Applications, 12 (Part A)(9):987– 990, September 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900234.

Harary:1988:NGG

[HH88] Frank Harary and D. Frank Hsu. Node-graceful graphs. Computers and Mathematics with Applications, 15(4):291–298, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902143.

Holden:1988:NMF

[HHHK88] H. Holden, L. Holden, and R. Høegh-Krohn. A numerical method for first order nonlinear scalar conservation laws in one-dimension. Computers and Mathematics with Applications, 15(6–8):595–602, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902829.

Harn:1989:PAU

[HHL89] L. Harn, D. Huang, and C. S.

Laih. Password authentication using public-key cryptography. Computers and Mathematics with Applications, 18(12): 1001–1017, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990028X.

Harary:1988:STH

[HHW88] Frank Harary, John P. Hayes, and Horng-Jyh Wu. A survey of the theory of hypercube graphs. Computers and Mathematics with Applications, 15 (4):277–289, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902131.

Hill:1989:MS

[Hil89] A. Hill. Modern symmetry. Computers and Mathematics with Applications, 17(4–6):709–713, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189902575.

Hatziavramidis:1985:ICE

[HK85] Dimitri Hatziavramidis and Hwar-Ching Ku. An integral Chebyshev expansion method for boundary-value problems of O.D.E. type. Computers and Mathematics with Applications, 11(6):581–586, June 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900409.

Hwang:1989:SSS

[HKW89] T.-W. Hwang, T.-H. Kuo, and C.-A. Wang. Similarity solutions for surface-tension driven flows in a slot with an insulated bottom. Computers and Mathematics with Applications, 17(12): 1573–1586, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900576.

Huang:1986:CPS

[HL86] Luping Huang and Tai-Ping Liu.

A conservative, piecewise-steady difference scheme for transonic nozzle flow. Computers and Mathematics with Applications, 12(4–5):377–388, April/May 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186901707.

Harary:1988:STW

[HL88] Frank Harary and Martin Lewinter. The starlike trees which span a hypercube. Computers and Mathematics with Applications, 15(4):299–302, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902155.

Hill:1989:CSP

[HL89] A. Hill and D. G. Larman.

A class of symmetric polytopes. Computers and Mathematics with Applications, 17(1– 3):133–136, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901545.

Hammel:1989:OOM

[HLHG89] S. E. Hammel, P. T. Liu, E. J.

Hilliard, and K. F. Gong. Optimal observer motion for localization with bearing measurements. Computers and Mathematics with Applications, 18(1– 3):171–180, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990134X.

Hosoya:1986:MSG

[Hos86] Haruo Hosoya. Matching and symmetry of graphs. Computers and Mathematics with Applications, 12 (Part B)(1–2):271– 290, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901549.

Hilton:1986:SM

[HP86] Peter Hilton and Jean Pedersen. Symmetry in mathematics. Computers and Mathematics with Applications, 12 (Part B)(1–2):315– 328, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901574.

Hickernell:1987:DCM

[HP87] F. J. Hickernell and W. Proskurowski. Design of a calculus microcomputer laboratory. Computers and Mathematics with Applications, 14(1):3–24, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122187901775.

Harary:1988:P

[HP88] Frank Harary and Keith Phillips.

Preface. Computers and Mathematics with Applications, 15 (4):vii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902064.

Hilton:1989:DDD

[HP89a] P. Hilton and J. Pedersen. Duality and the Descartes deficiency. Computers and Mathematics with Applications, 17 (1–3):73–88, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901491.

Hilton:1989:EBC

[HP89b] P. Hilton and J. Pedersen. Extending the binomial coefficients to preserve symmetry and pattern. Computers and Mathematics with Applications, 17(1– 3):89–102, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901508.

Heymann:1989:OSE

[HRA89] M. Heymann, N. Rajan, and M. Ardema. Optimal strategies in event-constrained differential games. Computers and Mathematics with Applications, 18(1– 3):127–131, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901302.

Haselgrove:1981:CBC

[HS81] Jenifer Haselgrove and S. M. Selim. A comparison of bounds for the chromatic number of a graph. Computers and Mathematics with Applications, 7(3): 259–260, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900857.

Hu:1987:AHS

[HS87] S. S. Hu and X. M. Shao. Adaptive hybridized spline differentiators for numerical solution of the advection equation. Computers and Mathematics with Applications, 13(9–11):839–849, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901672.

Hu:1988:AHS

[HS88] S. S. Hu and X. M. Shao. Adaptive hybridized spline differentiators for numerical solution of the advection equation. Computers and Mathematics with Applications, 15(6–8):525–534, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902787.

Hamner:1987:CSG

[HTY87] C. G. Hamner, D. W. Turner, and D. M. Young. Comparisons of several graphical methods for representing multivariate data. Computers and Mathematics with Applications, 13 (7):647–655, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901271.

Huang:1989:SRC

[Hua89] C. G. Huang. Several rigorous counterexamples about Saint-Venant’s principle. Computers and Mathematics with Applications, 18(8):729–738, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902290.

Hargittai:1988:P

[HV88] I. Hargittai and B. K. Vainshtein. Preface. Computers and Mathematics with Applications, 16(5–8):ix, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902234.

Hsueh:1989:SLS

[HWL89] C. T. Hsueh, J. F. Wang, and J. Y. Lee. Solving linear systems of determinant frequently zero over finite field GF (2). Computers and Mathematics with Applications, 18(4):373–379, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189902228.

Huang:1988:NST

[HWYL88] T.-C. Huang, J.-F. Wang, C.- S. Yang, and J.-Y. Lee. The number of spanning trees of the generalized Boolean n-cube network. Computers and Mathematics with Applications, 16 (9):715–725, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900077.

Haussmann:1986:AZP

[HZ86] W. Haussmann and K. Zeller. Approximate Zolotarev polynomials. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1133–1140, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186902373.

Ibe:1988:ALR

[IE88] O. C. Ibe and A. O. Esogbue. Application of Lagrangian relaxation to computer network control. Computers and Mathematics with Applications, 16(10– 11):845–850, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901940.

Iskandar:1980:NSE

[IJ80] Labib Iskandar and Padam C.

Jain. Numerical solution of equations having inelastic solitary wave interaction. Computers and Mathematics with Applications, 6(4):373–383, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900449.

Imado:1989:FEB

[IM89a] F. Imado and S. Miwa. Fighter evasive boundaries against missiles. Computers and Mathematics with Applications, 18 (1–3):1–14, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990120X.

Iskandar:1989:NSA

[IM89b] L. Iskandar and M. Sh. El-Deen Mohamedein. Numerical study of an analogue of the Boussinesq equation. Computers and Mathematics with Applications, 18(5):467–474, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902411.

Isser:1989:ALS

[IR89] A. Isser and A. Rosen. Applying the lifting surface method to the blade tip of a hovering helicopter. Computers and Mathematics with Applications, 17(11): 1455–1466, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900783.

Iraniparast:1987:QBS

[Ira87] N. Iraniparast. Qualitative behaviour of solutions of the Goursat problem for hyperbolic differential equations. Computers and Mathematics with Applications, 13(9–11):889–900, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901714.

Israel:1983:INF

[Isr83] David J. Israel. Interpreting network formalisms. Computers and Mathematics with Applications, 9(1):1–13, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900032.

Iwaarden:1987:CTT

[Iwa87] J. L. Van Iwaarden. The computer as a teaching tool in ordinary differential equations. Computers and Mathematics with Applications, 14(1):25–32, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901787.

Jha:1987:TEI

[JA87] N. K. Jha and J. A. Abraham. Techniques for efficiently implementing totally selfchecking checkers in MOS technology. Computers and Mathematics with Applications, 13(5– 6):555–566, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900836.

JaJa:1985:LBM

[JáJ85] Joseph JáJá. Lower bounds on monotone arithmetic circuits with restricted depths. Computers and Mathematics with Applications, 11(12):1155– 1164, December 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901038.

Jarmark:1987:CLC

[Jär87] B. Järmark. On closed-loop controls in pursuit–evasion. Computers and Mathematics with Applications, 13(1–3):157–166, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901015.

Jarmark:1989:TAA

[Jär89] B. Järmark. Two aggressive aircraft in a realistic short-range combat as a differential game study. Computers and Mathematics with Applications, 18(1– 3):101–105, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901272.

Jaynes:1992:CTA

[Jay92] E. T. Jaynes. Commentary on two articles by C. A. Los. Computers and Mathematics with Applications, 24(8–9):267–273, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812219290204U.

See [Los89b, Los89a].

Jensen:1986:CSP

[Jen86] William B. Jensen. Classification, symmetry and the periodic table. Computers and Mathematics with Applications, 12 (Part B)(1–2):487–510, January/April 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186901677.

Jen:1983:SSI

[JJS83] Erica Jen, Erica Jen, and R. P.

Srivastav. Solving singular integral equations using Gaussian quadrature and overdetermined systems. Computers and Mathematics with Applications, 9(4): 625–632, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901207.

Jones:1982:NSM

[JM82] H. M. Jones and S. McKee. The numerical stability of multistep methods for convolution Volterra integral equations: Nonsingular equations. Computers and Mathematics with Applications, 8(4):291–303, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900116.

Jena:1984:MCF

[JM84] S. K. Jena and M. N. Mathur.

Mixed convection flow of a micropolar fluid from an isothermal vertical plate. Computers and Mathematics with Applications, 10(3):291–304, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900579.

Jain:1989:OFS

[JM89] P. Jain and S. Maheshwari. Optimal feedback stabilization policy with asymmetric loss functions. Computers and Mathematics with Applications, 18(6– 7):651–662, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901156.

Johnson:1986:MMS

[Joh86]  Lee M. Johnson.  Milton’s

mathematical symbol of theodicy. Computers and Mathematics with Applications, 12(3–4): 617–627, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186904141.

Jazcilevich:1989:SCE

[JT89] A. Jazcilevich and R. P. Tewarson. A stability criterion for explicit boundary value Runge– Kutta methods. Computers and Mathematics with Applications, 17(7):1077–1083, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900370.

Julian:1988:BMP

[Jul88] William H. Julian. Banyan models for the parallel implementation of the fast Fourier transform algorithm. Computers and Mathematics with Applications, 15(4):311–317, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902179.

Jurkat:1983:EAA

[Jur83] M. Peter Jurkat. Error analysis of Aitken’s Δ^{2} process. Computers and Mathematics with Applications, 9(2):317–322, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901347.

Jha:1988:UED

[JV88] N. K. Jha and M. B. Vora. A t-unidirectional error-detecting systematic code. Computers and Mathematics with Applications, 16(9):705–714, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900065.

Kimmel:1989:SDE

[KA89a] M. Kimmel and O. Arino. A system of differential equations modeling the G _{1} phase of the cell cycle. Computers and Mathematics with Applications, 18(10– 11):907–917, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900084.

Kohn:1989:FSA

[KA89b] R. Kohn and C. F. Ansley. Filtering and smoothing algorithms for state space models. Computers and Mathematics with Applications, 18(6–7):515–528, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189901041.

Kamel:1987:BRP

[Kam87] K. Kamel. Book review: Proceedings ISMIII’82: First IEEE Computer Society International Symposium on Medical Imaging and Image Interpretation: Berlin, IEEE Computer Society Press (1982). 610 pp. Computers and Mathematics with Applications, 14(9–12):950, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187902446.

Kansa:1988:SCA

[Kan88] E. J. Kansa. Shock computations with adaptive mesh refinement and moving grids. Computers and Mathematics with Applications, 15(6–8):623–634, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188902866.

Kappraff:1986:CMD

[Kap86a] Jay Kappraff. A course in the mathematics of design.

Computers and Mathematics with Applications, 12(3–4):913– 948, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904347.

Kappraff:1986:GCS

[Kap86b] Jay Kappraff. The geometry of coastlines: a study in fractals. Computers and Mathematics with Applications, 12(3–4): 655–671, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904177.

Karplus:1981:PIT

[Kar81]  Kevin Karplus.  A proof

of the isomorphism of wxyztransformals and 2 × 2 integer matrices. Computers and Mathematics with Applications, 7(5):425–430, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901292.

Karakashian:1986:GSS

[Kar86] Ohannes A. Karakashian. Galerkin single-step methods for secondorder hyperbolic equations. Computers and Mathematics with Applications, 12(4–5):605– 613, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901847.

Karakashian:1987:RKN

[Kar87] O. A. Karakashian. Runge– Kutta–Nyström methods for hyperbolic problems with timedependent coefficients. Computers and Mathematics with Applications, 13(9–11):785–800, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187901647.

Karle:1988:OSU

[Kar88] J. Karle. Origin specification by use of phases. Computers and Mathematics with Applications, 16(5–8):397–406, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902301.

Katz:1980:OLS

[KC80] I. Norman Katz and Leon Cooper. Optimal location on a sphere. Computers and Mathematics with Applications, 6(2): 175–196, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900279.

Kerr:1989:SCT

[Ker89] S. Kerr. Symmetry in Christian time and space. Computers and Mathematics with Applications, 17(4–6):899–906, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902666.

Khater:1989:ASN

[KESC89] A. H. Khater, M. F. El-Sabbagh, and D. K. Callebaut. Analytical solutions for nonlinear magnetohydrodynamic atmospheres. Computers and Mathematics with Applications, 17(10): 1379–1384, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900217.

Kharab:1986:NSF

[Kha86] A. Kharab. A numerical solution of a free-boundary problem on the unsaturated flow of liquids in porous media. Computers and Mathematics with Applications, 12 (Part A)(12):1193– 1200, December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900040.

Khaliq:1989:PCS

[Kha89] A. Q. M. Khaliq. A predictor– corrector scheme for fourth order parabolic partial differential equations. Computers and Mathematics with Applications, 17(12):1563–1566, ???? 1989.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900552.

Kittredge:1983:SPT

[Kit83] Richard I. Kittredge. Semantic processing of texts in restricted sublanguages. Computers and Mathematics with Applications, 9(1):45–58, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900068.

Kitagawa:1989:NGS

[Kit89] G. Kitagawa. Non-Gaussian seasonal adjustment. Computers and Mathematics with Applications, 18(6–7):503–514, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812218990103X.

Kagiwada:1986:WRB

[KK86] Harriet H. Kagiwada and Robert E. Kalaba. The work of Richard Bellman — I. Optimization. Computers and Mathematics with Applications, 12 (Part A) (6):785–790, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900635.

Krishnamoorthy:1987:FDI

[KK87] M. S. Krishnamoorthy and B. Krishnamurthy. Fault diameter of interconnection networks. Computers and Mathematics with Applications, 13(5– 6):577–582, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218790085X.

Kawahara:1988:FEM

[KKK88] M. Kawahara, T. Kodama, and M. Kinoshita. Finite element method for tsunami wave propagation analysis considering the open boundary condition. Computers and Mathematics with Applications, 16(1–2):139–152, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900284.

Kagiwada:1984:NEU

[KKRS84] H. Kagiwada, R. Kalaba, N. Rasakhoo, and K. Spingarn. Numerical experiments using Sukhanov’s initial-value method for nonlinear two-point boundary value problems. Computers and Mathematics with Applications, 10(4– 5):327–330, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900609.

Kagiwada:1986:ASN

[KKS86] H. H. Kagiwada, R. Kalaba, and K. Spingarn. Automatic solution of nonlinear integral equations. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1169–1184, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186902403.

Kannan:1983:PSN

[KL83] R. Kannan and V. Lakshmikantham. Periodic solutions of nonlinear hyperbolic problems. Computers and Mathematics with Applications, 9(3): 479–486, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900615.

Kothari:1984:CMW

[KL84] S. Kothari and S. Lakshmivarahan. On the concealability of messages by the Williams publickey encryption scheme. Computers and Mathematics with Applications, 10(1):15–24, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122184900828.

Kulenovic:1989:GAP

[KLS89] M. R. S. Kulenović, G. Ladas, and Y. G. Sficas. Global attractivity in population dynamics. Computers and Mathematics with Applications, 18(10– 11):925–928, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900102.

Kaufman:1980:CRD

[KLT80] E. H. Kaufman, Jr., D. J. Leeming, and G. D. Taylor. A combined Remes-differential correction algorithm for rational approximation: experimental results. Computers and Mathematics with Applications, 6(2): 155–160, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900243.

Knoke:1986:REC

[Kno86] James D. Knoke. The robust estimation of classification error rates. Computers and Mathematics with Applications, 12 (Part A) (2):253–260, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900787.

Kim:1988:NNA

[KO88] S. J. Kim and J. T. Oden. A note on the numerical analysis of material damage based on the theory of materials of type N. Computers and Mathematics with Applications, 15(3):169–174, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812218890168X.

Kobayashi:1989:ADI

[Kob89a] M. Kobayashi. An algorithm for discontinuous inverse Sturm– Liouville problems with symmetric potentials. Computers and Mathematics with Applications, 18(4):349–356, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902198.

Kobayashi:1989:EDS

[Kob89b] M. Kobayashi. Eigenvalues of discontinuous Sturm–Liouville problems with symmetric potentials. Computers and Mathematics with Applications, 18 (4):357–364, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902204.

Kohn:1989:CMN

[Koh89] M. C. Kohn. Computer modeling at the national biomedical simulation resource. Computers and Mathematics with Applications, 18(10–11):919–924, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122189900096.

Kopsky:1988:TSS

[Kop88a] V. Kopský. Towards a system in space group representations. Computers and Mathematics with Applications, 16(5– 8):493–505, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902404.

Koptsik:1988:GSC

[Kop88b] V. A. Koptsik. Generalized symmetry in crystal physics. Computers and Mathematics with Applications, 16(5–8):407–424, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902313.

Kostova:1988:NSH

[Kos88] T. V. Kostova. Numerical solutions of a hyperbolic differentialintegral equation. Computers and Mathematics with Applications, 15(6–8):427–436, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902702.

Kouyoumdjian:1989:STA

[Kou89] E. R. Kouyoumdjian. A set theoretic approach to the symmetry analysis of hexcadecahedrane. Computers and Mathematics with Applications, 17(1– 3):375–395, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901697.

See errata [Kou90].

Kouyoumdjian:1990:EST

[Kou90] E. R. Kouyoumdjian. Errata to “A set theoretic approach to the symmetry analysis of hexadecahedrane”: Computers math. applic. 17 (1–3), 375– 395 (1989). Computers and Mathematics with Applications, 19(2):97, ???? 1990. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812219090013A.

See [Kou89].

Kovenya:1980:IDS

[Kov80] V. M. Kovenya. The implicit difference scheme for numerical solving the equations of gas dynamics. Computers and Mathematics with Applications, 6(1):51–59, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900590.

Krajewski:1982:SPU

[Kra82] Witold F. Krajewski. The sequential problem under uncertainty: the development of water systems. Computers and Mathematics with Applications, 8(4):313–318, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218290013X.

Kramer:1987:DME

[Kra87] I. Kramer. A dynamical model estimating the upper and lower bounds of the infectivity of the HTLV–III/LAV virus for the male bisexual and homosexual population in the absence of preventative measures. Computers and Mathematics with Applications, 14(9–12):817–828, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902306.

Kriegsmann:1985:MDN

[Kri85] Gregory A. Kriegsmann. A multiscale derivation of a new parabolic equation which includes density variations. Computers and Mathematics with Applications, 11(7–8):817–821, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0898122185901762.

Kroto:1989:CBO

[Kro89] H. W. Kroto. C_{60}^{B} buckminsterfullerene, other fullerenes and the icospiral shell. Computers and Mathematics with Applications, 17(1–3):417–423, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901715.

Krzanowski:1986:MDA

[Krz86]  W. J. Krzanowski.  Multi-

ple discriminant analysis in the presence of mixed continuous and categorical data. Computers and Mathematics with Applications, 12 (Part A)(2): 179–185, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900714.

Kreith:1985:KCC

[KS85] K. Kreith and C. A. Swanson. Kiguradze classes for characteristics initial value problems. Computers and Mathematics with Applications, 11(1–3): 239–247, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218590149X.

Knop:1984:CGA

[KSTK84] J. V. Knop, K. Szymanski, N. Trinajstić, and P. Krivka. Computer generation of all 1factors for a class of graphs with all vertices of degree two or three. Computers and Mathematics with Applications, 10(4– 5):369–382, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900658.

Kunisch:1985:NAD

[KSW85] K. Kunisch, W. Schappacher, and G. F. Webb. Nonlinear age-dependent population dynamics with random diffusion. Computers and Mathematics with Applications, 11(1–3): 155–173, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901440.

Kalaba:1986:ADF

[KT86] R. Kalaba and L. Tesfatsion. Automatic differentiation of functions of derivatives. Computers and Mathematics with Applications, 12 (Part A)(11):1091– 1103, November 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900143.

Kalaba:1989:TVL

[KT89] R. Kalaba and L. Tesfatsion. Time-varying linear regression via flexible least squares. Computers and Mathematics with Applications, 17(8–9):1215– 1245, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122189900916.

Kuo:1985:NIE

[KTPB85] J. T. Kuo, Y. C. Teng, P. Pecholcs, and J. Blair. A note on the influence of the elasticity on wave propagation in an acoustic/elastic coupled medium. Computers and Mathematics with Applications, 11(7–8):887– 896, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218590183X.

Kalaba:1983:DFA

[KTW83] Robert Kalaba, Asher Tishler, and Jia Song Wang. Decell’s finite algorithm for generalized inverses and tests to determine the rank of a matrix. Computers and Mathematics with Applications, 9(5):679–686, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901256.

Kumar:1989:RMF

[KTW89] S. K. Kumar, W. I. Thacker, and L. T. Watson. Rotating magnetohydrodynamic flow about a stationary disk in a circular magnetic field. Computers and Mathematics with Applications, 17 (11):1435–1439, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122189900758.

Kupershmidt:1986:DIF

[Kup86] Boris A. Kupershmidt. Discretization of ideal fluid dynamics in the Eulerian representation. Computers and Mathematics with Applications, 12(4–5): 389–412, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901719.

Kupershmidt:1987:CIF

[Kup87] B. A. Kupershmidt. Cylindrically invariant flows in relativistic fluid dynamics. Computers and Mathematics with Applications, 13(9–11):881–887, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901702.

Kiefer:1981:CTM

[KW81] Jamer E. Kiefer and George H.

Weiss. A comparison of two methods for accelerating the convergence of Fourier series. Computers and Mathematics with Applications, 7(6):527–535, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122181900365.

Kuhn:1989:CES

[KW89] H. Kuhn and H.-W. Wohltmann. Controllability of economic systems under alternative expectations hypotheses — the discrete case. Computers and Mathematics with Applications, 18(6– 7):617–628, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901120.

Lambert:1986:SSS

[Lam86] J. D. Lambert. A stable sequence of steplengths for Euler’s rule applied to stiff systems of differential equations. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1141–1151, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902385.

Lambert:1988:FPS

[Lam88] J. L. Lambert. Finding a partial solution to a linear system of equations in positive integers. Computers and Mathematics with Applications, 15 (3):209–212, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901721.

Lantos:1989:GRV

[Lan89] F. Lantos. From geometrical rigor to visual experience. Computers and Mathematics with Applications, 17(4–6):931–953, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189902708.

Lau:1980:NBR

[Lau80] William W. Lau. On nodal blocking of a retransmitted packet in a computer communication network. Computers and Mathematics with Applications, 6(3): 355–371, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900437.

Layton:1983:SUN

[Lay83]  William J. Layton.  Stable

and unstable numerical boundary conditions for Galerkin approximations to hyperbolic systems. Computers and Mathematics with Applications, 9(3): 559–566, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218390069X.

Layton:1985:BLT

[Lay85] William Layton. On the behavior over long time intervals of finite difference and finite element approximations to first order hyperbolic systems. Computers and Mathematics with Applications, 11(1–3):93–112, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122185901403.

Layton:1986:HAF

[Lay86] William Layton. High-accuracy finite-element methods for positive symmetric systems. Computers and Mathematics with Applications, 12(4–5):565–579, April/ May 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://

www.sciencedirect.com/science/
article/pii/0898122186901823.

Lacombe:1986:IFG

[LB86] Claude Lacombe and Claude Bédard. Interpolation function of a general triangular mid-edge finite element. Computers and Mathematics with Applications, 12(3):363–373, March 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901951.

Lardner:1982:FDM

[LBC82] R. W. Lardner, M. S. Belen, and H. M. Cekirge. Finite difference model for tidal flows in the Arabian Gulf. Computers and Mathematics with Applications, 8(6):425–444, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900189.

Lardner:1986:NST

[LCG86] R. W. Lardner, H. M. Cekirge, and N. Gunay. Numerical solution of the two-dimensional tidal equations using the method of characteristics. Computers and Mathematics with Applications, 12 (Part A)(10):1065– 1080, October 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690012X.

Lanteigne:1985:MPG

[LdT85] Jacques Lanteigne and Claude de Tourreil. The mechanical performance of GRP used in electrical suspension insulators. Computers and Mathematics with Applications, 11(10): 1007–1021, October 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900215.

Lipitakis:1981:NIM

[LE81] Elias A. Lipitakis and David J.

Evans. Normalized implicit methods for the solution of nonlinear elliptic boundary value problems. Computers and Mathematics with Applications, 7(6): 553–562, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900390.

Lengyel:1989:CKT

[Len89] S. Lengyel. Chemical kinetics and thermodynamics: a history of their relationship. Computers and Mathematics with Applications, 17(1–3):443–455, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189901739.

Letessier:1987:NRB

[Let87] J. Letessier. The numerical resolution of birth and death Kolmogorov equations. Computers and Mathematics with Applications, 13(7):595–600, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901209.

Levit:1982:BRP

[Lev82] Creon Levit. Book review: Programmable calculators, how to use them: By Roger J. Sippl and Charles J. Sippl. Matrix Publishers, Champaign Illinois 1978. approx 148 figures, 526 pp., ISBN 0-916460-08-8, hard cover. Price $15.95. Computers and Mathematics with Applications, 8(2):165, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122182900566.

Levine:1983:MAD

[Lev83] Daniel S. Levine. Models of agedependent predation and cannibalism via the McKendrick equation. Computers and Mathematics with Applications, 9(3): 403–414, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218390055X.

Li:1988:PEC

[Li88] Jia Li. Persistence and extinction in continuous age-structured population models. Computers and Mathematics with Applications, 15(6–8):511–523, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902775.

Li:1989:TEP

[Li89] S. Li. A tutorial on the extended perturbation analysis. Computers and Mathematics with Applications, 17(11):1467–1473, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189900795.

Liebermann:1986:GF

[Lie86] Raanan Liebermann. The gapon formalism. Computers and Mathematics with Applications, 12(3– 4):877–881, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904311.

Lindquist:1986:CST

[Lin86] W. B. Lindquist. Construction of solutions for twodimensional Riemann problems. Computers and Mathematics with Applications, 12(4–5):615– 630, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901859.

Lin:1989:PAG

[Lin89] C.-J. Lin. A parallel algorithm for generating combinations. Computers and Mathematics with Applications, 17(12): 1523–1533, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900527.

Lipitakis:1984:SEB

[Lip84] Elias A. Lipitakis. Solving elliptic boundary-value problems on parallel processors by approximate inverse matrix semi-direct methods based on the multiple explicit Jacobi iteration. Computers and Mathematics with Applications, 10(2):171–184, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122184900476.

Lirov:1989:ECD

[Lir89] Y. Lirov. Electronic circuit diagnostic expert systems — a survey. Computers and Mathematics with Applications, 18 (4):381–398, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990223X.

Lee:1984:IIG

[LJP84] Ding Lee, Kenneth R. Jackson, and Stanley Preiser. An improved implementation of generalized Adams methods for underwater wave propagation problems. Computers and Mathematics with Applications, 10 (3):257–263, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900531.

Lockman:1983:CIN

[LK83] Abe Lockman and David Klappholz. The control of inferencing in natural language understanding. Computers and Mathematics with Applications, 9(1):59–70, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218390007X.

Lee:1985:NTM

[LK85] Ding Lee and Robert M. Kennedy. A numerical treatment of a mixed type dynamic motion equation arising from a towed acoustic antenna in the ocean. Computers and Mathematics with Applications, 11(7– 8):807–816, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901750.

Lee:1988:CFG

[LL88a] C. S. Lee and G. Leitmann. Continuous feedback guaranteeing uniform ultimate boundedness for uncertain linear delay systems: an application to river pollution control. Computers and Mathematics with Applications, 16(10–11):929–938, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902039.

Lee:1988:CFN

[LL88b] E. S. Lee and R.-J. Li. Comparison of fuzzy numbers based on the probability measure of fuzzy events. Computers and Mathematics with Applications, 15(10):887–896, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901241.

Lee:1988:MFE

[LL88c] R. L. Lee and J. M. Leone, Jr.

A modified finite element model for mesoscale flows over complex terrain. Computers and Mathematics with Applications, 16(1–2):41–56, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900235.

Li:1989:AFQ

[LL89] R.-J. Li and E. S. Lee. Analysis of fuzzy queues. Computers and Mathematics with Applications, 17(7):1143–1147, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900448.

Lin:1984:DSD

[LLM84] Ming-Liang Lin, Ernst Leiss, and Bayliss McInnis. Division and sign detection algorithms for residue number systems. Computers and Mathematics with Applications, 10(4–5):331–342, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122184900610.

Lee:1980:TDM

[LM80] H. N. Lee and R. E. Meyers.

On time dependent multi-grid numerical technique. Computers and Mathematics with Applications, 6(1):61–65, ???? 1980.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900607.

Leung:1986:CPF

[LM86a] A. W. Leung and D. A. Murio. L_{2} convergence for positive finite-difference solutions of the diffusive logistic equation in twodimensional bounded domains. Computers and Mathematics with Applications, 12 (Part A) (9):991–1005, September 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900246.

Lim:1986:DCC

[LM86b] J.-T. Lim and S. M. Meerkov. Distributed communications in collision channels with errors. Computers and Mathematics with Applications, 12 (Part A) (6):791–802, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900647.

Lee:1987:P

[LM87] D. Lee and S. T. McDaniel. Preface. Computers and Mathematics with Applications, 14(5):ix, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902069.

Locher:1989:PVS

[LN89] P. Locher and C. Nodine. The perceptual value of symmetry. Computers and Mathematics with Applications, 17(4– 6):475–484, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902460.

Lusk:1982:ERB

[LO82] E. L. Lusk and Ross A.

Overbeek. Experiments with resolution-based theorem-proving algorithms. Computers and Mathematics with Applications, 8(2):141–152, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900530.

Liu:1986:NSE

[LO86] K. M. Liu and E. L. Ortiz. Numerical solution of eigenvalue problems for partial differential equations with the Taulines method. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1153–1168, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902397.

Lewin:1989:IRS

[LO89] J. Lewin and G. J. Olsder. The isotropic rocket — a surveillance evasion game. Computers and Mathematics with Applications, 18(1–3):15–34, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901211.

Loeb:1986:SM

[Loe86a]  A. L. Loeb.  Symmetry

and modularity. Computers and Mathematics with Applications, 12 (Part B)(1–2):63– 75, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901392.

Loeb:1986:SCC

[Loe86b] Arthur L. Loeb. Symmetry in court and country dance. Computers and Mathematics with Applications, 12(3–4):629– 639, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904153.

Loeb:1989:MPD

[Loe89] A. L. Loeb. The magic of the pentangle: Dynamic symmetry from Merlin to Penrose. Computers and Mathematics with Applications, 17(1–3):33–48, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901466.

Los:1989:ILS

[Los89a] C. A. Los. Identification of a linear system from inexact data: a three-variable example. Computers and Mathematics with Applications, 17(8–9):1285–1304, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189900965.

See commentary [Jay92, Zel92] and reply [Los92].

Los:1989:PLS

[Los89b] C. A. Los. The prejudices of least squares, principal components and common factors schemes. Computers and Mathematics with Applications, 17(8– 9):1269–1283, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900953.

See commentary [Jay92, Zel92] and reply [Los92].

Los:1992:RJZ

[Los92] Cornelis A. Los. Reply to E.

T. Jaynes’ and A. Zellner’s comments on my two articles. Computers and Mathematics with Applications, 24(8–9):277–288, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812219290206W.

Lotti:1987:PPB

[Lot87] G. Lotti. A parallel perturbed biharmonic solver. Computers and Mathematics with Applications, 13(8):681–686, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900411.

Loukakis:1983:NBA

[Lou83] E. Loukakis. A new backtracking algorithm for generating the family of maximal independent sets of a graph. Computers and Mathematics with Applications, 9(4):583–589, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183901153.

Louchard:1984:BEA

[Lou84] G. Louchard. The Brownian excursion area: a numerical analysis. Computers and Mathematics with Applications, 10(6): 413–417, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122184900713.

Love:1982:BRP

[Lov82] John Thomas Love. Book review: A programming logic: by Robert L. Constable and Michael J. O’Donnell. Winthrop Publishers, Cambridge, Mass. 1978. 398 pp. $15.96. Computers and Mathematics with Applications, 8(2):165–166, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900578.

Lowrey:1986:SVS

[Low86] Barbara E. Lowrey. A space view of a symmetric object. Computers and Mathematics with Applications, 12 (Part B)(1–2):477– 485, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901665.

Lowrey:1989:ME

[Low89] A. H. Lowrey. Mind’s eye. Computers and Mathematics with Applications, 17(4–6):485–503, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189902472.

Lee:1981:GAM

[LP81] Ding Lee and Stanley Preiser.

Generalized Adams methods for solving underwater wave propagation problems. Computers and Mathematics with Applications, 7(2):195–202, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812218190119X.

Lakshmikantham:1985:PSH

[LP85] V. Lakshmikantham and S. G. Pandit. Periodic solutions of hyperbolic partial differential equations. Computers and Mathematics with Applications, 11(1–3): 249–259, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901506.

Liebman:1989:SSD

[LRBK89] J. F. Liebman, J. M. Richter, M. J. Bienlein, and S. S. Kulkarni. Square spirals, dimensionality and biopolymers. Computers and Mathematics with Applications, 17(4–6):595–611, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189902514.

Lirov:1989:ALT

[LRG89] Y. Lirov, E. Y. Rodin, and B. K. Ghosh. Automated learning by tactical decision systems in air combat. Computers and Mathematics with Applications, 18(1– 3):151–160, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901326.

Lai:1987:PND

[LT87] H.-C. Lai and K. Tanaka. An n-person noncooperative discounted vector valued dynamic game with a stopped set. Computers and Mathematics with Applications, 13(1–3):227–237, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901064.

Luke:1983:SDE

[LTW83] Yudell L. Luke, Bing Y. Ting, and Jet Wimp. Solution of the difference equation for Gauss’ hypergeometric function by use of the tau method. Computers and Mathematics with Applications, 9(5):659–668, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901232.

Luke:1981:SSO

[Luk81] Yudell L. Luke. Solution of second order difference equations via the τ-method. Computers and Mathematics with Applications, 7(5):439–445, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901310.

Lumb:1987:LDR

[Lum87] J. R. Lumb. Lymphocyte differentiation, repertoire development and migration: the need for mathematical models. Computers and Mathematics with Applications, 14(9–12):657–697, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902215.

Lirov:1986:SAS

[LV86] Yuval Lirov and Michael W. Vannier. System analysis of stereotactic neurosurgery based on CT scans. Computers and Mathematics with Applications, 12 (Part A)(7):839–858, July 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900301.

Lemczyk:1988:EEI

[LY88] T. F. Lemczyk and M. M. Yovanovich. Efficient evaluation of incomplete elliptic integrals and functions. Computers and Mathematics with Applications, 16(9):747–757, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900107.

MacCamy:1983:SPM

[Mac83] R. C. MacCamy. Simple population models with diffusion. Computers and Mathematics with Applications, 9(3):341–344, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122183900214.

MacGillavry:1986:SMC

[Mac86a] Caroline H. MacGillavry. The symmetry of M. C. Escher’s “impossible” images. Computers and Mathematics with Applications, 12 (Part B)(1–2):123– 138, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690146X.

Mackay:1986:WS

[Mac86b] Alan L. Mackay. But what is symmetry? Computers and Mathematics with Applications, 12 (Part B)(1–2):19– 20, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901367.

Mackay:1986:GC

[Mac86c] Alan L. Mackay. Generalised crystallography. Computers and Mathematics with Applications, 12 (Part B)(1–2):21– 37, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901379.

Makovicky:1986:SAC

[Mak86a] Emil Makovicky. Symmetrology of art: coloured and generalized symmetries. Computers and Mathematics with Applications, 12(3–4):949–980, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186904359.

Maksic:1986:SHB

[Mak86b] Zvonimir B. Maksić. Symmetry, hybridization and bonding in molecules. Computers and Mathematics with Applications, 12(3– 4):697–723, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904190.

Makovicky:1989:OBT

[Mak89] E. Makovicky. Ornamental brickwork: Theoretical and applied symmetrology and classification of patterns. Computers and Mathematics with Applications, 17(4–6):955–999, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990271X.

Malikov:1989:STS

[Mal89] A. V. Malikov. Structural theory of space-time and intrapoint symmetry. Computers and Mathematics with Applications, 17(1– 3):279–299, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122189901636.

Mamedov:1986:CP

[Mam86] Kh. S. Mamedov. Crystallographic patterns. Computers and Mathematics with Applications, 12(3–4):511–529, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186904086.

Mansfield:1981:MFE

[Man81] Lois Mansfield. On mixed finite element methods for elliptic equations. Computers and Mathematics with Applications, 7(1):59–66, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900079.

Mannheim:1986:SSB

[Man86] Philip D. Mannheim. Symmetry and spontaneously broken symmetry in the physics of elementary particles. Computers and Mathematics with Applications, 12 (Part B)(1–2):169– 183, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901495.

Mansfield:1987:FEA

[Man87] L. Mansfield. Finite element approximation of stationary interior viscous flow problems. Computers and Mathematics with Applications, 14(4):249–254, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901325.

Mao:1986:OOS

[Mao86] Jianqin Mao. Optimal orthonormalization of the strapdown matrix by using singular value decomposition. Computers and Mathematics with Applications, 12 (Part 1)(3):353–362, March 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218690194X.

Marcati:1983:SCM

[Mar83] Pierangelo Marcati. Some considerations on the mathematical approach to nonlinear age dependent population dynamics. Computers and Mathematics with Applications, 9(3):361–370, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218390024X.

Marcati:1985:NHS

[Mar85a] P. Marcati. On a nonconservative huyperbolic system describing the nonlinear age dependent populations growth. Computers and Mathematics with Applications, 11(1–3):207–222, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901476.

Martinez:1985:PMS

[Mar85b] JoséMario Martínez. The projection method for solving nonlinear systems of equations under the “most violated constraint” control. Computers and Mathematics with Applications, 11(10): 987–993, October 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900197.

Marcati:1986:AST

[Mar86a] Pierangelo Marcati. Abstract stability theory and applications to hyperbolic equations with time dependent dissipative force fields. Computers and Mathematics with Applications, 12(4–5): 541–550, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901793.

Marchi:1986:CE

[Mar86b] E. Marchi. Cooperative equilibrium. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1185–1186, September/ December 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0898122186902415.

Marsaglia:1986:IFC

[Mar86c] G. Marsaglia. The incomplete Γ function as a continuous Poisson distribution. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1187–1190, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186902427.

Martinez:1988:ASO

[Mar88] J. M. Martínez. An accelerated successive orthogonal projections method for solving large-scale linear feasibility problems. Computers and Mathematics with Applications, 15(5):367–373, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188901587.

Marcus:1989:SSC

[Mar89] S. Marcus. Symmetry in the simplest case: the real line. Computers and Mathematics with Applications, 17(1–3):103–115, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812218990151X.

McKee:1980:RFN

[MB80] S. McKee and H. Brunner. The repetition factor and numerical stability of Volterra integral equations. Computers and Mathematics with Applications, 6(3):339–347, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900413.

Mukhopadhyay:1989:EDI

[MB89] B. Mukhopadhyay and R. K.

Bera. Effect of distributed instantaneous and continuous heat sources in an infinite conducting magneto-thermo-viscoelastic solid with thermal relaxation. Computers and Mathematics with Applications, 18(8):723– 728, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902289.

Moschopoulos:1984:DFL

[MC84] P. G. Moschopoulos and W. B.

Canada. The distribution function of a linear combination of chi-squares. Computers and Mathematics with Applications, 10(4–5):383–386, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218490066X.

McCalla:1985:TID

[MC85] Gordon McCalla and Nick Cercone. Techniques and issues in the design of applied artificial intelligence systems. Computers and Mathematics with Applications, 11(5):421–430, May 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900483.

McAllister:1988:FIG

[McA88] M. L. N. McAllister. Fuzzy intersection graphs. Computers and Mathematics with Applications, 15(10):871–886, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218890123X.

McCalla:1983:AOK

[McC83] Gordon I. McCalla. An approach to the organization of knowledge and its use in natural language recall tasks. Computers and Mathematics with Applications, 9(1):201–214, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900159.

McCoy:1985:HBV

[McC85] Peter A. McCoy. Hyperbolic boundary value problems arising from symmetric Poisson processes. Computers and Mathematics with Applications, 11 (1–3):307–313, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122185901555.

McCoy:1986:SHP

[McC86] Peter A. McCoy. Singularities of hyperbolic PDEs in two complex variables. Computers and Mathematics with Applications, 12 (4–5):551–556, April/May 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690180X.

McDonald:1983:DDC

[McD83] David D. McDonald. Description directed control: Its implications for natural language generation. Computers and Mathematics with Applications, 9(1): 111–129, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900111.

McDaniel:1985:AEM

[McD85] Suzanne T. McDaniel. Applications of energy methods to finite-difference solutions of the parabolic wave equation. Computers and Mathematics with Applications, 11(7–8):823– 829, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901774.

McLachlan:1986:APA

[McL86] G. J. McLachlan. Assessing the performance of an allocation rule. Computers and Mathematics with Applications, 12 (Part A) (2):261–272, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900799.

McWhinnie:1989:IIJ

[McW89] H. J. McWhinnie. Influences of the ideas of jay hambidge on art and design. Computers and Mathematics with Applications, 17(4–6):1001–1008, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902721.

Martinez:1986:GTR

[MD86] Y. Martinez and A. Pham Ngoc Dinh. A generalization of Tikhonov’s regularization of zero and first order. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1203–1208, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186902440.

Minambres:1986:ANM

[MdlS86] J. J. Miñambres and M. de la Sen. Application of numerical methods to the acceleration of the convergence of the adaptive control algorithms: the one-dimensional case. Computers and Mathematics with Applications, 12 (Part A)(10): 1049–1056, October 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900106.

Mangalgiri:1985:FCP

[MDR85] P. D. Mangalgiri, B. Dattaguru, and A. K. Rao. Fasteners in composite plates: effects of interfacial friction. Computers and Mathematics with Applications, 11(10):1057–1068, October 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900240.

Meade:1986:BRR

[Mea86a] Kevin P. Meade. Book review:

Review of PDE software: Modules, interfaces and systems: Elsevier Science Publishing Company, Inc., New York, 1984, B. Engquist and T. Smedsaas, Eds. Computers and Mathematics with Applications, 12(4– 5):632, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901884.

Meade:1986:RNP

[Mea86b] Kevin P. Meade. Review of Nonlinear Partial Differential Equations in Engineering and Applied Science, Lecture Notes in Pure and Applied Mathematics: Vol. 54, Marcel Dekker, Inc., New York and Basel, 1984, R. L. Sternberg, A. J. Kalinowski and J. S. Papadakis, Eds. Computers and Mathematics with Applications, 12(4–5): 631–632, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901872.

Medhin:1988:MMMa

[Med88a] N. G. Medhin. A mixed min-max control problem governed by integral equations. Computers and Mathematics with Applications, 16(10–11):867–875, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901964.

Medhin:1988:MMMb

[Med88b] N. G. Medhin. A mixed minmax control problem with state restrictions I. Computers and Mathematics with Applications, 16(10–11):877–886, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901976.

Meinardus:1986:ATA

[Mei86a] G. Meinardus. An approximation theoretic alternative to asymptotic expansions for special functions. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1191–1201, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902439.

Meinhardt:1986:FSA

[Mei86b] Hans Meinhardt. Formation of symmetric and asymmetric structures during development of higher organisms. Computers and Mathematics with Applications, 12 (Part B)(1–2):419– 433, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690163X.

Menzala:1986:SPW

[Men86] G. Perla Menzala. Scattering properties of wave equations with time-dependent potentials. Computers and Mathematics with Applications, 12(4–5): 457–475, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901744.

Merz:1987:SGL

[Mer87] A. W. Merz. Stochastic guidance laws in satellite pursuit– evasion. Computers and Mathematics with Applications, 13(1– 3):151–156, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901003.

Metropolis:1980:SIN

[Met80] N. Metropolis. Summation of imprecise numbers. Computers and Mathematics with Applications, 6(3):297–299, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900371.

Murio:1989:NIF

[MH89] D. A. Murio and D. Hinestroza. Numerical identification of forcing terms by discrete mollification. Computers and Mathematics with Applications, 17(11): 1441–1447, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218990076X.

Mouradd:1987:ESF

[MHM87] S. Mouradd, J. L. A. Hughes, and E. J. McCluskey. Effectiveness of single fault tests to detect multiple faults in parity trees. Computers and Mathematics with Applications, 13(5– 6):455–459, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122187900757.

Milder:1982:G

[Mil82] Benjamin Milder. The golem.

Computers and Mathematics with Applications, 8(6):467, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122182900207.

Miller:1986:OUD

[Mil86]  J. J. H. Miller.  Optimal

uniform difference schemes for linear initial-value problems. Computers and Mathematics with Applications, 12 (Part 2) (5–6):1209–1215, September/ December 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0898122186902452.

Mittnik:1989:MTS

[Mit89a] S. Mittnik. Multivariate time series analysis with state space models. Computers and Mathematics with Applications, 17(8– 9):1189–1201, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900898.

Mittnik:1989:P

[Mit89b] S. Mittnik. Preface. Computers and Mathematics with Applications, 18(6–7):vii, ???? 1989.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901016.

McDonald:1985:TDS

[MK85] B. Edward McDonald and W. A. Kuperman. Timedomain solution of the parabolic equation including nonlinearity. Computers and Mathematics with Applications, 11(7–8): 843–851, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901798.

MacKay:1986:TGI

[MK86] A. L. MacKay and J. Klinowski. Towards a grammar of inorganic structure. Computers and Mathematics with Applications, 12(3– 4):803–824, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904268.

Meunier:1983:FSC

[ML83] J. G. Meunier and F. Lepage. Formal semantic and computer text processing, 1982. Computers and Mathematics with Applications, 9(1):83–95, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900093.

McCahon:1988:PNA

[ML88] C. S. McCahon and E. S. Lee.

Project network analysis with fuzzy activity times. Computers and Mathematics with Applications, 15(10):829–838, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901204.

Mills:1987:CSM

[MLDR87] P. L. Mills, S. Lai, M. P.

Duduković, and P. A. Ramachandran. Comparison of solution methods for a mathematical model of dilute solute mass transfer with chemical reaction on a nonuniform surface. Computers and Mathematics with Applications, 14(2):139–151, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901441.

Martin:1984:RMC

[MM84] Robert J. Martin and Gary L.

Mullen. Reducing multinomial coefficients modulo a prime power. Computers and Mathematics with Applications, 10 (1):37–41, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900841.

Molnar:1986:SMB

[MM86a] V. Molnar and F. Molnar. Symmetry-making and -breaking in visual art. Computers and Mathematics with Applications, 12 (Part B)(1–2):291– 301, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901550.

Murphy:1986:PKD

[MM86b] B. J. Murphy and M. A. Moran. Parametric and kernel density methods in discriminant analysis: Another comparison. Computers and Mathematics with Applications, 12 (Part A)(2): 197–207, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900738.

Mitas:1988:GVA

[MM88] L. Mitás and H. Mitásová. General variational approach to the interpolation problem. Computers and Mathematics with Applications, 16(12):983–992, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188902556.

Mitra:1983:EPA

[MN83] A. K. Mitra and Paul Nelson. An eigenlength problem arising from a model of patch formation. Computers and Mathematics with Applications, 9(4): 591–593, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901165.

Manohar:1988:AIF

[MN88] R. Manohar and T. Y. Ngai.

Application of iteration functions to elliptic partial differential equations. Computers and Mathematics with Applications, 16(4):279–285, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901447.

Moller:1986:SIU

[Møl86] Aage R. Møller. Systems identification using pseudorandom noise applied to a sensorineural system. Computers and Mathematics with Applications, 12 (Part A)(6):803–814, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900659.

Molnar:1988:MPC

[Mol88] E. Molnár. Minimal presentation of crystallographic groups by fundamental polyhedra. Computers and Mathematics with Applications, 16(5–8):507–520, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188902416.

Morawetz:1981:FHD

[Mor81a] Cathleen Synge Morawetz. A formulation for higher dimensional inverse problems for the wave equation. Computers and Mathematics with Applications, 7(4):319–331, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900614.

Mori:1981:AND

[Mór81b] Tamás F. Móri. On the asymptotic network delay in a model of packet switching. Computers and Mathematics with Applications, 7(2):167–172, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181901164.

Morhac:1986:PDU

[Mor86] Miroslav Morhác. Precise deconvolution using the Fermat number transform. Computers and Mathematics with Applications, 12(3):319–329, March 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901914.

Morhac:1989:DEF

[Mor89] M. Morhác. k-dimensional errorfree deconvolution using the fermat number transform. Computers and Mathematics with Applications, 18(12):1023–1032, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189900308.

Modi:1986:MJS

[MP86] J. J. Modi and J. D. Pryce. Mobile Jacobi schemes for parallel computation. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1217–1224, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902464.

Manthey:1988:UHS

[MP88] Michael J. Manthey and Keith Phillips. A unique hierarchy on the successive condensations of a digraph. Computers and Mathematics with Applications, 15(4):319–323, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902180.

Miloh:1989:SCA

[MP89] T. Miloh and M. Pachter. Ship collision-avoidance and pursuit– evasion differential games with speed-loss in a turn. Computers and Mathematics with Applications, 18(1–3):77–100, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901260.

Marchesin:1986:RPG

[MPL86] D. Marchesin and P. J. Paes-Leme. A Riemann problem in gas dynamics with bifurcation. Computers and Mathematics with Applications, 12(4–5): 433–455, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901732.

Moritz:1987:PEM

[MPW87] K. Moritz, R. Polis, and K. H. Well. pursuit–evasion in medium-range air-combat scenarios. Computers and Mathematics with Applications, 13(1– 3):167–180, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901027.

Mills:1988:MMC

[MR88a] P. L. Mills and P. A. Ramachandran. Mathematical modelling of chemical engineering systems by finite element analysis using PDE/PROTRAN. Computers and Mathematics with Applications, 15(9):769–794, ???? 1988.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901290.

Misra:1988:FFV

[MR88b] J. C. Misra and S. Roy. Free and forced vibrations of the cranial vault: Study of a mathematical model. Computers and Mathematics with Applications, 16(3):247–266, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901848.

Molnarka:1988:ISS

[MR88c] G. Molnárka and B. Ráczkevi.

Implicit single step methods by spline-like functions for solution of ordinary differential equations. Computers and Mathematics with Applications, 16 (9):701–704, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900053.

Murio:1988:ISI

[MR88d] D. A. Murio and C. C. Roth. An integral solution for the inverse heat conduction problem after the method of Weber. Computers and Mathematics with Applications, 15(1):39–51, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188900703.

Mayo:1989:SPV

[MR89a] C. F. Mayo and C. E. Roberts. Sequential, parallel and vector solution of ordinary differential equations on a hypercube. Computers and Mathematics with Applications, 18(9):797–808, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189901788.

Misra:1989:SNE

[MR89b] J. C. Misra and S. Roy. Stresses in the neighbourhood of an exterior crack in a bone specimen. Computers and Mathematics with Applications, 17(11): 1493–1502, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900825.

Mokari:1985:CSA

[MRV85] Shahram Aivaz-Zadeh Mokari, Epiphane Ramahefarison, and Georges Verchery. Composite structure analysis with microcomputer using classical and interface finite elements. Computers and Mathematics with Applications, 11(10):1023– 1042, October 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122185900227.

Mahoney:1981:CSO

[MS81] J. F. Mahoney and B. D. Sivazlian. Computation of stockouts in a multi-installation inventory system. Computers and Mathematics with Applications, 7(5):369–386, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218190122X.

MacCamy:1985:NPP

[MS85] R. C. MacCamy and Eduardo Socolovsky. A numerical procedure for the porous media equation. Computers and Mathematics with Applications, 11(1–3): 315–319, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901567.

Medhin:1988:NSS

[MS88a] N. Medhin and M. Sambandham. Numerical solution of a system of random Volterra integral equations-1. successive approximation method. Computers and Mathematics with Applications, 15(2):97–105, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890079X.

Misra:1988:FTB

[MS88b] J. C. Misra and B. K. Sahu. Flow through blood vessels under the action of a periodic acceleration field: a mathematical analysis. Computers and Mathematics with Applications, 16(12):993–1016, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902568.

Misra:1988:MAV

[MS88c] J. C. Misra and S. Samanta. A mathematical analysis of the vibration characteristics of the human tibia. Computers and Mathematics with Applications, 16(12):1017–1026, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218890257X.

Misra:1988:MMB

[MS88d] J. C. Misra and S. Samanta. A mathematical model for the biomechanics of the body of a vertebra. Computers and Mathematics with Applications, 15 (2):85–96, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900788.

Misra:1988:SMA

[MS88e] J. C. Misra and S. I. Singh. Study on the mechanics of aneurysms in the left ventricle of the heart. Computers and Mathematics with Applications, 15 (1):17–27, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900685.

Mugler:1988:FFT

[MS88f] D. H. Mugler and R. A. Scott.

Fast Fourier Transform method for partial differential equations, case study: the 2-D diffusion equation. Computers and Mathematics with Applications, 16 (3):221–228, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901824.

McMath:1980:CVD

[MSH80] John S. McMath, S. E. Sims, and C. R. Hallum. A comparison of various derivatives for matrix valued functions. Computers and Mathematics with Applications, 6(2):161–166, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900255.

McKenna:1985:FMV

[MT85] Michael McKenna and Godfried T. Toussaint. Finding the minimum vertex distance between two disjoint convex polygons in linear time.

Computers and Mathematics with Applications, 11(12):1227– 1242, December 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901099.

McCormick:1982:EFM

[MTP82] S. F. McCormick, G. D. Taylor, and D. V. Pryor. Evaluation of functions on microcomputers: ln(x). Computers and Mathematics with Applications, 8(5):389–392, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122182900323.

Murrell:1982:CMS

[MU82] H. C. Murrell and Abraham Ungar. From Cagniard’s method for solving seismic pulse problems to the method of the differential transform. Computers and Mathematics with Applications, 8(2):103–118, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122182900505.

Mukai:1981:PAS

[Muk81] H. Mukai. Parallel algorithms for solving systems of nonlinear equations. Computers and Mathematics with Applications, 7(3):235–250, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900833.

Murphy:1983:DDC

[Mur83] Lea Murphy. Density dependent cellular growth in an age structured colony. Computers and Mathematics with Applications, 9(3):383–392, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900536.

Murio:1987:AND

[Mur87] D. A. Murio. Automatic numerical differentiation by discrete mollification. Computers and Mathematics with Applications, 13(4):381–386, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218790006X.

Murio:1988:NRI

[Mur88] D. A. Murio. Noise reconstruction for the inverse heat conduction problem. Computers and Mathematics with Applications, 16(12):1027–1033, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902581.

Murio:1989:MMN

[Mur89] D. A. Murio. The mollification method and the numerical solution of the inverse heat conduction problem by finite differences. Computers and Mathematics with Applications, 17(10): 1385–1396, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900229.

Messer:1989:SPS

[MW89] P. W. Messer and M. J. Wenninger. Symmetry and polyhedral stellation — II. Computers and Mathematics with Applications, 17(1–3):195–201, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901594.

Mazzola:1989:SOM

[MWBM89] G. Mazzola, H.-G. Wieser, V. Brunner, and D. Muzzulini. A symmetry-oriented mathematical model of classical counterpoint and related neurophysiological investigations by depth EEG. Computers and Mathematics with Applications, 17(4– 6):539–594, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902502.

Myers:1987:FCM

[Mye87] R. E. Myers. Fundamental concepts of microcomputer graphics. Computers and Mathematics with Applications, 14 (1):33–42, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901799.

Neta:1988:TFA

[ND88] B. Neta and C. L. DeVito.

The transfer function analysis of various schemes for the twodimensional shallow-water equations. Computers and Mathematics with Applications, 16(1– 2):111–137, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900272.

Neta:1980:TI

[Net80] Beny Neta. On three inequalities. Computers and Mathematics with Applications, 6(3): 301–304, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900383.

Neuberger:1980:TII

[Neu80] J. W. Neuberger. A typeindependent iterative method for systems of nonlinear partial differential equations: Application to the problem of transonic flow. Computers and Mathematics with Applications, 6(1):67–78, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900619.

Ng:1981:SMF

[Ng81] Edward W. Ng. Simplification of multiple Fourier series: an example of algorithmic approach. Computers and Mathematics with Applications, 7(5): 355–359, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901206.

Nicolau:1987:BRT

[Nic87] E. Nicolau. Book review: Technical, technological and social implications of bioengineering (implicatii tehnice, tehnologice si sociale ale bioingineriei): By E. Niculescu-Mizil. Editura Stiintifică si enciclopedică, Bucuresti (1982) 240 pp. 13 lei. Computers and Mathematics with Applications, 14(9–12):952–953, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902483.

Neta:1988:I

[NN88] B. Neta and I. M. Navon. Introduction. Computers and Mathematics with Applications, 16 (1–2):vii, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900193.

Neittaanmaki:1981:FEM

[NP81] P. Neittaanmäki and R. Picard.

On the finite element method for time-harmonic acoustic boundary value problems. Computers and Mathematics with Applications, 7(2):127–138, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901115.

Ng:1986:DFP

[NR86] Tie-Hua Ng and Ronald H. Randles. Distribution-free partial discrimination procedures. Computers and Mathematics with Applications, 12 (Part A)(2): 225–234, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900751.

Ng:1988:DPR

[NS88] K. Y. K. Ng and N. G. F. Sancho.

Dynamic programming, reduction of dimensionality and eigenvalue problems. Computers and Mathematics with Applications, 16(10–11):789–799, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901897.

Neta:1985:FPC

[NT85] Beny Neta and Heng-Ming Tai. LU factorization on parallel computers. Computers and Mathematics with Applications, 11 (6):573–579, June 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900392.

Obaid:1988:SFC

[OB88] S. A. Obaid and A. Benharbit. Stream functions of certain rotating cylinders. Computers and Mathematics with Applications, 15(6–8):437–446, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902714.

OConnor:1987:BRIb

[O’C87a] C. O’Connor. Book review: An introduction to multivariate statistical analysis: 2nd edn. by T. W. Anderson. 675 pp. Wiley, New York (1984). Computers and Mathematics with Applications, 14(9–12):951, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218790246X.

OConnor:1987:BRIa

[O’C87b] C. O’Connor. Book review:

Image analysis and mathematical morphology: By J. Serra. academic press, New York (1982). Computers and Mathematics with Applications, 14 (9–12):949, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902434.

Odanaka:1988:IDO

[Oda88] T. Odanaka. Information and decision in optimal inventory processes. Computers and Mathematics with Applications, 16(10– 11):815–822, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901915.

Oden:1980:P

[Ode80] J. T. Oden. Preface. Computers and Mathematics with Applications, 6(1):iii, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900541.

Outlaw:1986:SOI

[ODS86] Curtis Outlaw, Leroy Derr, and Diran Sarafyan. A sixth-order imbedded Runge–Kutta algorithm with continuously variable weights. Computers and Mathematics with Applications, 12 (Part A)(6):815–824, June 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900660.

Oden:1982:IPM

[OK82] J. T. Oden and S. J. Kim. Interior penalty methods for finite element approximations of the Signorini problem in elastostatics. Computers and Mathematics with Applications, 8 (1):35–56, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900384.

Ovchinnikova:1988:SNS

[OK88] E. N. Ovchinnikova and R. N. Kuz’min. Symmetry in the nuclear solid state physics. Computers and Mathematics with Applications, 16(5–8):657–661, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902532.

Omodeo:1982:LCM

[Omo82] Eugenio G. Omodeo. The linked conjunct method for automatic deduction and related search techniques. Computers and Mathematics with Applications, 8(3):185–203, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122182900426.

ONeill:1986:VER

[O’N86] Terence J. O’Neill. The variance of the error rates of classification rules. Computers and Mathematics with Applications, 12 (Part A) (2):273–287, February 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900805.

ONeill:1989:CTO

[O’N89] W. D. O’Neill. The coding theorem and ordinary least-squares models in economics. Computers and Mathematics with Applications, 17(8–9):1259–1264, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218990093X.

Ortiz:1986:BTE

[OP86] E. L. Ortiz and K.-S. Pun.

A bidimensional tau-elements method for the numerical solution of nonlinear partial differential equations with an application to Burgers’ equation. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1225–1240, September/ December 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0898122186902476.

Orlowski:1987:LPS

[OP87] M. Orlowski and M. Pachter. Linear programming in R^{3} and the skeleton and largest incircle of a convex polygon. Computers and Mathematics with Applications, 13(4):401–405, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900083.

Orlowski:1989:ADL

[OP89] M. Orlowski and M. Pachter. An algorithm for the determination of a longest increasing subsequence in a sequence. Computers and Mathematics with Applications, 17(7):1073–1075, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189900369.

Orlowski:1985:DAP

[Orl85] M. Orlowski. On the determination of all pairs of symmetric furthest neighbor vertices of a simple polygon. Computers and Mathematics with Applications, 11(4):347–353, April 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185900598.

Ortiz:1986:PYL

[Ort86] Eduardo L. Ortiz. Professor Yudell L. Luke: From tables to algorithms. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1047–1052, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902294.

Ortiz:1984:NSP

[OS84] E. L. Ortiz and H. Samara. Numerical solution of partial differential equations with variable coefficients with an operational approach to the tau method. Computers and Mathematics with Applications, 10(1):5–13, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900816.

Osborne:1986:SAF

[Osb86] Harold Osborne. Symmetry as an aesthetic factor. Computers and Mathematics with Applications, 12 (Part B)(1–2):77– 82, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901409.

Otter:1989:SSA

[OvD89] P. W. Otter and R. van Dal.

State-space approximation of multi-input multi-output systems with stochastic exogenous inputs. Computers and Mathematics with Applications, 18(6– 7):529–538, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901053.

Ozawa:1981:EMV

[Oza81] K. Ozawa. Estimation for the mean and the variance of the round-off error. Computers and Mathematics with Applications, 7(4):285–291, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900572.

Petrodulas:1985:ACC

[PAC85] T. Petrodulas, R. Aris, and R. W. Carr. Analysis of the counter-current movingbed chromatographic reactor. Computers and Mathematics with Applications, 11(1–3):5– 34, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901361.

Pachter:1987:SMP

[Pac87]  M. Pachter.  Simple-motion

pursuit–evasion in the half plane. Computers and Mathematics with Applications, 13 (1–3):69–82, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122187900940.

Pach:1989:DS

[Pac89]  J. Pach.  Delicate symme-

try. Computers and Mathematics with Applications, 17(1– 3):117–124, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901521.

Pan:1981:BOC

[Pan81a] V. Ya. Pan. The bit-operation complexity of matrix multiplication and of all pair shortest path problem. Computers and Mathematics with Applications, 7(5):431–438, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901309.

Pan:1981:NCM

[Pan81b] V. Ya. Pan. New combinations of methods for the acceleration of matrix multiplications. Computers and Mathematics with Applications, 7(1):73–125, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900092.

Pan:1982:BOC

[Pan82a] V. Pan. The bit-operation complexity of approximate evaluation of matrix and polynomial products using modular arithmetic. Computers and Mathematics with Applications, 8(2): 137–140, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900529.

Pan:1982:FMM

[Pan82b] V. Pan. Fast matrix multiplication without APA-algorithms. Computers and Mathematics with Applications, 8(5):343– 366, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900293.

Pan:1982:TAI

[Pan82c] V. Ya. Pan. Trilinear aggregating with implicit canceling for a new acceleration of matrix multiplication. Computers and Mathematics with Applications, 8(1):23–34, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900372.

Pan:1983:PPM

[Pan83] V. Ya. Pan. The projective power method for the algebraic eigenvalue problem. Computers and Mathematics with Applications, 9(6):735–745, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901384.

Pan:1985:FFM

[Pan85a] V. Pan. Fast finite methods for a system of linear inequalities. Computers and Mathematics with Applications, 11 (4):355–394, April 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900604.

Pan:1985:CPS

[Pan85b] V. Pan. On the complexity of a pivot step of the revised simplex algorithm. Computers and Mathematics with Applications, 11 (11):1127–1140, November 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901907.

Pan:1985:BCM

[Pan85c] V. Y. Pan. The bit complexity of matrix multiplication and of related computations in linear algebra. The segmented λ algorithms. Computers and Mathematics with Applications, 11(9): 919–928, September 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900951.

Pan:1985:ASR

[Pan85d] V. Y. Pan. On application of some recent techniques of the design of algebraic algorithms to the sequential and parallel evaluation of the roots of a polynomial and to some other numerical problems. Computers and Mathematics with Applications, 11(9):911–917, September 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218590094X.

Panossian:1986:RBS

[Pan86] Hagop Panossian. Richard Bellman and stochastic control systems. Computers and Mathematics with Applications, 12 (Part A)(6):825–829, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900672.

Pan:1987:ACC

[Pan87a] V. Pan. Algebraic complexity of computing polynomial zeros. Computers and Mathematics with Applications, 14(4): 285–304, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901374.

Pan:1987:SPC

[Pan87b] V. Pan. Sequential and parallel complexity of approximate evaluation of polynomial zeros. Computers and Mathematics with Applications, 14(8):591–622, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901866.

Pan:1989:FEPa

[Pan89] V. Pan. Fast and efficient parallel evaluation of the zeros of a polynomial having only real zeros. Computers and Mathematics with Applications, 17(11): 1475–1480, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900801.

Partridge:1988:CNQ

[Par88] D. Partridge. Connectionist networks qua graphs. Computers and Mathematics with Applications, 15(4):325–331, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902192.

Pasciak:1982:IAV

[Pas82] Joseph E. Pasciak. An iterative algorithm for the volume integral method for magnetostatics problems. Computers and Mathematics with Applications, 8(4):283–290, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122182900104.

Pauling:1989:ICI

[Pau89] L. Pauling. Interpretation of socalled icosahedral and decagonal quasicrystals of alloys showing apparent icosahedral symmetry elements as twins of an 820-atom cubic crystal. Computers and Mathematics with Applications, 17(1–3):337–339, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901661.

Presler:1981:CFSa

[PB81a] W. H. Presler and R. Broucke. Computerized formal solutions of dynamical systems with two degrees of freedom and an application to the contopoulos potential — I: the exact resonance case. Computers and Mathematics with Applications, 7(6): 451–471, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900304.

Presler:1981:CFSb

[PB81b] W. H. Presler and R. Broucke. Computerized formal solutions of dynamical systems with two degrees of freedom and an application to the Contopoulos potential — II: the near-resonance case. Computers and Mathematics with Applications, 7(6):

473–485, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900316.

Piessens:1982:ABF

[PB82] R. Piessens and Maria Branders.

Approximation for Bessel functions and their application in the computation of Hankel transforms. Computers and Mathematics with Applications, 8(4): 305–311, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900128.

Piessens:1986:CFL

[PB86] R. Piessens and M. Branders. Computation of Fourier and Laplace transforms of singular functions using modified moments. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1241–1248, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902488.

Pearson:1986:SRC

[Pea86] Ralph G. Pearson. Symmetry rules for chemical reactions. Computers and Mathematics with Applications, 12 (Part B)(1–2):229–236, January/April 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186901525.

Periti:1987:STA

[Per87] P. Periti. A system theoretical approach to the optimization of cancer treatment. Computers and Mathematics with Applications, 14(9–12):803–816, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812218790229X.

Petkovic:1987:SII

[Pet87]  M. S. Petković.  Some in-

terval iterations for finding a zero of a polynomial with error bounds. Computers and Mathematics with Applications, 14 (6):479–495, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900605.

Petukhov:1989:NEG

[Pet89] S. V. Petukhov. Non-Euclidean geometries and algorithms of living bodies. Computers and Mathematics with Applications, 17(4– 6):505–534, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902484.

Perez-Gomez:1987:FRM

[PG87] R. Pérez-Gómez. The four regular mosaics missing in the Alhambra. Computers and Mathematics with Applications, 14 (2):133–137, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218790143X.

Pickover:1989:HMT

[Pic89a] C. A. Pickover. Halley maps for a trigonometric and rational function. Computers and Mathematics with Applications, 17(1– 3):125–132, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901533.

Pickover:1989:MBV

[Pic89b] C. A. Pickover. Mathematics and beauty — VIII: Tesselation automata derived from a single defect. Computers and Mathematics with Applications, 17(1– 3):321–336, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990165X.

Piessens:1984:CBF

[Pie84] Robert Piessens. The computation of Bessel functions on a small computer. Computers and Mathematics with Applications, 10(2):161–166, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900452.

Pierce:1985:NRW

[Pie85] Allan D. Pierce. The natural reference wavenumber for parabolic approximations in ocean acoustics. Computers and Mathematics with Applications, 11(7–8): 831–841, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901786.

Pitcher:1984:TSE

[PM84] N. Pitcher and S. McKee. Two sided error bounds for discretisation methods in special q th order ordinary differential equations. Computers and Mathematics with Applications, 10 (3):265–277, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900543.

Pachter:1987:GAC

[PM87] M. Pachter and T. Miloh. The geometric approach to the construction of the barrier surface in differential games. Computers and Mathematics with Applications, 13(1–3):47–67, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900939.

Petkovic:1986:SHO

[PMS86] M. S. Petković, G. V. Milovanović, and L. V. Stefanović. Some higher-order methods for the simultaneous approximation of multiple polynomial zeros. Computers and Mathematics with Applications, 12 (Part A)(9):951–962, September 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900209.

Peggion:1985:EFD

[PO85] Germana Peggion and James J.

O’Brien. An explicit finitedifference scheme for solving the ocean acoustic parabolic wave equation. Computers and Mathematics with Applications, 11(9): 937–942, September 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900975.

Pourbabai:1987:MQS

[Pou87] B. Pourbabai. Markovian queueing systems with retrials and heterogeneous servers. Computers and Mathematics with Applications, 13(12):917–923, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187900642.

Pourbabai:1988:TBM

[Pou88] B. Pourbabai. Tandem behavior of an M/M/1/N → G/M/1 queuing system. Computers and Mathematics with Applications, 16(3):215–220, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901812.

Pourbabai:1989:ACA

[Pou89] B. Pourbabai. An assembly cell with an automated quality control station: a finite capacity and generally distributed processing times. Computers and Mathematics with Applications, 18(4):321–328, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902174.

Petkovic:1989:CTE

[PP89] M. S. Petković and L. D. Petković. A computational test for the existence of polynomial zero. Computers and Mathematics with Applications, 17(7): 1109–1114, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900412.

Parisi:1989:SBP

[PPR^{+}89] J. Parisi, J. Peinke, B. Röhricht, K. M. Mayer, and U. Rau. Symmetry-breaking pattern formation in semiconductor physics: Spatio-temporal current structures during avalanche breakdown. Computers and Mathematics with Applications, 17(1– 3):467–473, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901752.

Phillips:1985:FDS

[PR85] Timothy N. Phillips and Milton E. Rose. A finite difference scheme for a class of firstorder elliptic partial differential equations. Computers and Mathematics with Applications, 11 (4):411–417, April 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900628.

Pan:1986:FEL

[PR86a] V. Pan and J. Reif. Fast and efficient linear programming and linear least-squares computations. Computers and Mathematics with Applications, 12 (Part A) (12):1217–1227, December 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900064.

Parihar:1986:PCI

[PR86b] K. S. Parihar and M. P. Ramachandran. Piecewise cubic interpolatory polynomials and approximate solution of singular integral equations. Computers and Mathematics with Applications, 12 (Part A)(12):1201– 1215, December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900052.

Prasad:1987:APE

[PR87] U. R. Prasad and N. Rajan. Aircraft pursuit–evasion problems with variable speeds. Computers and Mathematics with Applications, 13(1–3):111–121, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122187900976.

Pan:1989:FEPb

[PR89] V. Pan and J. Reif. Fast and efficient parallel solution of dense linear systems. Computers and Mathematics with Applications, 17(11):1481–1491, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900813.

Prasad:1980:VPM

[Pra80] B. Prasad. Variable penalty methods for constrained minimization. Computers and Mathematics with Applications, 6 (1):79–97, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900620.

Prasad:1982:HVP

[Pra82] B. Prasad. A hybrid variable penalty method for nonlinear programming. Computers and Mathematics with Applications, 8(6):413–424, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900177.

Pruss:1983:QBP

[Prü83] Jan Prüss. On the qualitative behaviour of populations with agespecific interactions. Computers and Mathematics with Applications, 9(3):327–339, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900202.

Petkovic:1984:NSG

[PS84] Miodrag S. Petković and Lidija V.

Stefanović. The numerical stability of the generalized root iterations for polynomial zeros. Computers and Mathematics with Applications, 10(2):97–106, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122184900415.

Pavlovic:1986:SAL

[PT86] B. Pavlović and N. Trinajstić. On symmetry and asymmetry in literature. Computers and Mathematics with Applications, 12 (Part B)(1–2):197– 227, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901513.

Puglisi:1989:PAA

[Pug89] C. Puglisi. Parallel algorithms and architectures for matrix multiplication. Computers and Mathematics with Applications, 17(12):1567–1572, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900564.

Purdy:1983:BAS

[Pur83a] George B. Purdy. A binary algorithm for solving n by n systems of integer equations. Computers and Mathematics with Applications, 9(6):715–719, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183901360.

Purdy:1983:CFA

[Pur83b] George B. Purdy. A carry-free algorithm for finding the greatest common divisor of two integers. Computers and Mathematics with Applications, 9(2): 311–316, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901335.

Purdy’s extended binary algorithm for the gcd runs in O(n^{2}) time, where n is the number of bits in the inputs; see [BB87] for an O(n) algorithm.

Pachter:1989:P

[PY89] M. Pachter and Y. Yavin. Preface. Computers and Mathematics with Applications, 18 (1–3):vii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901193.

Qin:1984:CNC

[Qin84] Huashu Qin. On the controllability of a nonlinear control system. Computers and Mathematics with Applications, 10 (6):441–451, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900749.

Qin:1986:SSC

[Qin86] H. S. Qin. Some sufficient conditions for occurring bifurcation. Computers and Mathematics with Applications, 12 (Part A)

(12):1171–1182, December 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900027.

See note [Van87].

Rabinowitz:1986:SGK

[Rab86]  P. Rabinowitz.  A stable

Gauss–Kronrod algorithm for Cauchy principal-value integrals. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1249–1254, September/ December 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/089812218690249X.

Rajan:1989:TOC

[Raj89] N. Rajan. Two-on-one combat games. Computers and Mathematics with Applications, 18(1– 3):107–116, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901284.

Rall:1987:ISC

[Ral87] L. B. Rall. An introduction to the scientific computing language Pascal-SC. Computers and Mathematics with Applications, 14(1):53–69, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901817.

Ranai:1987:SPI

[Ran87] K. Ranai. Simulation performance of information placement strategies. Computers and Mathematics with Applications, 14 (8):645–654, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901891.

Rand:1988:HVN

[Ran88] O. Rand. Harmonic variables — a new approach to nonlinear periodic problems. Computers and Mathematics with Applications, 15(11):953–961, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900405.

Rao:1984:CHC

[Rao84] T. Mahadeva Rao. Conversion of Hensel codes to rational numbers. Computers and Mathematics with Applications, 10(2):185–189, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900488.

Rascle:1985:RPN

[Ras85] Michel Rascle. The Riemann problem for a nonlinear nonstrictly hyperbolic system arising in biology. Computers and Mathematics with Applications, 11 (1–3):223–238, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122185901488.

Reiter:1983:SRI

[RC83] Raymond Reiter and Giovanni Criscuolo. Some representational issues in default reasoning. Computers and Mathematics with Applications, 9(1):15–27, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900044.

Reghbati:1987:F

[Reg87] H. K. Reghbati. Foreword. Computers and Mathematics with Applications, 13(5–6):vii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218790071X.

Reiman:1981:CAC

[Rei81] Allan Reiman. Computer-aided closure of the Lie algebra associated with a nonlinear partial differential equation. Computers and Mathematics with Applications, 7(5):387–393, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901231.

Reid:1987:NFN

[Rei87] J. G. Reid. Normal functions of normal random variables. Computers and Mathematics with Applications, 14(3):157–160, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901477.

Rosen:1989:RSL

[RI89] A. Rosen and A. Isser. Recent studies on the lifting surface theory. Computers and Mathematics with Applications, 17(7): 1115–1130, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900424.

Rice:1981:IPP

[Ric81] Bart Rice. The 0–1 Integer Programming Problem in a finite ring with identity. Computers and Mathematics with Applications, 7(6):497–502, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218190033X.

Richman:1988:CTA

[Ric88] F. Richman. Computers, trees and Abelian groups. Computers and Mathematics with Applications, 15(4):333–336, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902209.

Rigney:1987:IRM

[Rig87] D. R. Rigney. Inherited rate model of the cell cycle: Kinetics of related cells, EPI-genetics of ribosomal DNA transcription and the evaluation of cancertherapy fractionation schedules and doses. Computers and Mathematics with Applications, 14(9– 12):699–739, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902227.

Rivlin:1986:ORM

[Riv86]  T. J. Rivlin.  Optimal re-

covery of multivariate contractions. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1255–1258, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902506.

Ranai:1988:TRA

[RL88] K. Ranai and P. H. Leng. Topological reliability analysis of local-area computer networks. Computers and Mathematics with Applications, 15 (2):119–129, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122188900818.

Rodin:1987:AIA

[RLM^{+}87] E. Y. Rodin, Y. Lirov, S. Mittnik, B. G. McElhaney, and L. Wilbur. Artificial intelligence in air combat games. Computers and Mathematics with Applications, 13(1–3):261–274, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218790109X.

Robins:1987:FCE

[RM87] J. M. Robins and H. Morgenstern. The foundations of confounding in epidemiology. Computers and Mathematics with Applications, 14(9– 12):869–916, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902367.

See errata [Ano89o].

Robert:1986:MIL

[Rob86] F. Robert. Machine iteration for a linear function. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1259–1274, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902518.

Robins:1987:ANA

[Rob87a] J. M. Robins. Addendum to “A new approach to causal inference in mortality studies with a sustained exposure period — application to control of the healthy worker survivor effect”. Computers and Mathematics with Applications, 14(9–12):923–945, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902380.

See errata [Ano89n].

Robins:1987:ENA

[Rob87b] J. M. Robins. Errata to “A new approach to causal intefence in mortality studies with a sustained exposure period — application to control of the healthy worker survivor effect” mathl modelling 7(9-12), 1393–1512 (1986). Computers and Mathematics with Applications, 14(9– 12):917–921, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902379.

Rockenbauer:1988:IUP

[Roc88] A. Rockenbauer. Invariance and the uncertainty principle in EPR spectroscopy of crystals and liquids. Computers and Mathematics with Applications, 16(5– 8):629–635, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122188902519.

Rodin:1984:BRa

[Rod84a] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 10 (2):191–205, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218490049X.

Rodin:1984:BRb

[Rod84b] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 10 (3):305–314, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900580.

Rodin:1984:BRc

[Rod84c] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 10(4– 5):399–404, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900683.

Rodin:1985:BRa

[Rod85a] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 11 (4):419–420, April 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218590063X.

Rodin:1985:BRb

[Rod85b] Ervin Y. Rodin. Book reports. Computers and Mathematics with Applications, 11(9): 983–986, September 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901014.

Rodin:1986:F

[Rod86] Ervin Y. Rodin. Foreword. Computers and Mathematics with Applications, 12 (Part B)(1–2):xi– xii, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901331.

Rodin:1987:Fa

[Rod87a] E. Y. Rodin. Foreword. Computers and Mathematics with Applications, 14(1):vii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901751.

Rodin:1987:Fb

[Rod87b] E. Y. Rodin. Foreword. Computers and Mathematics with Applications, 14(5):vii, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122187902057.

Rodin:1987:LE

[Rod87c] E. Y. Rodin. Letters to the editor. Computers and Mathematics with Applications, 14 (8):655–656, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901908.

Rodin:1987:PEB

[Rod87d] E. Y. Rodin. A pursuit–evasion bibliography — version 1. Computers and Mathematics with Applications, 13(1–3):275–340, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901106.

Rodin:1988:IBR

[Rod88a] E. Y. Rodin. The importance of being Robert Maxwell: an appraisal. Computers and Mathematics with Applications, 15 (5):v–vi, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890154X.

Rodin:1988:LE

[Rod88b] Ervin Y. Rodin. Letter from the editor. Computers and Mathematics with Applications, 16(9):769, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900132.

Rodin:1989:PEB

[Rod89a] E. Y. Rodin. A pursuit–evasion bibliography — version 2. Computers and Mathematics with Applications, 18(1–3):245–320, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189901399.

Rodin:1989:SSS

[Rod89b] Ervin Y. Rodin. Software survey section. Computers and Mathematics with Applications, 17(11):i–ii, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900850.

Rohdin:1983:NMM

[Roh83] H. Rohdin. Numerical modeling of a moving boundary diffusion/drift problem. Computers and Mathematics with Applications, 9(2):293–305, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183901311.

Rokhlin:1985:AVI

[Rok85] Vladimir Rokhlin. Application of volume integrals to the solution of partial differential equations. Computers and Mathematics with Applications, 11(7– 8):667–679, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901634.

Rose:1980:NMF

[Ros80] Michael E. Rose. Numerical methods for flows through porous media — II. Computers and Mathematics with Applications, 6(1):99–122, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900632.

Rossi:1987:GSA

[Ros87] C. Rossi. General software for the analysis of cytofluorimetric data by E-M methods. Computers and Mathematics with Applications, 14(9–12):771–782, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902264.

Rosen:1989:SFS

[Ros89] J. Rosen. Symmetry at the foundations of science. Computers and Mathematics with Applications, 17(1–3):13–15, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901442.

Roth:1986:QTO

[Rot86]  Robert S. Roth.  Quasilin-

earization and terminal observation. Computers and Mathematics with Applications, 12 (Part A)(6):831–838, June 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900684.

Rozsa:1986:SMA

[Róz86] Erzsébet Rózsa. Symmetry in Muslim arts. Computers and Mathematics with Applications, 12(3–4):725–750, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186904207.

Ranai:1987:OFA

[RPL87] K. Ranai, C. Pursglove, and P. H. Leng. Optimal file allocation in a loop topology local area network. Computers and Mathematics with Applications, 13(7):601–610, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901210.

Rao:1988:MCA

[RR88a] S. V. R. Rao and J. A. Rasmussen. Micro-computer applications in water resource engineering. Computers and Mathematics with Applications, 16 (9):677–682, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890003X.

Rao:1988:SMW

[RR88b] S. V. R. Rao and J. A. Rasmussen. Stochastic models in water resource engineering utilizing micro-computers. Computers and Mathematics with Applications, 16(9):671–676, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900028.

Rozsondai:1989:SAB

[RR89] M. Rozsondai and B. Rozsondai.

Symmetry aspects of bookbindings. Computers and Mathematics with Applications, 17(4– 6):837–885, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902642.

Rodabaugh:1981:NRE

[RT81] D. J. Rodabaugh and S. Thompson. A note on the relative efficiency of Adams methods. Computers and Mathematics with Applications, 7(5):401–403, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122181901255.

Ronto:1988:SSB

[RTF^{+}88] Gy. Rontó, K. Tóth, L. A. Feigin, D. I. Svergun, and A. T. Dembo. Symmetry and structure of bacteriophage t7. Computers and Mathematics with Applications, 16(5–8):617–628, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902507.

Rudnev:1988:AER

[Rud88]  S. V. Rudnev.  Application

of elliptic Reimannian geometry to problems of crystallography. Computers and Mathematics with Applications, 16(5– 8):597–616, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902490.

Rustem:1989:SCC

[Rus89] B. Rustem. A superlinearly convergent constrained min-max algorithm for rival models of the same system. Computers and Mathematics with Applications, 17(8–9):1305–1316, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900977.

Rao:1987:MCA

[RW87] S. V. R. Rao and K. F. Willis. Micro-computer applications in environmental engineering. Computers and Mathematics with Applications, 14(3):153–156, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901465.

Rodriguez:1986:SME

[RZ86] R. Rodríguez and P. E. Zadunaisky.

A stable method to estimate perturbations in differential equations. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1275–1286, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690252X.

Somani:1987:DHF

[SA87] A. K. Somani and V. K. Agarwal.

Diagnosis in hybrid fault situations under aim and a unified tcharacterization theorem. Computers and Mathematics with Applications, 13(5–6):567–576, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187900848.

Salzer:1984:CGF

[Sal84a]  Herbert E. Salzer.  Conflu-

ent general-factorial difference and derivative operators and polynomial-exponential interpolation. Computers and Mathematics with Applications, 10 (6):433–439, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122184900737.

Salzer:1984:GLL

[Sal84b] Herbert E. Salzer. A generalization of Layman–Lotockiĭ exponential interpolation. Computers and Mathematics with Applications, 10(1):71–75, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122184900877.

Sander:1986:BSI

[San86] Klaus Sander. Bilateral symmetry in insects: Could it derive from circular asymmetries during early embryogenesis? Computers and Mathematics with Applications, 12 (Part B)(1–2):413– 418, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901628.

Sarafyan:1984:CAS

[Sar84] Diran Sarafyan. Continuous approximate solution of ordinary differential equations and their systems. Computers and Mathematics with Applications, 10 (2):139–159, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122184900440.

Sarafyan:1988:FCA

[Sar88] D. Sarafyan. Families of continuous approximate processes for the solution of ordinary differential equations. Computers and Mathematics with Applications, 16(10–11):887–894, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901988.

Satty:1985:PDE

[Sat85] ThomasL. Satty. Partial differential equations: an outline. Computers and Mathematics with Applications, 11(1–3):1– 4, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218590135X.

Sridharan:1985:KSP

[SB85] N. S. Sridharan and J. L.

Bresina. Knowledge structures for planning in realistic domains. Computers and Mathematics with Applications, 11 (5):457–480, May 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900513.

Safwat:1988:PGC

[SB88] K. N. A. Safwat and B. D. Brademeyer. Proof of global convergence of an efficient algorithm for predicting trip generation, trip distribution, modal split and traffic assignment simultaneously on large-scale networks. Computers and Mathematics with Applications, 16(4):269–277, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901435.

Shampine:1983:OEC

[SBB83] L. F. Shampine, L. S. Baca, and H.-J. Bauer. Output in extrapolation codes. Computers and Mathematics with Applications, 9(2):245–255, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901281.

Stolevik:1986:SPG

[SBCC86] Reidar Stølevik, Jon Brunvoll, Bjørg Cyvin, and Sven Cyvin. Strange patterns generated from simple rules with possible relevance to crystallization and other phenomena. Computers and Mathematics with Applications, 12(3–4):789–801, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186904256.

Sakhnenko:1988:SMS

[SC88] V. P. Sakhnenko and G. M.

Chechin. Symmetry methods and space group representations in the theory of phase transitions. Computers and Mathematics with Applications, 16(5– 8):453–464, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902350.

Schultz:1985:MAP

[Sch85] Martin H. Schultz. Multiple array processors for ocean acoustic problems. Computers and Mathematics with Applications, 11(7– 8):777–785, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901725.

Scharfenberg:1986:SGA

[Sch86a] Peter Scharfenberg. On symmetries in the graphic art of Horst Barting. Computers and Mathematics with Applications, 12(3– 4):883–893, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904323.

Schattschneider:1986:BWH

[Sch86b] Doris Schattschneider. In black and white: how to create perfectly colored symmetric patterns. Computers and Mathematics with Applications, 12(3– 4):673–695, May/August 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904189.

Schneider:1986:STA

[Sch86c] W. Schneider. A simple technique for adding tension to parabolic blending interpolation. Computers and Mathematics with Applications, 12 (Part A) (11):1155–1160, November 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900192.

Shannon:1987:CRI

[SCH87] A. G. Shannon, J. H. Clarke, and L. J. Hills. Contingency relations for infectious diseases. Computers and Mathematics with Applications, 14(9–12):829–833, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902318.

Sticklen:1985:MMM

[SCSS85] Jon Sticklen, B. Chadrasekaran, J. W. Smith, and John Svirbely. MDX–MYCIN: the MDX paradigm applied to the mycin domain. Computers and Mathematics with Applications, 11 (5):527–539, May 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900550.

Scull:1989:BCG

[Scu89] S. C. Scull. Bézier curves on groups. Computers and Mathematics with Applications, 18 (8):691–700, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902265.

Shinar:1987:TTG

[SD87] J. Shinar and A. Davidovitz. A two-target game analysis in lineof-sight coordinates. Computers and Mathematics with Applications, 13(1–3):123–140, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187900988.

Shukla:1989:SBX

[SDB89] R. Shukla, J. A. Deddens, and C. R. Buncher. Survival benefits of X-ray screening for lung cancer after bias adjustments. Computers and Mathematics with Applications, 18(10–11):937–948, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189900126.

Schvaneveldt:1988:GTF

[SDD88] R. W. Schvaneveldt, D. W. Dearholt, and F. T. Durso. Graph theoretic foundations of pathfinder networks. Computers and Mathematics with Applications, 15(4):337–345, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902210.

Sacks-Davis:1981:TIO

[SDS81] R. Sacks-Davis and L. F. Shampine. A type-insensitive ODE code based on second derivative formulas. Computers and Mathematics with Applications, 7(6):487–495, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900328.

Speziale:1981:NFM

[SE81] Charles G. Speziale and A. C. Eringen. Nonlocal fluid mechanics description of wall turbulence. Computers and Mathematics with Applications, 7 (1):27–41, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900055.

Sahimi:1988:GEM

[SE88a] M. S. Sahimi and D. J. Evans. Group explicit methods for the numerical solution of the wave equation. Computers and Mathematics with Applications, 15(6– 8):699–709, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902891.

Sanugi:1988:NFO

[SE88b] B. B. Sanugi and D. J. Evans. A new fourth order Runge–Kutta formula for y ’ = ay with stepsize control. Computers and Mathematics with Applications, 15(12):991–995, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901344.

Senechal:1986:GCS

[Sen86]  Marjorie Senechal.  Geom-

etry and crystal symmetry. Computers and Mathematics with Applications, 12(3–4):565– 578, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904116.

Sen:1988:RID

[Sen88a] M. De La Sen. A robust indirect discrete adaptive-control approach based on passivity results for nonlinear systems. Computers and Mathematics with Applications, 15(5):389–403, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188901605.

Senechal:1988:CS

[Sen88b] M. Senechal. Color symmetry. Computers and Mathematics with Applications, 16(5– 8):545–553, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902441.

Senechal:1989:SR

[Sen89a] M. Senechal. Symmetry revisited. Computers and Mathematics with Applications, 17 (1–3):1–12, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901430.

Sengupta:1989:DVP

[Sen89b] J. K. Sengupta. A dynamic view of the portfolio efficiency frontier. Computers and Mathematics with Applications, 18(6– 7):565–580, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901089.

Senior:1989:SWN

[Sen89c] A. Senior. Symmetry and wayward nature. Computers and Mathematics with Applications, 17(4–6):913–919, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990268X.

Serbin:1980:CAA

[Ser80] Steven M. Serbin. On the construction and analysis of approximations of arbitrarily high-order for proportionally-damped second order systems. Computers and Mathematics with Applications, 6(2):251–263, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900334.

Seydel:1987:NMC

[Sey87] R. Seydel. New methods for calculating the stability of periodic solutions. Computers and Mathematics with Applications, 14(7):505–510, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900459.

Schroeder:1988:LAN

[SF88] L. Schroeder and S. Fleeter. Locally analytic numerical method and application to viscous internal and external flows. Computers and Mathematics with Applications, 15(11):915–932, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188900351.

Saluja:1987:TDP

[SFK87] K. K. Saluja, H. Fujiwara, and K. Kinoshita. A testable design of programmable logic arrays with universal control and minimal overhead. Computers and Mathematics with Applications, 13(5–6):503–517, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900794.

Shinar:1989:OGL

[SGG89] J. Shinar, M. Guelman, and A. Green. An optimal guidance law for a planar pursuit–evasion game of kind. Computers and Mathematics with Applications, 18(1–3):35–44, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901223.

Shieh:1987:TCT

[SGN87] L. S. Shieh, S. Ganesan, and J. M. Navarro. Transformations of a class of time-varying multivariable control systems to block companion forms. Computers and Mathematics with Applications, 14(6):471–477, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900599.

Sharir:1981:SCA

[Sha81a] M. Sharir. A strong-connectivity algorithm and its applications in data flow analysis. Computers and Mathematics with Applications, 7(1):67–72, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900080.

Sharir:1981:SSM

[Sha81b] Micha Sharir. A strange sorting method inspired by formal differentiation. Computers and Mathematics with Applications, 7(4):293–298, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900584.

Shayman:1982:BRG

[Sha82] Mark A. Shayman. Book review:

Geometrical methods for the theory of linear systems: Edited by C. I. Byrnes and C. F. Martin. D. Reidel, Dordrecht, 1980. $39.50. Computers and Mathematics with Applications, 8(3): 241–242, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900487.

Shampine:1984:SER

[Sha84]  L. F. Shampine.  Stability

of explicit Runge–Kutta methods. Computers and Mathematics with Applications, 10(6): 419–432, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900725.

Shafranovskii:1986:MRT

[Sha86a] I. I. Shafranovskii. Magic rectangles in the table of the 32 classes of symmetry. Computers and Mathematics with Applications, 12 (Part B)(1–2):93– 96, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901422.

Shampine:1986:CLN

[Sha86b] L. F. Shampine. Conservation laws and the numerical solution of ODEs. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1287–1296, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902531.

Shampine:1986:GEE

[Sha86c] L. F. Shampine. Global error estimation with one-step methods. Computers and Mathematics with Applications, 12 (Part A)(7):885–894, July 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900325.

Shampine:1987:IVO

[Sha87] L. F. Shampine. Interpolation for variable order, Runge–Kutta methods. Computers and Mathematics with Applications, 14 (4):255–260, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901337.

Shah:1988:SMN

[Sha88] S. A. Shah. P -stable modification of the Numerov method. Computers and Mathematics with Applications, 16(12):1041– 1043, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890260X.

Shonkwiler:1980:ADB

[Sho80]  Ron Shonkwiler.  On age-

dependent branching processes with immigration. Computers and Mathematics with Applications, 6(3):289–296, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218090036X.

Sridhar:1987:ASP

[SHPT87] T. Sridhar, D. S. Ho, T. J. Powell, and S. M. Thatte. Analysis and simulation of parallel signature analyzers. Computers and Mathematics with Applications, 13(5–6):537–545, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900812.

Shubnikov:1988:AT

[Shu88a] A. V. Shubnikov. Antisymmetry of textures. Computers and Mathematics with Applications, 16(5–8):373–377, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902271.

Shubnikov:1988:WPC

[Shu88b] A. V. Shubnikov. On the works of Pierre Curie on symmetry. Computers and Mathematics with Applications, 16(5–8):357–364, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902258.

Shubnikov:1988:SS

[Shu88c] A. V. Shubnikov. Symmetry of similarity. Computers and Mathematics with Applications, 16(5– 8):365–371, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890226X.

Sheshinski:1989:CBA

[SI89] E. Sheshinski and M. D. Intriligator. Cost-benefit analysis with switching regimes: an application of the theory of planning. Computers and Mathematics with Applications, 17(8– 9):1317–1327, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900989.

Sidner:1983:WSM

[Sid83] Candace L. Sidner. What the speaker means: the recognition of speakers’ plans in discourse. Computers and Mathematics with Applications, 9 (1):71–82, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900081.

Sinestrari:1983:LGS

[Sin83] Eugenio Sinestrari. Local and global stability for the solutions of a nonlinear renewal equation. Computers and Mathematics with Applications, 9(3): 353–359, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900238.

Schmidt:1985:ENS

[SJ85] Henrik Schmidt and Finn B.

Jensen. Efficient numerical solution technique for wave propagation in horizontally stratified environments. Computers and Mathematics with Applications, 11(7–8):699–715, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218590166X.

Schultz:1985:F

[SL85a] Martin H. Schultz and Ding Lee. Foreword. Computers and Mathematics with Applications, 11(7– 8):vii, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901580.

Siegmann:1985:ATD

[SL85b] William L. Siegmann and Ding Lee. Aspects of threedimensional parabolic equation computations. Computers and Mathematics with Applications, 11(7–8):853–862, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122185901804.

Sugiyama:1988:SAP

[SL88] H. Sugiyama and E. S. Lee. Stochastic approximation — a powerful method for solving deterministic numerical problems. Computers and Mathematics with Applications, 15 (11):963–967, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900417.

Slanina:1986:IBP

[Sla86] Zdenek Slanina. An interplay between the phenomenon of chemical isomerism and symmetry requirements: a perennial source of stimuli for molecular-structure concepts, as well as for algebraic and computational chemistry. Computers and Mathematics with Applications, 12(3–4): 585–616, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218690413X.

Shyi:1989:FMR

[SLHC89] C.-N. Shyi, J.-Y. Lee, K.-C.

Hung, and C.-H. Chen. A fast method for robot location determination. Computers and Mathematics with Applications, 18 (9):771–777, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901764.

Shieh:1987:FSR

[SLPC87] L. S. Shieh, S. R. Lian, C. B.

Park, and N. P. Coleman. Fast and stable recursive algorithms for continuous-time and discretetime model conversions. Computers and Mathematics with Applications, 14(8):629–644, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812218790188X.

Scobelev:1980:SAN

[SM80] B. Yu. Scobelev and Yu. I. Molorodov. Subcritical autooscillations and nonlinear neutral curve for Poiseuille flow. Computers and Mathematics with Applications, 6(1): 123–133, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900644.

Sinha:1988:TIR

[SM88a] P. C. Sinha and A. K. Mitra. Tidally induced residual circulation. Computers and Mathematics with Applications, 16(1– 2):153–167, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900296.

Staniforth:1988:OMR

[SM88b] A. Staniforth and J. Mailhot. An operational model for regional weather forecasting. Computers and Mathematics with Applications, 16(1–2):1–22, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/089812218890020X.

Schultz:1987:CSP

[SMH87] F. W. Schultz, A. C. M. Martens, and A. Hagenbeek. Computer simulation of the progression of an acute myelocytic leukemia in the brown Norway rat. Computers and Mathematics with Applications, 14(9–12):751–761, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187902240.

Smith:1983:OPJ

[Smi83] A. C. Smith. One-parameter Jacobi transformations of sequences. Computers and Mathematics with Applications, 9(6): 721–733, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901372.

Smith:1989:CSM

[Smi89] H. L. Smith. Competing subcommunities of mutualists. Computers and Mathematics with Applications, 18(10–11):949–957, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189900138.

Sniedovich:1988:CPN

[Sni88] M. Sniedovich. C-programming:

a nonlinear parametric optimization method. Computers and Mathematics with Applications, 16(10–11):837–844, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901939.

Snyman:1989:CDM

[Sny89] J. A. Snyman. A convergent dynamic method for large minimization problems. Computers and Mathematics with Applications, 17(10):1369–1377, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189900205.

Socolovsky:1988:LNO

[Soc88] E. A. Socolovsky. Lagrangian non-oscillatory and FEM schemes for the porous media equation. Computers and Mathematics with Applications, 15(6– 8):611–617, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902842.

Sod:1985:RCM

[Sod85]  Gary A. Sod.  A random

choice method with application to reaction-diffusion systems in combustion. Computers and Mathematics with Applications, 11(1–3):129–144, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901427.

Sod:1987:DAR

[Sod87] G. A. Sod. A dictionary approach to reaction-diffusion systems with nonlinear diffusion coefficients. Computers and Mathematics with Applications, 13(9– 11):771–783, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901635.

Sokolov:1986:CWS

[Sok86] V. I. Sokolov. Chirality in the world of stereochemistry. Computers and Mathematics with Applications, 12(3–4):547– 563, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904104.

Sowa:1983:GLC

[Sow83]  John F. Sowa.  Generat-

ing language from conceptual graphs. Computers and Mathematics with Applications, 9 (1):29–43, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900056.

Shih:1989:FBI

[SP89] K. G. Shih and A. N. Palazotto. Formulation of boundary integral equations for linear elastic shells. Computers and Mathematics with Applications, 18(12):985–999, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900278.

Spanos:1980:NSV

[Spa80] P-T. D. Spanos. Numerical simulations of a van der Pol oscillator. Computers and Mathematics with Applications, 6(1): 135–145, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900656.

Srivastava:1986:NER

[SPF86] A. Srivastava and S. L. Paveri-Fontana. Numerical experiments on the rational Runge– Kutta method. Computers and Mathematics with Applications, 12 (Part A)(12):1161– 1170, December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900015.

Squire:1981:DLQ

[Squ81] William Squire. In defense of linear quadrature rules. Computers and Mathematics with Applications, 7(2):147–149, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901139.

Subramanian:1987:CCF

[SRP87] G. Subramanian, C. V. Raghava Rao, and K. Pramadavalli. A circular cylinder in a flow field with parabolic velocity distribution — a numerical study. Computers and Mathematics with Applications, 14(2):127–132, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901428.

Skowronski:1987:BPE

[SS87] J. M. Skowronski and R. J.

Stonier. The barrier in a pursuit– evasion game with two targets. Computers and Mathematics with Applications, 13(1– 3):37–45, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900927.

Shafranovskii:1988:LCS

[SS88] I. I. Shafranovskii and G. I.

Shafranovskii. The law of correlation of symmetry and statistics in the mineral world. Computers and Mathematics with Applications, 16(5–8):593–596, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188902489.

Skowronski:1989:TPQ

[SS89] J. M. Skowronski and R. J. Stonier. Two-person qualitative differential games with two objectives. Computers and Mathematics with Applications, 18(1– 3):133–150, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901314.

Sambandham:1988:ANT

[SSB88] M. Sambandham, T. S. Srivatsan, and K. B. Bota. Approximation and numerical treatment of an autoregressive equation with stochastic coefficients. Computers and Mathematics with Applications, 16(10–11):921–927, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188902027.

Saluja:1987:SAT

[SSS87] K. K. Saluja, L. Shen, and S. Y. H. Su. A simplified algorithm for testing microprocessors. Computers and Mathematics with Applications, 13(5– 6):431–441, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122187900733.

StMary:1985:AIF

[St.85] D. F. St. Mary. Analysis of an implicit finite-difference scheme for a third-order partial differential equation in three dimensions. Computers and Mathematics with Applications, 11(7– 8):873–885, July/August 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901828.

Schuhmacher:1988:STN

[ST88] K. Schuhmacher and H. Thieme.

Some theoretical and numerical aspects of modelling dispersion in the development of ectotherms. Computers and Mathematics with Applications, 15(6– 8):565–594, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902817.

Stakhov:1989:GSM

[Sta89] A. P. Stakhov. The golden section in the measurement theory. Computers and Mathematics with Applications, 17(4– 6):613–638, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902526.

Sambandham:1988:NSR

[STB88] M. Sambandham, V. Thangaraj, and K. B. Bota. Numerical solution of random Love’s integral equation. Computers and Mathematics with Applications, 16(10– 11):915–920, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902015.

Steinberg:1981:BRA

[Ste81a] David I. Steinberg. Book review: Algorithms for network programming: By Jeff L. Kennington and Richard V. Helgason. Wiley–Interscience, New York 1980. 291 pp., $23.50. Computers and Mathematics with Applications, 7(4):350, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900651.

Steinberg:1981:BRC

[Ste81b] David I. Steinberg. Book review: Computer modelling in agriculture: By N. R. Brockington. Clarendon Press, Oxford, England 1979. 156pp. (£8.95 net in U.K.). Computers and Mathematics with Applications, 7(4):349, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900638.

Steinberg:1981:BRH

[Ste81c] David I. Steinberg. Book review:

Home computers: 2^{10} questions and answers. vol. 1: hardware and vol. 2: software: By Rich Didday, Dilithium Press (Distributed by ISBS, Inc., P.O. Box 555, Forest Grove, OR 97116, U.S.A.) 1977. Vol. 1, 255pp + appendices, $7.95, Vol. 2, 175pp + appendices, $6.95, Both Soft Cover. Computers and Mathematics with Applications, 7(4): 350–351, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181900663.

Steinberg:1981:BRO

[Ste81d] David I. Steinberg. Book review:

Optimal strategies in sports: Edited by Shaul P. Ladany and Robert E. Machol. North-Holland, Amsterdam 1977. 231 pp. $20 (Dfl. 50.00). Computers and Mathematics with Applications, 7(4):349–350, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218190064X.

Steinberg:1982:BRP

[Ste82] David I. Steinberg. Book review: Proceedings of the International Symposium on Optimum Structural Design: (The 11th Naval Structural Mechanics Symposium). 587 pp. Price: $40, postage paid (U.S. only); $47 prepaid, surface mail (Foreign). Office of Special Professional Education, Harvill Building, University of Arizona, Tucson, AZ 85721. Computers and Mathematics with Applications, 8(6):469, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122182900219.

Stefanek:1987:BRH

[Ste87]  G. Stefanek.  Book review:

The handbook of artificial intelligence: Vol. 1. by A. Barr and E. A. Feigenbaum. Kaufman, Calif. (1981). 409 pp. Computers and Mathematics with Applications, 14(9–12):947, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187902392.

Steppeler:1988:GFE

[Ste88] J. Steppeler. A Galerkin finite element-spectral weather forecast model in hybrid coordinates. Computers and Mathematics with Applications, 16 (1–2):23–30, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900211.

Singh:1987:FDM

[STM87a] H. P. Singh, D. D. Tripathi, and R. B. Mishra. Finitedifference methods for boundary value problems at high grashof number. Computers and Mathematics with Applications, 14 (7):511–518, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900460.

Singh:1987:SPH

[STM87b] H. P. Singh, D. D. Tripathi, and R. B. Mishra. Structure of a point heat source near an interface. Computers and Mathematics with Applications, 14 (8):623–627, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901878.

Singh:1987:SMR

[STM87c] H. P. Singh, D. D. Tripathi, and R. B. Mishra. The structure of magnetized rotating polytropes. Computers and Mathematics with Applications, 14 (6):465–469, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900587.

Singh:1988:PSS

[STM88] H. P. Singh, D. D. Tripathi, and R. B. Mishra. Physical study of steady electromagnetofluiddynamic viscous flows. Computers and Mathematics with Applications, 15(3):161–168, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188901678.

Singh:1988:DCP

[STP88] H. P. Singh, D. D. Tripathi, and H. K. Pandey. The development of circulation-preserving MHD flows via spherical mapping. Computers and Mathematics with Applications, 15 (5):347–350, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901551.

Strejc:1989:RES

[Str89] V. Strejc. Remarks on equilibrium state achievement in state space control. Computers and Mathematics with Applications, 18(6–7):643–649, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901144.

Shieh:1983:BTC

[STY83] Leang S. Shieh, Yih T. Tsay, and R. E. Yates. Block transformations of a class of multivariable control systems to four basic block companion forms. Computers and Mathematics with Applications, 9(6):703–714, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901359.

Seaman:1987:PGD

[STY87] J. W. Seaman, Jr., D. W. Turner, and D. M. Young. Polyhedron graphs for displaying multivariate data. Computers and Mathematics with Applications, 14 (4):269–277, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901350.

Smirnova:1988:FLC

[SU88] N. L. Smirnova and V. S. Urusov.

Fundamental law of crystal chemistry by Shubnikov, its applications and restrictions. Computers and Mathematics with Applications, 16(5–8):563–567, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902465.

Schultz:1989:THM

[SvDHH89] F. W. Schultz, J. J. M. van Dongen, K. Hählen, and A. Hagenbeek. Time-history of the malignant population in the peripheral blood of a child with t-cell acute lymphoblastic leukemia: a pilot study. Computers and Mathematics with Applications, 18(10– 11):929–936, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189900114.

Snyman:1988:CDG

[SvR88] J. A. Snyman and M. van Rooyen. A conjugate direction gradient method with reconnaissance steps for unconstrained minimization. Computers and Mathematics with Applications, 16(9):737–745, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900090.

Swart:1988:HBS

[Swa88] J. H. Swart. Hopf bifurcation and the stability of nonlinear age-dependent population models. Computers and Mathematics with Applications, 15(6– 8):555–564, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902805.

Shampine:1988:EIO

[SZ88] L. F. Shampine and W. Zhang. Efficient integration of ordinary differential equations by transformation. Computers and Mathematics with Applications, 15 (3):213–220, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188901733.

Shampine:1989:CLW

[SZ89] L. F. Shampine and W. Zhang.

Convergence of LMM when the solution is not smooth. Computers and Mathematics with Applications, 18(4):365–372, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189902216.

Szablowski:1987:GLL

[Sza87a] P. J. Szabłowski. Generalized laws of large numbers and auxiliary results concerning stochastic approximation with dependent disturbances — II. Computers and Mathematics with Applications, 13(12):973–987, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218790068X.

Szablowski:1987:SAD

[Sza87b] P. J. Szabłowski. Stochastic approximation with dependent disturbances — I. Computers and Mathematics with Applications, 13(12):951–972, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900678.

Szablowski:1988:ORP

[Sza88] P. J. Szabłowski. Optimal, recursive procedures of identification. Computers and Mathematics with Applications, 16 (3):229–246, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901836.

Szablowski:1989:CFT

[Sza89] P. J. Szabłowski. Can the first two conditional moments identify a mean square differentiable process? Computers and Mathematics with Applications, 18 (4):329–348, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902186.

Tandon:1989:SNT

[TA89] P. N. Tandon and R. Agarwal. A study of nutritional transport in a synovial joint. Computers and Mathematics with Applications, 17(7):1131–1141, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900436.

Takche:1988:CSM

[Tak88] J. Takche. Complexities of special matrix multiplication problems. Computers and Mathematics with Applications, 15 (12):977–989, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901332.

Tanaka:1989:CSS

[Tan89] K. Tanaka. The closest solution to the shadow minimum of a cooperative dynamic game. Computers and Mathematics with Applications, 18(1–3):181–188, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901351.

Tarjan:1988:CSI

[Tar88] I. Tarján. Cooperation with the Shubnikov Institute of the Soviet Academy of Sciences. Computers and Mathematics with Applications, 16(5–8):437–441, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188902337.

Tarnai:1989:BPS

[Tar89] T. Tarnai. Buckling patterns of shells and spherical honeycomb structures. Computers and Mathematics with Applications, 17(4– 6):639–652, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902538.

Toscani:1987:EBE

[TB87] G. Toscani and N. Bellomo. The Enskog–Boltzmann equation in the whole space R^{3}: Some global existence, uniqueness and stability results. Computers and Mathematics with Applications, 13(9– 11):851–859, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901684.

Theil:1989:ITL

[TC89] H. Theil and C.-F. Chung. The information tableau of a linear allocation model. Computers and Mathematics with Applications, 17(8–9):1265–1268, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900941.

Torabi:1988:SLP

[TD88] M. Torabi and J. A. Dracup. A sequential linear programming approach to solve mixed integer programming problems. Computers and Mathematics with Applications, 15(12):1029– 1039, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901381.

Telegdi:1989:AMV

[Tel89] L. Telegdi. Arrangements of minimal variance — multidimensional scaling in the symmetrical case. Computers and Mathematics with Applications, 17(1– 3):137–145, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901557.

Turner:1988:AFE

[TG88] J. C. Turner, Jr. and M. D. Gunzburger. An analysis of finite element approximations for an algebraic model of turbulence. Computers and Mathematics with Applications, 15(11):945–951, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900399.

Travis:1980:ESE

[TH80] C. C. Travis and Glen Haddock.

Estimating system equilibrium from response to an input of short duration. Computers and Mathematics with Applications, 6(2):247–249, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900322.

Thomas:1980:CEE

[Tho80] G. Thomas. The consistence error of extended variational principles in the finite element method. Computers and Mathematics with Applications, 6(4): 401–409, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900462.

Thompson:1986:EPP

[Tho86a] S. Thompson. The effect of partial pivoting in sparse ordinary differential equation solvers. Computers and Mathematics with Applications, 12 (Part A) (12):1183–1191, December 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900039.

Thompson:1986:RCS

[Tho86b] S. Thompson. Remarks on the choice of a stiff ordinary differential equation solver. Computers and Mathematics with Applications, 12 (Part A)(11):1125– 1141, November 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186900179.

Thompson:1987:SSC

[Tho87] S. Thompson. Stability of some continuously imbedded Runge– Kutta methods of sarafyan. Computers and Mathematics with Applications, 13(8):677– 680, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218790040X.

Thompson:1989:AME

[Tho89] J. R. Thompson. AIDS: the mismanagement of an epidemic.

Computers and Mathematics with Applications, 18(10–11): 965–972, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900151.

Tse:1989:ISMa

[TK89a] E. Tse and A. Khilnani. An integrated system model for a fishery management process — I: Model structure. Computers and Mathematics with Applications, 17(8– 9):1329–1343, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900990.

Tse:1989:ISMb

[TK89b] E. Tse and A. Khilnani. An integrated system model for a fishery management process — II. A case study. Computers and Mathematics with Applications, 18(6–7):675–690, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218990117X.

Tanikawa:1987:NLF

[TM87] A. Tanikawa and H. Mukai. New Lagrangian function for nonconvex primal-dual decomposition. Computers and Mathematics with Applications, 13 (8):661–676, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900393.

Tsokos:1987:PMB

[TO87] C. P. Tsokos and M. N. Oğuztöreli. A probabilistic model for breast cancer survival data. Computers and Mathematics with Applications, 14(9– 12):835–840, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218790232X.

Tompkins:1985:AAI

[Tom85] David H. Tompkins. Affine approximation of isoparametric elements. Computers and Mathematics with Applications, 11(12): 1211–1225, December 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901087.

Toth:1986:SIE

[Tót86] L. Fejes Tóth. Symmetry induced by economy. Computers and Mathematics with Applications, 12 (Part B)(1–2):83– 91, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901410.

Toussaint:1983:SAF

[Tou83] Godfried T. Toussaint. The symmetric all-furthest- neighbor problem. Computers and Mathematics with Applications, 9(6):747–754, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901396.

Takle:1988:AFE

[TR88] E. S. Takle and R. D. Russell.

Applications of the finite element method to modeling the atmospheric boundary layer. Computers and Mathematics with Applications, 16(1–2):57–68, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188900247.

Trojan:1980:TBC

[Tro80] J. M. Trojan. Tight bounds on the complexity index of onepoint iterations. Computers and Mathematics with Applications, 6(4):433–441, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900504.

Therneau:1989:MMD

[TSJ89] T. M. Therneau, L. A. Solberg, Jr., and R. B. Jenkins. Modeling megakaryocyte development as a branching process. Computers and Mathematics with Applications, 18(10–11):959–964, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812218990014X.

Tsai:1988:BPM

[TSS88] J. S. H. Tsai, L. S. Shieh, and T. T. C. Shen. Block power method for computing solvents and spectral factors of matrix polynomials. Computers and Mathematics with Applications, 16(9):683–699, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900041.

Tsai:1988:FSA

[TSY88] J. S. H. Tsai, L. S. Shieh, and R. E. Yates. Fast and stable algorithms for computing the principal n th root of a complex matrix and the matrix sector function. Computers and Mathematics with Applications, 15(11):903–913, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218890034X.

Thompson:1986:BFF

[TT86] S. Thompson and P. G. Tuttle. Benchmark fluid flow problems for continuous simulation languages. Computers and Mathematics with Applications, 12 (3):345–351, March 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901938.

Toth:1989:GAP

[TT89] G. Fejes Tóth and L. Fejes Tóth.

A geometrical analogue of the phase transformation of crystals. Computers and Mathematics with Applications, 17(1– 3):251–254, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901612.

Turner:1988:BLD

[Tur88] W. J. Turner. Better leak detection in gas pipelines. Computers and Mathematics with Applications, 15(1):69–75, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188900739.

Tuzhilin:1988:AMM

[Tuz88] A. Tuzhilin. Atomatics as mathematical modelling and knowledge representation: Applications to systems analysis in computer and information sciences. Computers and Mathematics with Applications, 15(9):723–742, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188901265.

Traub:1980:CCI

[TW80] J. F. Traub and H. Woźniakowski. Convergence and complexity of interpolatory-newton iteration in a Banach space. Computers and Mathematics with Applications, 6(4):385–400, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900450.

Usmani:1985:SET

[UA85] Riaz A. Usmani and Ravi P. Agarwal. An A-stable extended trapezoidal rule for the integration of ordinary differential equations. Computers and Mathematics with Applications, 11(12): 1183–1191, December 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901063.

Ubhaya:1983:LTA

[Ubh83] Vasant A. Ubhaya. Linear time algorithms for convex and monotone approximation. Computers and Mathematics with Applications, 9(4):633–643, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183901219.

Ubhaya:1984:ADP

[Ubh84] Vasant A. Ubhaya. O ( n )

algorithms for discrete n-point approximation by quasi-convex functions. Computers and Mathematics with Applications, 10(4– 5):365–368, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900646.

Ubhaya:1987:ALS

[Ubh87] V. A. Ubhaya. An O ( n ) algorithm for least squares quasiconvex approximation. Computers and Mathematics with Applications, 14(8):583–590, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901854.

Ungar:1985:DTA

[UM85] Abraham Ungar and H. C. Murrell. The differential transform and its application to an electrostatics image problem. Computers and Mathematics with Applications, 11(6):565–572, June 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122185900380.

Ungar:1982:NRL

[Ung82] Abraham Ungar. A note on Ranger’s line rotelet near a plane. Computers and Mathematics with Applications, 8 (2):95–102, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218290027X.

Urmantsev:1986:SSS

[Urm86] Yu. A. Urmantsev. Symmetry of system and system of symmetry. Computers and Mathematics with Applications, 12 (Part B)(1– 2):379–405, January/April 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901604.

Usmani:1985:SNF

[Usm85] Riaz A. Usmani. Some new finite difference methods for computing eigenvalues of twopoint boundary-value problems. Computers and Mathematics with Applications, 11(9):903– 909, September 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900938.

Usmani:1980:QSS

[UW80a] R. A. Usmani and S. A. Warsi. Quintic spline solutions of boundary value problems. Computers and Mathematics with Applications, 6(2):197–203, ???? 1980. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122180900280.

Usmani:1980:SSS

[UW80b] R. A. Usmani and S. A. Warsi.

Smooth spline solutions for boundary value problems in plate deflection theory. Computers and Mathematics with Applications, 6(2):205–211, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122180900292.

Vainshtein:1986:SBM

[Vai86] B. K. Vainshtein. Symmetry of biological macromolecules and their associations. Computers and Mathematics with Applications, 12 (Part B)(1–2):237– 269, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901537.

Vainshtein:1988:VSH

[Vai88] B. K. Vainshtein. A. V. Shubnikov and his ideas in modern crystallography. Computers and Mathematics with Applications, 16(5–8):351–356, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188902246.

Vanderbauwhede:1987:NPH

[Van87] A. Vanderbauwhede. Note on a paper by H. S. Qin. Computers and Mathematics with Applications, 14(3):185–187, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901519.

See [Qin86].

Venturino:1987:NIA

[Ven87] E. Venturino. Nonlinearly interacting age-dependent populations. Computers and Mathematics with Applications, 13(9– 11):901–911, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901726.

Verwer:1982:IES

[Ver82] Jan G. Verwer. Instructive experiments with some Runge–Kutta– Rosenbrock methods. Computers and Mathematics with Applications, 8(3):217–229, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900451.

Verwer:1986:SPB

[Ver86] J. G. Verwer. On the shift parameter in the backward beam method for parabolic problems for preceding times. Computers and Mathematics with Applications, 12(3):309–318, March 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122186901902.

Verheyen:1989:CSJ

[Ver89] H. F. Verheyen. The complete set of jitterbug transformers and the analysis of their motion. Computers and Mathematics with Applications, 17(1–3):203–250, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901600.

Vichnevetsky:1985:PSR

[Vic85]  R. Vichnevetsky.  Propaga-

tion and spurious reflection in finite-element approximations of hyperbolic equations. Computers and Mathematics with Applications, 11(7–8):733– 746, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901683.

Vlachonikolis:1986:EEP

[Vla86] Ioannis G. Vlachonikolis. On the estimation of the expected probability of misclassification in discriminant analysis with mixed binary and continuous variables. Computers and Mathematics with Applications, 12 (Part A)(2):187–195, February 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186900726.

Venkataramanan:1989:IMP

[VMO89] M. A. Venkataramanan, J. Mote, and D. L. Olson. An interactive MOLP procedure using epsilonconstraints. Computers and Mathematics with Applications, 17(12):1535–1543, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900539.

vanNiekerk:1987:NLO

[vN87] F. D. van Niekerk. Non-linear one-step methods for initial value problems. Computers and Mathematics with Applications, 13 (4):367–371, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900046.

vanNiekerk:1988:ROS

[vN88] F. D. van Niekerk. Rational onestep methods for initial value problems. Computers and Mathematics with Applications, 16 (12):1035–1039, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902593.

vanNiekerk:1989:GMU

[vNvN89] F. D. van Niekerk and A. van Niekerk. A Galerkin method using rational basis functions. Computers and Mathematics with Applications, 17(7):1085– 1093, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900382.

Volin:1985:ACD

[VO85] Yu. M. Volin and G. M. Ostrovskii. Automatic computation of derivatives with the use of the multilevel differentiating technique-1. algorithmic basis. Computers and Mathematics with Applications, 11(11): 1099–1114, November 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901889.

Vichnevetsky:1985:HON

[VP85] R. Vichnevetsky and E. C.

Pariser. High order numerical Sommerfeld boundary conditions: Theory and experiments. Computers and Mathematics with Applications, 11(1– 3):67–78, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901385.

Voss:1988:TSH

[VS88] D. A. Voss and S. M. Serbin. Two-step hybrid methods for periodic initial value problems. Computers and Mathematics with Applications, 15 (3):203–208, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218890171X.

Wong:1989:EEE

[WA89] P. J. Y. Wong and R. P. Agarwal. Explicit error estimates for quintic and biquintic spline interpolation. Computers and Mathematics with Applications, 18(8):701–722, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902277.

Wachspress:1984:GAP

[Wac84a] Eugene L. Wachspress. Generalized ADI preconditioning. Computers and Mathematics with Applications, 10(6):457–461, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122184900762.

Wachspress:1984:SLI

[Wac84b] Eugene L. Wachspress. Splitlevel iteration. Computers and Mathematics with Applications, 10(6):453–456, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900750.

Wachspress:1985:EIA

[Wac85] Eugene L. Wachspress. Editor’s introduction to Automatic indexed calculation of Wachspress’ rational finite element functions” by G. R. Dalton. Computers and Mathematics with Applications, 11(6):621–623, June 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185900446.

Wainwright:1981:TDP

[Wai81] Roger L. Wainwright. Three dimensional x-projection method (with acceleration techniques) for solving systems of linear equations. Computers and Mathematics with Applications, 7(3): 211–223, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218190081X.

Wainwright:1982:FDX

[Wai82] Roger L. Wainwright. Four dimensional x-projection method (with acceleration techniques) for solving systems of linear equations. Computers and Mathematics with Applications, 8(4): 329–337, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122182900153.

Walsh:1985:SQG

[Wal85] T. R. Walsh. On the size of quadtrees generalized to ddimensional binary pictures. Computers and Mathematics with Applications, 11(11):1089– 1097, November 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901877.

Walsh:1987:BRM

[Wal87] K. M. Walsh. Book review: Microcomputers and laboratory instruments: By D. J. Malcolme-Lawes. Plenum Press, New York, (1984). 246 pp. Computers and Mathematics with Applications, 14(9–12):947, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187902409.

Wang:1987:MNS

[Wan87] C.-A. Wang. Multiple numerical solutions of buoyancy induced flows of a vertical ice wall melting in saturated porous media. Computers and Mathematics with Applications, 14(7):527–540, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187900484.

Wang:1988:BIF

[Wan88] C.-A. Wang. Buoyancy-induced flows in porous media. Computers and Mathematics with Applications, 16(4):297–305, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901460.

Wasilkowski:1984:SNP

[Was84] G. W. Wasilkowski. Some nonlinear problems are as easy as the approximation problem. Computers and Mathematics with Applications, 10(4–5):351–363, ???? 1984. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122184900634.

Washburn:1986:PSC

[Was86] Dorothy K. Washburn. Pattern symmetry and colored repetition in cultural contexts. Computers and Mathematics with Applications, 12(3–4):767– 781, May/August 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186904232.

Wada:1988:AMA

[WAYU88] T. Wada, H. Akaike, Y. Yamada, and E. Udagawa. Application of multivariate autoregressive modelling for analysis of immunologic networks in man. Computers and Mathematics with Applications, 15(9):713–722, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901253.

Wu:1988:TPE

[WBZB88] J.-Y. Wu, V. Van Brunt, W.- R. Zhang, and J. C. Bezdek. Tower packing evaluation using linguistic variables. Computers and Mathematics with Applications, 15(10):863–869, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901228.

Werby:1985:SNT

[WCB85] Michael F. Werby and Stanley A. Chin-Bing. Some numerical techniques and their use in the extension of T -matrix and null-field approaches to scattering. Computers and Mathematics with Applications, 11(7–8): 717–731, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901671.

Wang:1986:GMQ

[WD86] K. Wang and P. J. Davis. Group matrices for the quaternion and generalized dihedral groups. Computers and Mathematics with Applications, 12 (Part 2)(5–6):1297–1301, September/December 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186902543.

Webb:1983:SAN

[Web83] G. F. Webb. The semigroup associated with nonlinear age dependent population dynamics. Computers and Mathematics with Applications, 9(3):487–497, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122183900627.

Webb:1986:LMS

[Web86a] G. F. Webb. Logistic models of structured population growth. Computers and Mathematics with Applications, 12 (4–5):527–539, April/May 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901781.

Webb:1986:RTT

[Web86b] Glenn Webb. Review of Trends in Theory and Practice of Nonlinear Differential Equations, Lecture Notes in Pure and Applied Mathematics: Vol. 90, Marcel Dekker, Inc., New York and Basel, 1984, V. Lakshmikantham, Ed. Computers and Mathematics with Applications, 12(4– 5):631, April/May 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122186901860.

Webb:1987:DSP

[Web87] G. F. Webb. Dynamics of structured populations with inherited properties. Computers and Mathematics with Applications, 13(9–11):749–757, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812218790160X.

Webb:1988:NPK

[Web88] W. A. Webb. A nonlinear public key cryptosystem. Computers and Mathematics with Applications, 15(2):81–84, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900776.

Webb:1989:CSS

[Web89]  G. F. Webb.  α- and β-

curves, sister–sister and mother– daughter correlations in cell population dynamics. Computers and Mathematics with Applications, 18(10–11):973–984, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122189900163.

Werner:1984:NLM

[Wer84] Wilhelm Werner. Newton-like methods for the computation of fixed points. Computers and Mathematics with Applications, 10(1):77–86, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900889.

Werner:1986:ISS

[Wer86] Wilhelm Werner. Iterative solution of systems of nonlinear equations based upon quadratic approximations. Computers and Mathematics with Applications, 12(3):331–343, March 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901926.

Worm:1981:CFF

[WFW81] G. H. Worm, C. D. Fiskeaux, and P. S. Wludyka. A computational form for finding the integral solutions to a system of linear equations. Computers and Mathematics with Applications, 7(5):405–406, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901267.

Woods:1982:LCN

[WH82] Dale Woods and Joel Huenemann. Larger Carmichael numbers. Computers and Mathematics with Applications, 8(3): 215–216, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218290044X.

Williams:1982:ICD

[Wil82] H. C. Williams. The influence of computers in the development of number theory. Computers and Mathematics with Applications, 8(2):75–93, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900268.

Wilson:1986:SLH

[Wil86] Dana Wilson. Symmetry and its “love–hate” role in music. Computers and Mathematics with Applications, 12 (Part B)(1–2):101– 112, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901446.

Williams:1987:MLA

[Wil87] G. Williams. The microcomputer in linear algebra. Computers and Mathematics with Applications, 14(1):43–52, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901805.

Wimp:1983:IRR

[Wim83] Jet Wimp. Irreducible recurrences and representation theorems for _{3}F_{2}(1). Computers and Mathematics with Applications, 9(5):669–678, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183901244.

Winkle:1985:I

[Win85] William A. Von Winkle. Introduction. Computers and Mathematics with Applications, 11(7– 8):ix, July/August 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901592.

Witten:1983:F

[Wit83a] Matthew Witten. Foreword.

Computers and Mathematics with Applications, 9(3):vii– viii, ???? 1983. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900184.

Witten:1983:HPD

[Wit83b] Matthew Witten. Hyperbolic partial differential equations: a few opening comments. Computers and Mathematics with Applications, 9(3):323–325, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122183900196.

Witten:1983:SFH

[Wit83c] Matthew Witten. On stochasticity in the von foerster hyperbolic partial differential equation system: Further applications to the modeling of an asynchronously dividing cellular system. Computers and Mathematics with Applications, 9(3):447–457, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122183900597.

Witten:1985:F

[Wit85a] Matthew Witten. Foreword. Computers and Mathematics with Applications, 11(1–3):xi– xii, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901336.

Witten:1985:M

[Wit85b] Matthew Witten. In memoriam. Computers and Mathematics with Applications, 11(1–3): xiii, January/March 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901348.

Witschi:1986:M

[Wit86] Werner Witschi. Moirés. Computers and Mathematics with Applications, 12 (Part B)(1–2):363– 378, January/April 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186901598.

Witten:1987:MA

[Wit87a] M. Witten. Math advantage. Computers and Mathematics with Applications, 14 (9–12):955, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902495.

Witten:1987:MMC

[Wit87b] M. Witten. Mathematical modeling and computers in medicine: Editor’s remarks. Computers and Mathematics with Applications, 14(9–12):xi–xii, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902203.

Webber:1983:ENL

[WJMM83] Bonnie Webber, Aravind Joshi, Eric Mays, and Kathleen McKeown. Extended natural language data base interactions. Computers and Mathematics with Applications, 9(1):233–244, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900172.

Wuyts:1988:RCC

[WLL88] P. Wuyts, R. Lowen, and E. Lowen. Reflectors and coreflectors in the category of fuzzy topological spaces. Computers and Mathematics with Applications, 16(10–11):823–836, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901927.

Wosilait:1987:NSI

[WLR87] W. D. Wosilait, R. H. Luecke, and M. P. Ryan. Numerical simulations of the infusion of adriamycin using a physiological flow model. Computers and Mathematics with Applications, 14(9– 12):861–868, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187902355.

Wu:1987:CEE

[WO87] S. R. Wu and J. T. Oden. Convergence and error estimates for finite element solutions of elastohydrodynamic lubrication. Computers and Mathematics with Applications, 13(7):583–593, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901192.

Wu:1987:ASP

[WSR87] H. Wu, P. J. H. Sharpe, and E. J. Rykiel, Jr. Age-size population dynamics: Derivation of a general matrix methodology. Computers and Mathematics with Applications, 13(9– 11):759–766, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901611.

Wu:1988:GFS

[Wu88] Y. Wu. The generating function for the solution of ODEs and its discrete methods. Computers and Mathematics with Applications, 15(12):1041–1050, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901393.

Wang:1988:MSS

[WW88] C.-A. Wang and S.-J. Wang.

Multiple similarity solutions of buoyancy induced flows for ice melting in cold pure water. Computers and Mathematics with Applications, 15(12): 1007–1017, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901368.

Wen:1989:ECM

[WW89a] K. A. Wen and J. F. Wang. Efficient computing methods for parallel processing: an implementation of the Viterbi algorithm. Computers and Mathematics with Applications, 17(12): 1511–1521, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900515.

Wendroff:1989:SSN

[WW89b] Burton Wendroff and Andrew B. White, Jr. A supraconvergent scheme for nonlinear hyperbolic systems. Computers and Mathematics with Applications, 18 (8):761–767, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189902320.

Xu:1987:STD

[XS87] S. Xu and S. Y. H. Su. A systematic technique for detecting and locating bridging and stuckat faults in I/O pins of LSI/VLSI chips. Computers and Mathematics with Applications, 13(5– 6):461–474, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900769.

Yachter:1982:EPS

[Yac82]  M. Yachter.  On the ex-

istence of periodic solutions of the D.E.: χ̈ + F (χ) = ϵf(χχ̇, t; ϵ), … (1). Computers and Mathematics with Applications, 8(4):257–271, ???? 1982. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900086.

Yakowitz:1989:SFM

[Yak89] S. Yakowitz. A statistical foundation for machine learning, with application to Go-Moku. Computers and Mathematics with Applications, 17(7):1095–1102, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189900394.

Yanik:1987:DMP

[Yan87] E. Greenwell Yanik. A discrete maximum principle for collocation methods. Computers and Mathematics with Applications, 14(6):459–464, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900575.

Yanik:1989:SCD

[Yan89] E. G. Yanik. Sufficient conditions for a discrete maximum principle for high order collocation methods. Computers and Mathematics with Applications, 17(11):1431–1434, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189900746.

Yavin:1984:NST

[Yav84] Y. Yavin. The numerical solution of three stochastic differential games. Computers and Mathematics with Applications, 10(3):207–234, ???? 1984. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122184900506.

Yavin:1987:PED

[Yav87a] Y. Yavin. Pursuit–evasion differential games with deception or interrupted observation. Computers and Mathematics with Applications, 13(1–3):191–203, ???? 1987. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122187901040.

Yavin:1987:STT

[Yav87b] Y. Yavin. A stochastic two-target pursuit–evasion differential game with three players moving in a plane. Computers and Mathematics with Applications, 13(1– 3):141–149, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/089812218790099X.

Yavin:1989:GTC

[YdV89] Y. Yavin and R. de Villiers. The game of two cars with a containment probability as a cost function: the case of variable speed. Computers and Mathematics with Applications, 18 (1–3):61–67, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901247.

Yousif:1986:IAS

[YE86] W. S. Yousif and D. J. Evans.

An iterative algorithm for the solution of a tridiagonal linear system of equations. Computers and Mathematics with Applications, 12 (Part A)(10): 1019–1027, October 1986. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122186900088.

Yousif:1987:AAA

[YE87a] W. S. Yousif and D. J. Evans.

Application of the AGE algorithm to the SLOR and ADI methods. Computers and Mathematics with Applications, 14 (3):225–228, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901544.

Yousif:1987:MAG

[YE87b] W. S. Yousif and D. J. Evans. The modified alternating group explicit (MAGE) algorithm for solving tridiagonal linear equations. Computers and Mathematics with Applications, 14 (3):169–172, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187901490.

Yee:1986:LFI

[Yee86] H. C. Yee. Linearized form of implicit TVD schemes for the multidimensional Euler and Navier– Stokes equations. Computers and Mathematics with Applications, 12(4–5):413–432, April/May 1986. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122186901720.

Yavin:1981:OCM

[YJ81] Yaakov Yavin and Alida M. Jordaan. Optimal controls that maximize the probability of hitting a moving target. Computers and Mathematics with Applications, 7(1):17–26, ???? 1981. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900043.

Ying:1987:UFC

[YK87] X. Ying and I. N. Katz. A uniform formulation for the calculation of stress singularities in the plane elasticity of a wedge composed of multiple isotropic materials. Computers and Mathematics with Applications, 14 (6):437–458, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187900563.

Young:1989:SAR

[YNA89] P. Young, C. Ng, and P. Armitage. A systems approach to recursive economic forecasting and seasonal adjustment. Computers and Mathematics with Applications, 18(6–7):481–501, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189901028.

Yu:1989:SCA

[Yu89] L.-X. Yu. Symmetry in Chinese arts and crafts. Computers and Mathematics with Applications, 17(4–6):1009–1026, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902733.

Yavin:1981:OCS

[YV81] Y. Yavin and A. Venter. Optimal control of stochastic systems with interrupted observations. Computers and Mathematics with Applications, 7(6): 509–525, ???? 1981. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181900353.

Yavin:1989:PNG

[YV89] Y. Yavin and R. De Villiers. Proportional navigation and the game of two cars: the case of a pursuer with variable speed. Computers and Mathematics with Applications, 18 (1–3):69–76, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189901259.

Yang:1987:RPH

[YWLB87] C. S. Yang, J. F. Wang, J. Y. Lee, and F. T. Boesch. Reliability properties of the hypercube network. Computers and Mathematics with Applications, 14(7):541–548, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122187900496.

Zung:1988:SFD

[ZA88] E. Zung and J. Angeles. Simulation of finite-dimensional linear dynamical systems using zero-order holds and numericalstabilization methods. Computers and Mathematics with Applications, 16(4):307–320, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122188901472.

Zadeh:1983:CAF

[Zad83] Lotfi A. Zadeh. A computational approach to fuzzy quantifiers in natural languages. Computers and Mathematics with Applications, 9(1):149–184, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122183900135.

Zadrozny:1989:ADE

[Zad89] P. A. Zadrozny. Analytic derivatives for estimation of linear dynamic models. Computers and Mathematics with Applications, 18(6–7):539–553, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189901065.

See errata [Zad92].

Zadrozny:1992:EAD

[Zad92] Peter A. Zadrozny. Errata to “Analytic derivatives for estimation of linear dynamic models”: Computers and Mathematics with Applications, Vol. 18, No. 6/7, pp. 539–553, 1989. Computers and Mathematics with Applications, 24(8–9):289–290, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/

                                        article/pii/089812219290207X.

See [Zad89].

Zamorzaev:1988:GA

[Zam88] A. M. Zamorzaev. Generalized antisymmetry. Computers and Mathematics with Applications, 16(5–8):555–562, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902453.

Zamani:1989:NQE

[Zam89] N. G. Zamani. A note on the quadratic element in one dimension. Computers and Mathematics with Applications, 18(12): 1019–1021, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900291.

Zlatev:1988:NTL

[ZB88] Z. Zlatev and R. Berkowicz. Numerical treatment of large-scale air pollution models. Computers and Mathematics with Applications, 16(1–2):93–109, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188900260.

Zedan:1987:SRT

[Zed87]  H. Zedan.  An AN-stable

Rosenbrock-type method for solving stiff differential equations. Computers and Mathematics with Applications, 13 (7):611–615, ???? 1987. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122187901222.

Zellner:1992:BCC

[Zel92] Arnold Zellner. Brief comment on C. A. Los’ papers. Computers and Mathematics with Applications, 24(8–9):275, October/November 1992. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812219290205V.

See [Los89b, Los89a].

Zeman:1985:USI

[Zem85] Marvin Zeman. Uniqueness of solutions to the initial-value problem for partial differential equations with multiple characteristics. Computers and Mathematics with Applications, 11(1–3): 293–305, January/March 1985. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122185901543.

Zenkin:1983:SEP

[Zen83] A. A. Zenkin. Some extensions of Pall’s theorem. Computers and Mathematics with Applications, 9(4):609–615, ???? 1983. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122183901189.

Zenkin:1989:WPF

[Zen89] A. A. Zenkin. Waring’s problem: g (1, 4) = 21 for fourth powers of positive integers. Computers and Mathematics with Applications, 17(11):1503–1506, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122189900837.

Zheludev:1988:STS

[Zhe88] I. S. Zheludev. Space and time symmetry. Computers and Mathematics with Applications, 16(5– 8):379–384, ???? 1988. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122188902283.

Zadunaisky:1980:IIA

[ZL80] P. E. Zadunaisky and G. Lafferriere. On an iterative improvement of the approximate solution of some ordinary differential equations. Computers and Mathematics with Applications, 6(1):147–154, ???? 1980. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122180900668.

Zlatev:1982:CTP

[Zla82] Zahari Zlatev. Comparison of two pivotal strategies in sparse plane rotations. Computers and Mathematics with Applications, 8(2):119–135, ???? 1982. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122182900517.

Zlatev:1988:SLS

[ZN88] Z. Zlatev and H. B. Nielsen. Solving large and sparse linear leastsquares problems by conjugate gradient algorithms. Computers and Mathematics with Applications, 15(3):185–202, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122188901708.

Zouni:1989:SCS

[Zou89] O. Zouni. Space ⇀ ↽ color ⇀ ↽ symmetry. Computers and Mathematics with Applications, 17(4– 6):1027–1040, ???? 1989. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122189902745.

Zvilna:1989:AI

[Zvi89] J. Zvilna. Ad infinitum. Computers and Mathematics with Applications, 17(4–6):1041–1066, ???? 1989. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122189902757.

Zvyagin:1988:PCS

[Zvy88] B. B. Zvyagin. Polytypism of crystal structures. Computers and Mathematics with Applications, 16(5–8):569–591, ???? 1988. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122188902477.

Zlatev:1985:ESS

[ZWS85] Zahari Zlatev, Jerzy Wasniewski, and Kjeld Schaumburg. Exploiting the sparsity in the solution of linear ordinary differential equations. Computers and Mathematics with Applications, 11(11): 1069–1087, November 1985. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122185901865.